A Complete Bibliography of Publications in the Electronic Journal of Linear Algebra (ELA)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 February 2015 Version 1.01

Title word cross-reference

(0, 1) [BH04, DS10, HP12a, MNP03]. (β) [Mec12]. (C) [Mec12]. (k, τ) [Car12]. (m + 1)^{2} − s [Sne03]. (P, Q) [MDKM14, ZCL10]. (v, k, λ, μ) [Elz03]. −2 [WSH12]. {1, 2, 3} [LY11]. {1, 2, 4} [LY11]. {1, 3} [CI12]. {1, 4} [CI12]. 2 [BGRV09, Bea99, EEH^{+}12, FKS14, MN10, NBdP11, XTC10]. 2n [Bay11]. 2 × 2 [CM05, MVPS09, MNP03]. 3 [AMS09, BSZ10, Kal12, RSS13]. 3 × 3 [LvdD14, CN10]. 4 [AHL^{+}10, CV07, tCN12, GS11, OV10, dFNP13]. 5 [dFNP13]. 5 × 5 [LMS04]. A [MT11, Ver08]. A^{∗} [Ver08]. A^{T} [Ver08]. A^{T} XA = D [XHZ09]. A_{T,S}^{(2)} [DX12]. AB [LS09]. AX = B [fXyHZ09].

AXB + CXD = E [fYwW12]. AXB + CY D = E [WZY08]. BA [LS09]. Ā [Ver08]. Z [Gow12]. C [Liu12, Uhl98]. C^{∗} [MD11a, MD11b, SW11b]. C [RV09]. χ [CF14a, CF14b]. D [ÁAFG12, NWZ98, ZCB03]. † [LHL11]. ℓ_{p} [Sch11a]. G [ZX08, AG12]. GF(2) [Bos10, Bea99]. GF(2^{2}) [OK09]. GL(2, R[x]) [Fay00]. H [FGVR14, ZzH06, ZLXX09, ZLHL10, ZZLL12]. J [GP03]. K [BT13, CP99, LZZ13, LD11, Liu12, LWYJ13, AM14, Bay11, CF14a, CF14b, GL10, Gon12, XY14]. {K, S + 1} [LT12]. K_{2,3} [YSH12]. L [MPT00]. λ [Elz03]. λ_{1} [LHY09]. LDL^{T} [Ova04]. LDU [BP12a]. M [BP12a, BPS05, CEM12, CLW12, GJOvdD07, LLYL11, Nab00, NX06, OTvdD09, SGL12, SST04, Stu98, TH10, ZZL11, Psa02]. M_{∨} [OTvdD09]. R^{N} [AG12]. R^{P} [KS09]. Z [Bal11]. μ [Elz03]. N [BSZ10, Liu12]. n = 4 [Dru04]. P [Bal11, DHHW03, JK96, BHL09, ZZ11]. P^{α} [HM98]. P_{0} [CDH^{+}02, CDH^{+}03]. PG(⋀^{K} V) [De 10]. PG(V) [De 10]. π [Gil07b]. Q [BP07, DN12, DHS09, Nik14, NY14, dAN13, dFNP13, CN10, MH96]. QR [SW11a]. R [BH04, BF08]. S [BPS06, BH04]. SL(2, C) [PLBG12]. sl_{2} [AC10]. SL_{n}(F [x]) [Ero04]. ⋆ [DD11a]. TP_{K} [AJN14]. U_{q}(ŝl_{2}) [AC05]. W [Son13]. X = Q + A^{H}(I ⊗ X − C)^{(} − δ)A [YLD10]. X^{2}AX = AXA [LM10]. XA + AX∗ = 0 [DD11b]. XAX = B [KL13]. ξ [Qi13]. Z [MOS^{+}98, RHC06].

* [CHMN15, FKS14, HSSM05]. *-algebras [CHMN15]. *-congruence [FKS14, HSSM05]. *-ideals [CHMN15]. *congruence [DD11b].

- [LY11]. -adic [MH96]. -ads [Uhl98]. -algebras [MD11a, MD11b]. -by- [tCN12, FKS14, RSS13]. -colored [Kal12]. -commuting [RV09]. -competition [SGL12]. -congruence [DD11a]. -cycle [dFNP13]. -cyclic [Liu12]. -D [BGRV09]. -dimensional [MN10, NBdP11]. -eliminated [Bay11]. -extremal [LHY09]. -factorization [Ova04]. -generalized [CI12]. -index [DN12, Nik14, NY14, dAN13, dFNP13]. -inverses [LY11, MDKM14, ZX08]. -involution [LD11]. -Laplacian [BHL09, ZZ11]. -Lie [BSZ10, Qi13]. -like [MT11]. -matrices [HP12a, MNP03, BP12a, BPS05, BH04, DS10, GJOvdD07, HM98, Nab00, NX06, OTvdD09, RHC06, Stu98, TH10, ZzH06, ZLXX09, ZLHL10, ZZLL12, ZZL11]. -matrix [CEM12, CLW12, CDH^{+}02, CDH^{+}03, DHHW03, DHS09, JK96, LLYL11, SST04].

-maximal [Gon12]. -minor [YSH12]. -modules [SW11b]. -monotone [CP99]. -numerical [CN10]. -optimal [ÁAFG12, NWZ98, ZCB03]. -orthogonal [ZCL10]. -Pencils [MOS^{+}98]. -positive [FGVR14]. -potent [BT13, LT12]. -Property [Bal11, BP07]. -regular [BF08, Car12]. -shape [LHL11]. -spaces [Sch11a]. -spectral [GP03]. -strictly [BPS06]. -symmetric [CF14a, CF14b]. -th [CF14a, CF14b, Psa02]. -Transformations [Bal11, Gow12]. -trees [EEH^{+}12]. -triangular [Gil07b]. -tridiagonal [AG12]. -weighted [Son13]. -ycle [dFNP13].

2008 [MST09]. 2011 [BBCP12].

3rd [HZ10b].

absolute [Gil07a, Roh09a]. Absolutely [JHH^{+}03]. abstract [DN12]. Accurate [BP12a, MM13]. action [Sze07]. actions [D̄T01, Qi13]. acyclic [BF04, BOvdD05]. Additional [WW01]. Additive [AH10, Lim12, OK09, CILW11]. Additivity [CJ08, LX11]. adic [MH96]. adjacency [FHJT96]. adjoint [FKS08]. adjointable [SW11b]. adjoints [KM11]. ads [Uhl98]. Affine [NT07, AC05]. agent [SST04]. aggregation [Pul12]. aggregation-disaggregation [Pul12]. algebra [AC05, BS05, CILW11, Dub08, GKMS13, HZ10b, Kol00, PLBG12, RV09]. Algebraic [BF05, Bra09, Fen01, Böt11, BSS14, DD10, FK98, HM12a, Kir01, KM05, LZL10, MNS00, MN01, NS12, RMdAJ10, Sch11b, SGL11, TT07, ZL11]. algebras [BSZ10, BDL^{+}09, BS07, Cha12, CJ08, CHMN15, Gol08, LX11, Mar14, MM12, MD11a, MD11b, NR11, PM12, Qi13, SZ09, Tao11]. algorithm [D̄SZ03, NX05, Roh09a, Sol02]. algorithms [MMD05, SW11a]. Allow [CF12, BCD^{+}09, CHOvdD09, CEH^{+}12].

Almost [Kol00]. alternant [Chu06]. alternative [GP03, Roh13]. Alto [MST09]. American [MST09]. among [hLSxS12]. analogy [Lin13a]. Analysis [BBCP12, HM12a, CGMSR11, DX12, gLgWhX13, Ova04, Rol03, See13, HZ10b]. Analytic [RS05]. Anderson [Lin13b]. angle [GSM14, KKP09, See13]. anti [XHZ09]. anti-symmetric [XHZ09]. Aperture [See13]. application [Bru02, KKP08, Kol00, NS12, YFW14, ZCB03]. Applications [BBCP12, BS05, MMW12, MST09, AV06, BP12b, BTM07, BL14a, CMRR13, HK98, KDTP12, LM98, LLZ13, MNS11, Per02, SS08b, Son13, TL07, HZ10b]. applied [JvdB12]. approach [KKP08, Kir04a, Sch11b]. approximate [HL07]. Approximating [Siv05]. approximation [Sch12, SA12, fXyHZ09, ZCL10]. approximations [FMMN11]. arbitrary [AHL^{+}10, CMRR13, CG13, CKSV05, CV07, CGK^{+}13, CDH^{+}07, DOvdD10, GSF09, GS11, Per07, SK00]. area [AGV11, MR12]. area-preserving [MR12]. arising [BLdP00, ZLLX12]. Arnol’d [vdH10]. array [Sch11b]. arrow [PES09, PES11]. associated [tCN12, LT12, MR12, MS04]. associative [BDL^{+}09, Gol08]. asymmetric [Xu08]. Atiyah [D̄ok02]. attached [CMRR13]. attain [hLSxS12]. automorphisms [YY13]. autonomous [Fen01, MR12]. avoidance [SST04].

backward [TG06]. balanced [FRM12, MNS00]. Banach [CIW09, CILW11, DX12]. Banachiewicz [TT05]. band [BG99]. banded [Bap07]. Bannai [Bro13]. Bannai/Itô [Bro13]. base [YY12]. based [OTvdD09]. Basic [Woź01]. basis [AC10, WWY10]. be [Cio10]. behaviour [Nof12]. Berkowitz [Sol02]. Bernstein [MM13]. Bessel [KL15]. Bethe [RM09]. between [GSM14, HM12b, PT00, Tlu07, Ver08, XTC10]. Bézout [CL09a]. Bézoutians [WWY10]. Bi [SK00]. Bi-CG [SK00]. Bicyclic [SHZ12, LL12, YFW14, Zhu11]. Bidiagonal [BP12b]. bidirected [BLP09]. bilinear [Sze05]. Binary [Zho12, MNS00]. biorthogonal [Mat03]. bipartite [BLS05, BOvdD05, Bro13, COvdD09, Cio10, hLSxS12, LWYJ13, MKR13]. bipartite/ abipartite [Bro13]. biquadratiques [El 11]. Birkhoff [Mar06]. Bishop [Mec12]. bivariate [BS12]. Block [BB07, COvdD09, KS11, PS13, BCR13, BZZ09, BZ10, BSZW12, CMS12, CvdD12, Gil07b, JN09, LvdD14, Sin12, TS01, Tia05, Tia10a, ZzH06, ZLHL10, ZB10]. block-matrix [TS01]. blocking [GT05]. Boolean [BF10]. bordered [BZ03]. bosons [BLdP00]. bound [BGRV09, BF11, BGH15, CLW12, Gil14, HK98, Kir01, Kir07, LZL10, Yu10, ZHS08, dF10]. boundary [CEG12, Loi12]. Bounded [Rod06, AKS12, CKLN14, CIW09, HP12b, LR98b, dAN13]. boundedly [Ero04]. Bounds [DFS13, HZ10a, NO03, Sch11a, YSH12, ZzH06, EG10, FFP12, LLYL11, LHL14, MC13, Mat08, MKR13, MNS00, Nie06, Sha12, XCL11, ZL11, DD08]. Boyle [GN02]. Brauer [LL11]. Braun [FR04]. Brualdi [BT12, Wan08]. Bruhat [Bru02, BH04]. Brunovsky [Jin05]. Burke [Kar06].

CA [MST09]. cacti [SWK14]. calculations [MPSS07]. can [God12]. Canonical [MMX00, BH13, DD10, FKS14, KKK11, KL13, MX99, XLJ13]. cardinality [SVP11]. carries [D̄SZ03]. case [BvdHL05, BK97, BOvdD04, Dru04, Fer09, JK11, MN10]. cases [BNSY14]. caterpillar [RMdAJ10, SGL11]. caterpillars [RMdAJ10]. Cauchy [Chu06, HK98]. Cayley [CL04, FT02]. Certain [MSS08, FFP12, LR02, Nie06, RMdAJ10, Stu98, YF11]. certification [SSV04]. CG [SK00]. chain [BL14b, GKMS13, Kir03, Kir04a, Kir14, Xu08]. Change [FKS14, DS07]. characterisation [Dic10]. characteristic [BDL^{+}09, FS98, Gon12, KM14, MPT00, XTC10, dF05]. Characterization [Bal11, KM05, Son13, LHC09, LHL11, Pek14, Roh10, Sci07, Xu10, YY12]. Characterizations [AH10, LMT10, AG14, CRU10, FR04, LZ14, Rum02]. characterized [Bou09]. Characterizing [Qi13]. Chatelier [FR04]. Chebyshev [LA02]. Checking [BO00]. Chief [Sha10, Sha13]. Choi [Nie13]. chordal [FM13]. circle [BF07, Per02]. circulant [BF10, LR13]. circum [LMT10]. circum-Euclidean [LMT10]. Class [BS07, BGRV09, BH04, BZ10, El 11, FKS08, HSSM05, Nab00, SS08b, SH12, ZLLX12]. classe [El 11]. Classes [BdPdP10, BLR01, CFP11, FT02, KM05, RMdAJ10, Stu98]. classic [LŠ08]. classical [Ber03]. classification [Ari02, Meh06b]. classifying [BDL^{+}09]. clique [JN09, MÇ12, SH08, dAN13]. closed [Shu11]. clow [Sol02]. cocycles [FG14]. cocyclic [ÁAFG12]. codimension [DKS14]. codimensions [KKK11]. coefficient [KKP08]. coefficients [Böt11, Brü09, FFP12, PP08]. Colin [BFH05, FM13]. Collatz [Oep03]. colored [Kal12]. coloring [Bru02]. Column [EPM13, BH04, DMLT03, Kir09]. columns [FS98]. combination [TCID11]. combinations [LWB11, Nik06]. Combinatorial [Dah07, BLM^{+}13, BNSY14, Kir04a, NM02, SS08c, KSvdD05]. combinatorially [JK96]. common [MS04]. commutant [JdGM05]. Commutative [CMS12, HKT14]. commutativity [DK08, JdGM05]. commutator [Bou11]. Commutators [dSP14]. commutes [Bou11]. Commuting [Han05, JW12a, XY14, BKNO07, DKO12, DKO13, Gil05, Gil11, MT10, RV09, TCID11]. compact [YY13]. companion [AV04, Aur13, KV04, KV14, vdDW96]. Comparison [CP99, Rod07, Woź01]. competition [SGL12]. complement [LNP14, YF11, ZZLL12]. complementarity [BP07, Roh09b]. complementation [BNS14, NX05]. complements [BOvdD05, FV08, LH10, RHC06, TT05, ZLXX09, yZLXJ09]. complete [BLM^{+}13, BBC^{+}14, Bou09, HKT14, Nof12, dF03]. completely [Dic10, DS08, Gow12, GS13]. Completing [Tia10a]. completion [AJN14, CDH^{+}02, CDH^{+}03, DHHW03, DHS09, FJS00, JK96, JK11]. completions [BK97, CH13, JN09, JTU98]. complex [AT06, Cof08, FGVR14, GSF09, GR09, GHK14, Gus13, HPP09, LLMZ12, Psa02, Rod07, Rum02, SSM10, SZ09, vdAtMM07]. complex-valued [vdAtMM07]. component [AGV11, Rol03]. components [DKO13]. composition [LLMZ12]. compound [LH10, RM09]. compounds [BLdP00]. Computation [vdAtMM07, HD10, PS13]. computations [MM13]. computer [BdPNdGPS08, LR98a]. Computing [KKK11, Jin05, KP08, LZFR12, NX05]. con [Ver08]. concave [DFS13, Per02]. concavity [KM14]. concerning [CF12]. condition [Bid07, BG99, Cio10, FR04, Gil14, Kar06, TG06, WZS09, XLJ13, ZX08]. Conditioning [Kir03, BF07, Dav04, Kir04a]. Conditions [JK11, CI12, DLRM^{+}13, LD11, Roh09c]. cone [Bal11, Dic10, DS08, JG14]. coneigenvalues [DL12]. cones [Gow12, GS13, MS04]. Conference [BBCP12, FFB12]. configurations [D̄ok02]. congruence [DD11a, D̄SZ03, FKS14, HSSM05, Sze07]. congruences [Rod07, Ver08]. conjecture [D̄ok02, Dru04, Găv13, GN02, JTU98, Wan08, dF05]. conjectures [BF04]. connected [AGV11, Cio10, DKO13, MN01]. connections [Dod07]. connectivities [RMdAJ10]. connectivity [BF05, FK98, Kir01, MNS00, MN01, SGL11, vdH10]. Consistency [DD11a, LS98]. constant [Bos10, Brü09, Kir14]. constrained [LS98, Rol03]. constraint [LX14, PL12]. constraints [FK98, ZJW11]. Constructing [JR08, Pol10]. construction [Gus13, LM10]. Constructions [Rea02]. containing [HZS06]. continuous [Dra11, LZL10, ZL11]. contraction [Ben02]. contribution [Oep03]. control [KM11, SST04]. Controllability [Dod07]. controllable [CFP11]. convergence [Hol00, LA02, Mat08, Uhl98]. converging [ERS09]. convex [MS04]. convexity [KM01, NBdP14]. Copies [CMRR13]. copositive [GSM14, Gow12, Jar13, JR08]. coprime [BCG96]. core [BSS14]. corona [HS10]. counting [CL04]. counts [DKS14]. Courant [Tlu07]. covariance [ZCB03]. cover [BFMN11, CCH^{+}12, WZ12]. CP [SM01, SM04]. CP-rank [SM04]. Craig [Mar08]. Cramer [JvdB12]. criterion [Sur00, yZLXJ09]. Critical [JW12b, LNP14]. Crossing [FRM12]. curvature [DD10]. curve [Fen01, TT07]. customer [GT05]. cut [DZ11, LZZ13]. cutpoints [Kir01]. cycle [JK11, Kal12, dFNP13, vdH09]. cycles [YWZ12]. cyclic [CvdD12, FW10, FN98, Hog09, Lin13a, Liu12].

d [Sch12, BGRV09]. DAE [KM11]. DAES [CH13]. Davis [Nie13]. December [MST09]. Decomposition [Rak12, BP12a, Ben10, BL14a, Bru02, CMS00, EPM13, Meh06a, SS08a, Stu98, TMR06, XLJ13, vdDW96]. decompositions [BP12b, BCG96, De 13, HPP09, Mol09]. decoupling [CM00]. Dedication [The98]. defined [Nie06]. définies [El 11]. definite [Dru09, WZ08, YLD10, vdH09]. definiteness [CN98]. deflating [MT04]. degenerate [Nik14]. degree [BHL09, BF12, OV10, SI10, ZZ11, Zhu12]. degrees [JS02]. delay [HM12a]. deleting [Lot07]. dependence [KLR11]. derivable [AH10]. derivations [AH10, Qi13, XZ10]. derivative [CF14b, Sch11b]. Derivatives [AGV11, BGJ13, CF14a, vdAtMM07]. derived [SZ09]. Description [Roh09b]. descriptor [Brü09, CM00]. design [ZCB03]. designs [ÁAFG12, FRM12, NWZ98]. Determinant [CT04, CD00, MPSS07]. determinantal [Lin14]. Determinants [FMS12]. determined [LLW11, MK10, OV10, SH12]. determining [CKSV05]. Deux [Ben02]. Diagonal [BMN^{+}13, BS05, GHH^{+}14, HS03]. diagonalizable [Gil10]. diagonally [BP12a, BCR13, BPS06, Mor03, ZLXX09, yZLXJ09, ZLHL10, ZZLL12, ZH07]. diagonals [CL09b, LMR11]. diameter [AGV11, LL08a, hLSxS12]. Diameters [DKO13, KSH11]. Diaz [MNS11]. dichotomy [AS06]. difference [Loi12, Rod06, Sin14]. differences [Chu06]. differential [HM12a, KM05, NS12, Sch11b, ZL10]. differential-algebraic [HM12a, KM05, NS12, Sch11b]. digraph [Sne03]. Digraphs [KOvdD00, BCH^{+}13, BOvdD04, Kal12, KSH11, LR13, SGL12, SHZ12, YY12]. Dihedral [YY13]. dimension [Ber03, BS07, Tlu07, DD08, DD11b]. dimensional [Gol08, MM12, MN10, NBdP11]. diods [HKT14]. Dirac [FKS08]. Directions [FFB12, KSvdD05]. disaggregation [Pul12]. discrepancy [Dah03]. discrete [BHL09, Brü09, FKS08, GR09]. discrete-time [Brü09]. discs [KS11]. Disjoint [Bou09, Kol00]. Distance [SI10, BB07, BLP09, BS12, CEG12, JM13, Kir07, LMT10, Zha14]. distance-regular [CEG12]. distances [DKO12]. distinct [AAC^{+}13, dF10]. distortion [BLR01]. distribution [Kir03, Kir04a, Kir14, ZLXX09, ZLHL10]. distributions [Sch12]. Disturbance [CM00]. divided [Chu06]. division [Cha12, KDTP12]. divisors [CS09, SC08]. domain [BLS05, CL09a, LA02]. dominant [BP12a, BPS06, Mor03, ZLXX09, yZLXJ09, ZLHL10, ZZLL12, ZH07]. double [Chu06]. doubly [LZ12, MMX00, PES09, ZH07]. Drazin [BZ10, CGMSR11, COvdD09, CvdD12, CIW09, CILW11, DW10, GKW02, KKP08, NX06, SK00, Son13, WW01]. Drazin-inverse [SK00]. dual [KL15]. duality [GS13]. Dunford [Mec12]. dynamic [ZJW11]. dynamics [WS12].

Edge [LLZ13, BBC^{+}09, BF05, HS10, MN01]. Edge-grafting [LLZ13]. edges [DZ11, LZZ13, Sne03]. Editor [Sha10, Sha13]. Editor-in-Chief [Sha10, Sha13]. Efficient [HD10, BP12a, DD11a]. eigen [Uhl98]. eigenproblems [MMD05, PES09]. eigenspace [BT12]. eigenstructure [Nof12]. eigensystem [vdAtMM07]. Eigenvalue [CH13, GHK14, Kar06, KV04, AT06, AKS12, BNPS08, BF04, BGK^{+}11, BLM^{+}13, BNSY14, BF11, CLW12, EG10, Fer09, HP12c, Hog05, Jar13, JS02, LŠ08, LLYL11, LL08b, LhLY09, LWYJ13, gLgWhX13, MS14, MKR13, MP09b, Oep03, Pat01, Rea02, SH10, TH10, Tlu07, WSH12, YGX13, YFW14, ZLXX09, ZLHL10, ZJW11, Zhu12, vdAtMM07]. Eigenvalues [Dod08, Fot10, Kou06, AAC^{+}13, BdPdP10, BHL09, CL09b, Cha12, DH13, HJS02, HZS06, Kir04b, Kir06, Kir07, Kir09, KM01, Laf04, LPW14, LLS14, LZL10, LR98b, MVPS09, MC13, MX08, Nik06, Sin14, Tam10, dF10]. eigenvector [Cio10, HL07, vdAtMM07]. eigenvectors [CG07, Kou06]. elasticity [NO03]. electronic [HZ10b]. elementary [CS09, Ero04, SC08, Sur00]. elements [Che12, MD09, MD11b]. eliminated [Bay11]. elimination [Dah07]. ellipsoids [See13]. elliptic [LA02]. elliptical [RSS13]. Embedding [ÁAFG12, LWZ04]. enclosures [HD10]. energies [LL08a]. energy [DS07, GLX14, SHZ12, Tri13, XD13]. entries [FJS00, FHR04, LZ12, MSS08, ZZ14]. entropies [KN14]. entry [AJN14, BMN^{+}13, Kir04a]. EP [BL14a, Che12, MD09, MD11b]. equal [BBC^{+}14]. equality [Tia12]. equals [MS14]. equation [AS06, Böt11, DD11a, GL11, HD10, KL13, LM10, LZL10, LX14, Mar08, PL12, RR02, Roh09a, WZ08, WZY08, WJ10, XHZ09, fXyHZ09, YLD10, fYwW12, ZL10, ZL11, ZCL10, DD11b]. equations [DKS14, HM12a, HKT14, HSSM05, JvdB12, KM05, LR02, LD11, Loi12, RMZ14, Rod06, SW11b, WZS09]. equilibrium [MMW12]. equitable [AC10]. Equivalence [Sze07, Cof08, CFP11, HP07a, HS03]. equivalences [Rod07]. equivalent [HP07b, ZX08]. ergodic [Sch12]. Erratum [LhLY09]. Error [Pul12, CMS12]. errors [TG06]. Essential [Meh06a]. Essentially [Dru06]. establish [TS01]. Estimates [Gil07a, Gil05, Gil11, YH05, ZL10]. Estimation [JS02, JPS12, PS11]. Estrada [DZ11, SWK14, Sha12]. Euclidean [Fie98, Fie05, JM13, LMT10, Tao11]. evaluations [MPSS07]. Even [DMS11, Wan08]. Eventual [HW12, BCD^{+}09, CEH^{+}12, EHT09, Hog09]. Eventually [Hog09, NM02, OTvdD09]. Exact [FFJM14, LX14, HD10]. exactly [Jar13, Kal12, SH10]. example [LŠ08]. existence [LMS04]. expansion [FN02, FS00]. expansions [NO03]. Explicit [Brü09, HPP09, Roh11, vdDW96, BZ10, Ero04, LM10]. exponent [BF08, KOvdD00]. exponential [TMR06]. exponentially [BG99]. Exponents [KSH11, BF10, JW12b]. extended [LX14]. Extension [KL15, CGMSR11, Tam99, ZLLX12]. external [Dru04]. Extremal [RMdAJ10, XD13, LHY09, SHZ12]. Extreme [PES11, WJ10]. Extremizing [FK98].

factor [BOvdD04]. factorization [Aur13, BH13, CRU09, GP03, Ova04]. factorizations [HP07b, TS01, Zho12]. families [Kol00, MMT11]. family [AC05, AV04]. Fast [FMMN11, KP08]. feasible [Elz03]. feedback [CM00]. Feichtinger [Găv13]. few [FJS00]. fewer [BJL^{+}10]. Fibonacci [MSS08, TMR06]. Fiedler [BF12]. field [BDL^{+}09, CDH^{+}07, DGH^{+}09, OK09, Son13]. fields [BvdHL05, BZZ09, CVV11, DK06, DMS11, Gro10, VV09, XTC10, ZB10]. find [HL07]. finite [BvdHL05, BDL^{+}09, DK06, DMS11, Gol08, Gro10, Gus13, VV09]. finite-dimensional [Gol08]. finitely [D̄T01, Elz03]. first [Rod06]. Fisher [Böt06]. five [NWZ98]. fixed [HKT14, SI10]. fixed-point [HKT14]. Flanders [CRU09]. flat [JHH^{+}03]. flexible [JvdB12]. Focal [HM05]. foci [LA02]. Forbidden [dFNP13]. forcing [BBC^{+}14, BCH^{+}13, BGH15, CCH^{+}12, EEH^{+}12, Pet12]. forest [VY12]. form [FKS14, GN02, JTU98, MMT11, Shu11, Sze05]. Formal [KM11]. formed [Dru09]. formes [El 11]. forms [AV04, El 11, GG04, Jin05, KKK11, KL13, LS09, MMMM11, MMX00, MX99, TT05, Tia05, YFW14]. formula [CL04, Kar06, Pul12]. Forty [Roh09c]. four [NWZ98, Tia12, VY12]. four-vertex [VY12]. Fourier [Dah07]. fraction [BCG96]. frame [KL15]. frames [RW02]. free [Bri07, YY12, YSH12]. French [Ben02, El 11]. Frobenius [ES08, ERS09, Per02]. Full [CRU09, JdGM05, KOS^{+}09, LMS04, XZ10]. fully [FN98]. function [BS12, DFS13, MS04]. functional [Bay11]. functionals [Nie06]. functions [Dav04, Dra11, ES10, Gil05, Gil07a, Gil10, Gil11, GP03, Maz10, PTZ14, RS05, RS04, SS08b]. fundamental [TMR06]. Further [Mar08, SS08c].

gap [Sta13]. gaps [BF10]. GCD [Ova04]. General [Cha12, DK08, ES08, FJS00, Mol11, TT07, WWY10, ZLXX09, ZZLL12, vdAtMM07]. generalization [BTM07, MÇ12, Mat03, NS12, OTvdD09]. Generalizations [LL11, MST09]. Generalized [BZ03, Bay11, BLdP00, LH10, MP09a, Bap07, CILW11, CI12, DW10, DS13, ES10, HJS02, KKK11, KKP08, KP11, LX11, Nie13, PTZ14, PS13, RHC06, RM09, RS04, SMB03, Tia05, TL07, Tia12, TCID11, Tya12, WZS09, XLJ13, XQ11, XZ11, ZZLL12, ZJW11]. generally [yZLXJ09]. generated [Ero04, SM04]. Generating [KOS^{+}09]. generation [BdPNdGPS08, LR98a]. generic [KKK11, Nof12]. genus [YY13]. Geometric [CFP11, CN09, Pol10, VU12]. geometry [Fie05, Hen10, MS04]. Gersgorin [JPS12]. Gersgorin-style [JPS12]. Gershgorin [KS11]. Gershgorin-type [KS11]. Girth [Kir04b, Zhu11]. given [DMLT03, DZ11, ERS09, HP12a, Kir09, LL08a, hLSxS12, LHC09, SH08, XD13, YWZ12, ZZ14, Zhu11, Zhu12]. Graded [Mar14, PM12]. grafting [LLZ13]. graph [BS12, BFH05, BBC^{+}14, Bru02, BGH15, CMRR13, Cio10, DS07, DN12, DGH^{+}09, DH13, DKO12, EG10, FK98, GS03, Hog05, JS02, LZ12, MKR13, MK10, NO03, Nik06, RV09, Tlu07, Yu10, Zhu12, dF03]. Graphs [BvdHL04, BvdHL05, BEH^{+}14, LLW11, Sta13, AAC^{+}13, AKS12, BNPS08, BFH05, BFMN11, BBC^{+}09, BJL^{+}10, BGK^{+}11, BLM^{+}13, BNSY14, BF11, Bou09, BOvdD05, But07, CEG12, Car12, COvdD08, CCH^{+}12, CL07, CG07, CL04, DKO13, DZ11, EFP12, Elz03, FM13, Fie05, FN02, FS00, Fri12, GL10, GLX14, HP12c, HS10, JM13, JN09, KM14, Kir01, KS08b, LR13, hLSxS12, LLZ13, LZZ13, LNP14, LLS14, LL08b, LHC09, LhLY09, LHY09, Liu12, LL12, LWYJ13, LZ14, MS14, Nik14, NY14, RM09, RRCM11, Sci07, Sha12, SH12, SGL11, Siv05, SH08, Tri13, WSH12, XD13, YL11, YSH12, YGX13, YFW14, YWZ12, Zha14, Zhu11, dAN13, vdH09, vdH10]. Grassmann [KKP09]. Group [CL09a, CG13, COvdD08, Mor11, ZB10, BZZ09, BSZW12, D̄T01, Fay00, LT12, LWB11, NX05, Sze05, TL07]. groups [BDL^{+}09, Nie06, YY13]. growing [BG99]. growth [TT07]. GUT [AM12]. GUT-majorization [AM12]. gyroscopic [Lan99].

Hadamard [ÁAFG12, BNS14, CLW12, HK98, HZ10a, LLYL11, MP00, Sch11a, YF11]. Hamiltonian [LZFR12, MX08, Rod08a, Rod14b, SW11a]. Handelman [GN02]. Hankel [FN06]. Harmonic [Mor13, VU12]. Hartwig [Böt06]. having [ES08, MSS08, RRCM11]. Hawkins [Bid07, FR04]. Herglotz [SS08b]. Hermite [KDTP12, SMB03]. Hermitian [BdPdP10, DK08, Dru06, Fot10, IT12, Lim12, Mat08, MX99, OK09, Rod08b, Rod14a, Tia10a, Tia10b, Tia12, WJ10]. Hermitian/skew [IT12]. Hermitian/ skew-Hermitian [IT12]. Hessenberg [DS10]. hidden [Mor03]. Hilbert [Dra11, SW11b, Sur00]. Hom [PM12]. Hom-algebras [PM12]. homogeneity [GS13]. homomorphisms [JG14]. Hopf [BH13]. Horn [CL09b, Tam10]. Horner [TMR06]. Hua [Lin13a]. hulls [AMS09]. Hurwitz [Lin13a, Tya12]. Hyers [HM12b]. hyperbolic [BTM07]. hypercube [MT11]. hypergeneralized [TCID11]. hypergraphs [Ref14]. hyperplane [dSP14]. hypo [BSZ10]. hypo-nilpotent [BSZ10].

Hyponormal [MRR04]. Hypothesis

[vdH10].

ideal [BSZ10]. ideals [CHMN15, JG14].

Idempotence [XTC10].

Idempotence-preserving [XTC10].

idempotency [Fan11]. idempotent [FV08]. idempotents [AH10, JHH^{+}03]. identical [ZB10]. identifying [CGK^{+}13]. identity [VY12]. II [BNSY14, FN02]. Image [MDKM14]. Image-kernel [MDKM14]. imaginary [MX08]. immanant [CF14a]. implications [BFH05]. Improved [AG14, Dic10]. Incidence [WZ12]. inclusion [Oep03]. inconsistent [SK00]. Indefinite [Jar13, BdP13, BLM^{+}02, FGVR14, LTU96, Mat08, MRR04, Meh06b, Meh06a, MX99, NBdP14]. independence [DGH^{+}09, DD10, EG10, MÇ12]. index [BGRV09, DN12, DZ11, GS03, Loi12, Nik14, NY14, SWK14, Sha12, SK00, WW01, dAN13, dFNP13]. index-2 [Loi12]. indices [BCRS04, DDM09, SGL12]. induced [Rak12]. Inequalities [KN14, PTZ14, Sin14, TH10, CZY02, DL12, Dra11, HS03, LWZ04, MT10, MP00, NX06, Nie10, Tia12]. inequality [AT06, BLM^{+}13, DS07, KM13, LM98, Lin13b, Lin14, MS04, MNS11, Nie13, YF11]. Inertia [BJL^{+}10, BGK^{+}11, BGH15, Tia10b]. Inertially [CV07, GS11, CGK^{+}13, DOvdD10]. inertias [GOvdD13, Tia10a]. Infinite [BE97, BF05, Gil07b, Hol00, MN01]. inner [BLM^{+}02, FGVR14, LTU96, MRR04, Meh06b, Meh06a, MX99, RD12]. Institute [MST09]. integer [FFP12, MMT11, Zho12]. integers [Dub08]. Integral [GKW02, RS04]. Interest [SA12]. Interior [DS08, Dic10, JR08]. Interlacing [But07, Lot07]. International [BBCP12, HZ10b]. Interpolation [Pin04, AB00, Ber03, KDTP12]. Interval [RMZ14, SPS12a, CL04, Roh09c, Roh10, Roh11, RF11, WZ12]. Invariance [XQ11]. Invariant [Rod14b, CM05, FGVR14, MRR04, XD13]. invariants [GG04, MM12]. Inverse [RF11, Aur13, BF04, BGK^{+}11, BLM^{+}13, BNSY14, BI02, Bri07, BSZW12, CEM12, CGMSR11, CG13, COvdD09, CvdD12, CLW12, CPU10, CIW09, CILW11, DW10, Fer09, FR04, GKW02, Hog05, HM12b, KKP08, KP08, KP11, KS08a, LŠ08, LS13, LLYL11, LWB11, MD11a, Nab00, NX05, PES09, Rea02, Roh11, SK00, Son13, Stu98, Tia10b, TCID11, WW11, WW01, XCL11, ZB10, ZJW11, ZZL11]. inverse-orthogonal [LS13]. inverse-positive [FR04]. inverses [BZ03, Bap07, BPS05, BZZ09, BZ10, CL09a, COvdD08, CI12, LY11, MDKM14, NX06, Nie13, PS13, Tia05, TL07, Tia12, XQ11, XZ11, ZX08]. inversion [BOvdD04, NS12]. invertible [GLR03, LWB11, RS05]. involution [Che12, LD11, MD09]. involutory [D̄SZ03]. involving [MP00, Sin14, XQ11]. irreducibility [GS13, yZLXJ09]. Irreducible [FN06, Oep03]. irregular [CG07]. Isometric [RW02]. isometries [Che12, MN10, MD09]. isoperimetric [Siv05]. Isotropic [Rod08b]. issue [BBCP12]. iterates [Mat08]. Iterations [Per02]. Iterative [PL12, Pul12, SK00].

Jacobi [CEG12, CEM12, Gus13, Mat08]. Jensen [KM13, Nie13]. Johnson [KS08b, LM98]. joint [NBdP14, PT00, Xu10]. Jordan [AH10, Fan11, JTU98, KLL^{+}11, LS09, MX99, Tao11, XZ10]. journal [HZ10b].

Kantorovich [Nie10]. Kemeny [Kir14]. kernel [MDKM14]. Khatri [CZY02]. kinds [fYwW12]. known [MMT11]. Krein [BdPNdGPS08, LR98a, NBdP11]. Kreĭn [RS04]. Kronecker [Har14]. labeled [JN09]. Langer [RS04]. Laplacian [AKS12, BF11, BHL09, Bou09, BSZW12, But07, DH13, HP12c, KM14, Kir06, Kir07, LLZ13, LLW11, Liu12, LWYJ13, Lot07, MC13, MK10, OV10, Pat01, SH12, Tlu07, XD13, YL11, Yu10, YGX13, YFW14, ZZ11, Zhu11]. Laplacian-energy-like [XD13]. large [KOvdD00]. Largest [BHL09, AKS12, BF11, LLS14, LL08b, LhLY09, MS14, MKR13]. law [BLZ10, MD11a, ZX08]. laws [BL14b, CI12, LY11, TL07, XZ11]. LCM [Ova04, WHT14]. leading [MMT11, PP08]. Least [ZCL10, CJ11, HL07, LLS14, LWYJ13, PL12, WSH12, YGX13, YFW14, fYwW12, Zhu12]. least-squares [CJ11, HL07]. Lebesgue [Mol09]. Left [MVPS09, Cha12, KRT08, KS09, KS08a, XZ10]. length [BS07]. Leonard [AC10, Bro13, NT07]. letter [HJS02]. level [Pul12]. levels [RRCM11]. Lewis [Kar06, Tam99]. Li [BT12]. Lie [BSZ10, BDL^{+}09, BS07, NR11, PLBG12, Qi13, SZ09]. like [MT11, NX06, XD13]. Limit [Kir06]. line [Tlu07]. Linear [CKLN14, Dub08, Fan11, KRT08, Nik06, AM12, AG12, BE12, BP07, BCRS04, Brü09, CJ11, Dah03, DD10, FG14, FFP12, FS98, HM12a, HR06, HSSM05, HZ10b, JvdB12, KKP08, KS09, Kol00, KM05, KM11, KLR11, LS98, LMT10, LWB11, Loi12, MR12, MNS11, NBdP11, NS12, Nie06, Nie10, Nie13, RR02, Roh09b, Sch11b, SK00, TCID11, WW01]. Linearization [Lan08]. Linearizations [AV06, DDM09, BF12, DD08]. list [BL14b]. Liu [Wan08]. local [Qi13]. localization [AT06, GHK14, LL11, Maz10]. Loewner [PS11]. Loewy [LM98]. log [KM14]. log-concavity [KM14]. London [LM98]. long [NY14]. look [SS08c]. loop [YY12]. loop-free [YY12]. Loops [BEH^{+}14, SGL12]. loss [GT05]. low [FMMN11, GR09]. Lower [Sha12, BK97, BF11, LLYL11, ZHS08, ZL11, dF10, DD08]. Lyapunov [AS06, HM98, LX14, MS04, ZL10].

magnitude [FFP12]. main [DH13]. Majorization [MC13, AM12, AG12, BE12, DL12, DS13, KRT08, KS09]. majorizations [HR06]. many [D̄T01, Elz03, Kir01]. map [MR12]. Maps [LPR07, Mol09, Mol11, AH10, CJ08, Fan11, GLR03, LX11, MNS11, OK09, Per02, RR10, XTC10, XY14]. Markov [BL14b, GKMS13, Kir03, Kir04a, Kir14, Xu08]. marriage [Bru02]. matching [XD13]. matchings [Fri12, Zha14]. Mathematics [MST09]. Matrices [Fie05, Han09, JCST09, SPS12a, AT06, AG14, AJN14, AB00, ÁAFG12, AV04, AV06, AM14, Ari02, BT13, BB07, BNS14, BZ03, Bap07, BO00, BF04, BFH^{+}09, BP12a, BMN^{+}13, BTM07, Bea99, Ben10, BL14a, BF07, BE97, BLR01, BLM^{+}02, Bos10, BG99, Böt06, Böt11, Bou11, BCR13, BOvdD05, BPS05, BPS06, BH04, BZZ09, BZ10, BSZW12, BF08, BF10, BS05, CVV11, CRU10, CZY02, CL09a, CL09b, CEG12, CEM12, CG13, COvdD08, CvdD12, CS09, CN98, CN10, tCN12, CKLN14, CM05, CP99, DGH^{+}09, DS10, Dod08, DMS11, DK08, DMLT03, Dru06, Dru09, EPM13, ES08, ERS09, ES10, Ero04, FvdD00, FT02, FW10, FMS12, FN06, Fot10, FGVR14, FHR04, FV08, FMMN11, FKS08, FR04, FS98, FKS14, GSF09, Gil05, Gil07b, Gil10, Gil11, Gil14, GR09, GSM14]. matrices [GK99, GJOvdD07, GHK14, GLR03, HP12a, Han05, HP12b, HR06, HM98, HS03, HK05, HP07b, Hog09, HW12, HPP09, HZ10a, JM13, Jar13, JR08, JTU98, KKP08, KV04, KV14, KS11, Kir04b, Kir09, KM01, Kou06, KL13, LM98, Laf04, LS13, LMS04, LMT10, LZ12, LLZ13, LHL14, Lim12, Lin14, LvdD14, LH10, LWB11, LZFR12, LR98b, MVPS09, MMT03, MM13, MNP03, Mat08, MMW12, MMX00, MRR04, Meh06b, Meh06a, MX99, MX08, VV09, MT11, MSS08, MP09a, MMT11, MP00, MT04, Mor03, Nab00, NM02, NX06, Oep03, OTvdD09, OK09, Ova04, dSP14, Per07, PES09, PES11, Pin04, Pin08, RD12, RSS13, Rea02, RHC06, RMZ14, Rod08a, Rod08b, RŠ11, Rod14a, Rod14b, Roh09c, Roh10, Sem06, SW11a, Sch12, SM04, SH10, SC08, Stu98, SPS12b, Sze07, Tam10, Tia05]. matrices [Tia10a, TH10, Tia12, Tya12, WHT14, XLJ13, XY14, YF11, ZzH06, ZLXX09, yZLXJ09, ZLHL10, ZLLX12, ZZLL12, Zho12, ZH07, ZZL11, dF10, vdDW96, BL04]. Matricial [SS08a, GLR03]. Matrix [BBCP12, BOvdD04, ES10, FFB12, Gre14, LWZ04, MST09, SSV04, Aur13, BLP09, BS12, Bay11, BLZ10, BH13, Bri07, BCG96, Bru02, BF12, BT12, BSS14, CW13, CT04, CEM12, COvdD09, CLW12, CJLP99, CDH^{+}02, CDH^{+}03, Cof08, Dav04, DHHW03, DHS09, DDM09, DKS14, D̄SZ03, ERS09, EG10, FJS00, FN02, FHJT96, Gil07a, GP03, Gus13, HD10, HKT14, Hol00, HK98, HSSM05, HZS06, IT12, JW01, JK96, JS02, KKK11, KKP08, KP08, KP11, KRT08, KS09, Kir07, KL13, Lan08, LT12, LR02, LM10, LX11, LLYL11, LD11, LLMZ12, LY11, LMR11, LX14, MMMM11, MS14, MH96, MR12, Maz10, NS12, Nof12, PTZ14, PL12, PP08, Pat01, PS11, PM12, Pol10, PT00, Psa02, RV09, RR02, RMZ14, RS05, Rod07, Roh11, RF11, Roh13, SMB03, SA12, Shu11, SST04]. matrix [TMR06, TS01, TL07, Tia10b, Uhl98, WZ08, WZY08, WJ10, Woź01, WZ12, XHZ09, fXyHZ09, XZ11, XTC10, XZ10, YH05, YLD10, fYwW12, ZL10, ZZ14, ZB10, ZCL10, ZX08, ZCB03, dF05, DD08, HZ10b, KSvdD05]. matroid [vdH09]. MatTriad [BBCP12]. Max [GKMS13, BS05, BSS14]. max-algebraic [BSS14]. Maxima [DN12, Nik14, NY14, dAN13, dFNP13]. Maximal [Bru02, BSZ10, DKO12, GSM14, Gon12, Sne03, Tia10a]. maximum [BBC^{+}14, BCH^{+}13, CCH^{+}12, EEH^{+}12, HP12a, Hog11, JS02, MNP03, OV10, Pet12, SI10, YY12, Zhu12, vdH09]. mean [SS08b]. means [LWZ04, Mol11, Pol10, Sin14]. measurement [CM00]. Mercer [KM13]. Merris [GS03]. metabelian [BDL^{+}09]. Metcalf [MNS11]. method [BH13, CPU10, Jin05, LA02, PL12, SK00]. methods [LZFR12, Mat08, Pul12]. metric [BdP13, GR09, KKP09]. midpoint [Roh11]. Minimal [BLR01, SM01, BvdHL04, BvdHL05, CKSV05, DDM09, DK06, GLX14, KL15, LL08a, LZ12, RR10, Tia10a, fXyHZ09]. minimization [LX14]. minimizing [YFW14]. Minimum [AAC^{+}13, BFMN11, BBC^{+}09, BCH^{+}13, BEH^{+}14, CDH^{+}07, DGK^{+}12, GHH^{+}14, LNP14, BFH05, BFH^{+}09, BGK^{+}11, BMN^{+}13, CW13, CJ11, DGH^{+}09, DMS11, Gro10, Hog11, LLYL11, hLSxS12, MNR10, SGL11, SH08, TH10]. minimum-norm [CJ11]. minor [YSH12]. minors [BP12b, FvdD00, GJOvdD07, MMT11, Pin08]. minus [Rak12]. Mirsky [CL09b]. mixed [TL07, XZ11]. mixed-type [TL07, XZ11]. model [TT07, Xu08, ZJW11]. modèles [Ben02]. models [Ben02, CMS12]. modified [CG13]. modularity [MS14]. modules [SW11b]. moment [LZZ13]. monotone [CP99]. monotonically [JN09]. Monotonicity [SPS12a, SPS12b]. Moore [BI02, Bri07, HM12b, KP08, Mat03, MD11a, Tia10b, TCID11, WW11, XCL11]. Motzkin [Dah07]. moving [GT05]. MR2390359 [LhLY09]. Multi [SST04, Pul12]. Multi-agent [SST04]. multi-level [Pul12]. multidiagonal [FMS12]. multigraphs [MNR10]. multiple [LY11, XZ11, ZX08]. Multiplicative [GLR03, LMR11, BP07]. multiplicities [KV04]. multiplicity [JS02, Tlu07]. multiplier [NR11]. multipoint [Loi12]. multivariate [gLgWhX13]. Mumford [Sur00].

Natural [D̄T01]. necessarily [BFH^{+}09]. necessary [Cio10, Roh09c]. negative [CRU10, FvdD00, Jar13]. Nested [AKS12, CL04, FRM12, GJOvdD07, WZ12]. nests [Bru02]. networks [GT05]. Neutral [RŠ11, Rod14b]. Nevanlinna [Ber03, RS04]. Newton [BH13, NX06]. Newton-like [NX06]. Nilpotent [Per07, BSZ10, BK97, BS07, BKNO07, CVV11, KOS^{+}09, VV09, NR11, SZ09]. Nodal [BLS05, Tlu07]. Non [LR13, LR02, LMS04, BdPdP10, Gil11, GS13, Laf04, LL08b, LhLY09, LHY09, LWYJ13, MR12, MMW12, MT10, Pul12]. non-autonomous [MR12]. non-bipartite [LWYJ13]. non-commuting [Gil11, MT10]. non-equilibrium [MMW12]. Non-existence [LMS04]. non-Hermitian [BdPdP10]. non-homogeneity [GS13]. non-real [Laf04]. non-regular [LL08b, LhLY09, LHY09]. Non-singular [LR13]. non-symmetric [Pul12]. Non-trivial [LR02]. nonincreasing [OK09]. Nonnegative [SC08, SSM10, AG14, BK97, BSS14, CDH^{+}03, CP99, DMLT03, ES10, FN98, FV08, HS03, HP07b, HZ10a, KS08a, KM01, LM98, Laf04, LŠ08, LHL14, NM02, Oep03, OTvdD09, SRMB07, SRM11, WZ08, YH05, YF11, Zho12, BL04, MST09]. Nonnegativity [FV08, CEH^{+}12, EHT09, Hog09]. nonpositive [CRU10]. nonsingular [BPS05, Kal12, LMS04, PP08, WHT14]. nonsingularity [BO00, FFJM14]. nonstrictly [ZLXX09, ZZLL12]. nonsymmetric [SK00]. nonzero [CV07, DOvdD10, VV09, ZZ14]. Norm [Gil05, Gil11, KLL^{+}11, CF14b, CJ11, LPR07]. Normal [PP08, Ari02, Che12, KS11, Meh06b, Meh06a]. normalized [But07, Kir06, Kir07]. norms [BGJ13, Böt06, Dru09]. Note [EEH^{+}12, Lot07, Sha13, BCR13, BSZW12, CW13, Car12, Dah03, HP12b, Hog11, KV14, Kir07, KM01, LS13, LL08b, LhLY09, LY11, MS14, NBdP14, NX06, Pek14, RHC06, SM04, YH05, Zhu12]. Notes [Lin13b, WSH12]. notion [KKP09]. nullity [BBC^{+}14, BCH^{+}13, CCH^{+}12, CL07, EEH^{+}12, Hog11, LHC09, Pet12, vdH09]. nullstellensatz [CHMN15]. number [AAC^{+}13, BFMN11, BBC^{+}14, BCH^{+}13, CCH^{+}12, DZ11, EEH^{+}12, EG10, LR98b, MNP03, Pet12, Siv05, SH08, XD13, dF10, dAN13]. numbers [BG99, CJLP99, Gil14, Kar06, MÇ12, TG06, XLJ13, ZZ14]. Numerical [BdP13, LTU96, AMS09, AM14, BLdP00, BdPNdGPS08, CN98, CN09, CN10, CPU10, DLRM^{+}13, Hen10, KLR11, LR98a, LPR07, NBdP11, NBdP14, PT00, RSS13, Rod14a, VU12, ZLLX12].

objects [NWZ98]. oblique [KL15]. observability [FS98]. obtained [BF12]. occasion [HZ10b]. occur [God12]. octonion [Rod14a]. odd [BF12, DMS11, MNP03]. old [JW12b]. One [HS03, AJN14, Bou11, BOvdD04, CJ11, Jar13, Kal12, Lim12, OK09, SH10]. One-sided [HS03]. only [Elz03]. open [Fri12]. opérateurs [Ben02]. operation [Pek14]. Operator [Dru09, CF14a, CF14b, CN09, Dra11, DFS13, KM13, KN14, LTU96, Sin14, SW11b, XQ11]. operators [BdPdP10, BdP13, Ben02, Bra09, BKNO07, CIW09, Dra11, Fan11, FN98, Gil07b, HR06, KM11, KLR11, LR98a, MT10, Mol09, MN10, Mol11, NBdP11, Nie06, Nie13, Sch11a, VU12]. Opinion [WS12]. Optimal [JPS12, Rol03, ÁAFG12, DGH^{+}09, KM11, NWZ98, fXyHZ09, XCL11, ZCL10, ZCB03]. orbits [D̄T01, DD11b]. order [AMS09, AHL^{+}10, BLZ10, BH04, CV07, CI12, EPM13, Fen01, GS11, HP12a, LY11, MD11a, Rak12, Rod06, Sch11b, TL07, XZ11, ZX08]. ordered [BL14b, Nie10]. Ordering [LZ12, PS11]. orderings [BLZ10]. orders [WZ12]. oriented [GLX14, Ref14]. ortho [XHZ09]. ortho-symmetric [XHZ09]. orthogonal [CMS12, CJLP99, D̄SS09, LS13, ZCL10]. orthogonality [SS08c]. Orthosymmetric [Sin12]. Ostrowski [LHL14]. outerplanar [BFMN11]. Overton [Kar06].

packings [Per02]. pair

[CFP11, DKS14, NT07, WZS09, dF05]. Pairs [Bou11, AC05, AC10, Bro13, JW12a, Rod08b, RŠ11]. Palindromic [BF12, MMMM11, Rod06]. Palo [MST09]. parabolic [AS06, Sze07]. paracontracting [BE97]. parallel [KDTP12, NX05]. parameter [MP09b]. parameterized [ZHS08]. parameters [BFH05, FM13]. Pareto [SVP11]. part [IT12]. Partial [MD09, Uhl98, AJN14, BLZ10, Che12, EPM13, EEH^{+}12, JTU98, Rak12]. partially [Dod08]. partitioned [SRM11]. Pascal [Bay11, MP09a]. Path [ZZL11, COvdD08, CCH^{+}12, dF03]. Paths [ERS09, EFP12, NY14]. pattern [CW13, GSF09, JdGM05]. Patterns [JdGM05, AHL^{+}10, BCD^{+}09, CHOvdD09, CEH^{+}12, CKSV05, CV07, CF12, CGK^{+}13, DOvdD10, EHT09, GS11, GOvdD13, Hog11, JW12a, KS08a, KOS^{+}09, VV09, Per07, Tya12]. PDEs [ZLLX12]. abipartite [Bro13]. combined [CFP11]. Itô [Bro13]. skew-Hermitian [IT12]. skewsymmetric [RŠ11]. pencil [KKP08]. Pencils [MOS^{+}98, BF12, Cof08, DKS14, MPT00, MMX00, Rod07]. pendant [BF05, GL10, Liu12, LWYJ13]. pendent [LHC09]. Penrose [BI02, Bri07, HM12b, KP08, Mat03, MD11a, Tia10b, TCID11, WW11, XCL11]. per-symmetric [CKLN14]. perfect [BF07, God12, Zha14]. Periodic [BCG96]. peripheral [FN98]. permanent [Wan08]. permanental [BLdP00]. Permanents [DS10, LLZ13]. perplectic [MMD05]. Perron [BT12, ES08, ERS09, NX05, Per02, RHC06]. Perturbation [DW10, DX12, MX08, CGMSR11, CS09, FFJM14, HP12c, LA02, XCL11]. Perturbations [Gil10, CMS00, FKS14, GR09, MX08]. Perturbing [Laf04]. physics [MMW12]. Pick [Ber03]. placed [JK11]. placement [CH13]. plane [De 13]. point [Fie98, HKT14]. points [tCN12, DS08, Kir06]. Polar [CMS00, HPP09, XLJ13, vdDW96]. polyhedral [Tlu07]. Polynomial [AMS09, MT10, AV04, AV06, BF12, Fen01, FS98, KDTP12, Nof12, Stu98, VY12, WWY10, dF03, dF05]. polynomially [Meh06a]. polynomials [AM14, BH13, Cha12, tCN12, DDL06, DDM09, FN02, Han09, KM14, Lan08, MMMM11, MPT00, Maz10, PP08, PT00, SMB03, Spi06, Tya12, DD08]. polytope [Mar06]. Positive [CN98, FHR04, HJS02, LvdD14, Pet12, SW11b, YLD10, BGRV09, Bal11, BCRS04, CW13, CS09, Dic10, Dru09, DS08, EEH^{+}12, El 11, FJS00, FGVR14, FR04, Gow12, GS13, Han09, HZ10a, Jar13, JCST09, JN09, JK11, KS08a, Lin14, MM13, Mol09, MN10, Mol11, Mor03, MNS11, Pin08, Sch11a, SH10, vdH09]. positive-definite [Dru09]. positives [El 11]. positivity [BCD^{+}09, CHOvdD09, EHT09, FT02, FFJM14]. Possible [ZZ14, CJLP99]. PostLie [PLBG12]. potent [BT13, IT12, LT12]. potentially [KOS^{+}09]. pour [Ben02]. power [BL14b, CF14a, CF14b, CHOvdD09, Gol08, GL11, HM05, HP12b, WHT14]. power-associative [Gol08]. power-positivity [CHOvdD09]. powers [BGJ13, DGK^{+}12, Dub08, Shu11, Uhl98]. Preface [BBCP12, BL04, FFB12, HB98, Her00, HZ10b, KSvdD05, MST09]. preliminary [BFH^{+}09]. prescribe [LPW14]. prescribed [Dod08, FS98, RRCM11, SC08, Tam10]. preserve [GLR03]. preservers [AM12, AG12, BE12, CKLN14, DK08, KRT08, KS09, KLL^{+}11, LMT10, Lim12]. preserving [BK97, CT04, ES10, Fan11, HR06, LPR07, LZFR12, MMD05, MR12, Mol09, Mol11, Sze05, XTC10]. primitive [BF08, BF10, SGL12, YY12]. Principal [CG07, GJOvdD07, MMT11, Rol03]. principle [FR04]. probabilities [Xu08]. problem [BF04, BNSY14, CEM12, CDH^{+}02, CDH^{+}03, CPU10, DHS09, FJS00, Gro10, Hog05, JK96, LŠ08, gLgWhX13, LX14, MS04, MP09b, PL12, RR10, Rea02, Roh09b, ZJW11]. Problems [BDL^{+}09, BP07, BGK^{+}11, BLM^{+}13, BLR01, CEG12, CF12, DHHW03, Fer09, Fri12, Loi12, Pul12]. product [BS12, BLM^{+}02, CZY02, CLW12, CMS00, FGVR14, HK98, HZ10a, KV14, LTU96, LLYL11, Lim12, Meh06b, Meh06a, MP00, RD12, TL07, TG06, XQ11, YF11, ZZL11]. Products [BKNO07, Gre14, GJOvdD07, BE97, D̄T01, Fan11, Hol00, KKK11, KV04, KSH11, KL13, KLL^{+}11, LY11, MRR04, MX99, PTZ14, Sch11a, XZ11, ZX08]. projection [KL15]. projections [Cof08, Sch12, Spi06]. projective [De 10]. projectors [TCID11]. Proof [D̄ok02, Ari02, Ero04, Sur00]. proofs [CL09b]. propagation [Pul12]. propeller [LZ14]. proper [SPS12b]. Properties [LT12, ZCB03, BE12, CF12, Dah07, DN12, DS13, EFP12, GLR03, HK05, HW12, JW01, Kir03, KL14, LLMZ12, MH96, Ref14, Tao11, TT07, XQ11]. Property [Bal11, BP07, BNPS08, ES08, ERS09, FG14, MPT00, Mec12, SST04]. pseudo [Fot10, HP07a]. pseudo-Hermitian [Fot10]. Pseudomonotonicity [Tao11]. pseudospectral [KL14]. pseudospectrum [AGV11]. purely [MX08].

Quadratic

[Mat08, YFW14, BdP13, HD10, Jin05, VY12]. quadratric [El 11]. quantum [AC05, BLdP00]. quartics [De 13]. quasi [DK08, GL10]. quasi-commutativity [DK08]. quasi-tree [GL10]. quasi-unicyclic [GL10]. Quasihyponormal [RD12]. quaternion [Son13, WZ08, WZY08, WJ10, fYwW12]. quaternionic [CD00, DDL06, Rod07, Rod08a]. queueing [GT05]. quotient [SZ09]. Quotients [Har14].

radii [GL10, Liu12, LL12, YSH12, YWZ12]. radius [CN09, HZ10a, KV14, KL14, LPR07, LHL14, Rum02, Sch11a, Sne03, SH08, SI10, Xu10, YH05, Yu10, ZzH06, ZZ11, Zha14, Zhu11]. Radon [Lin13a]. random [FG14]. range [AM14, CN98, CN10, CPU10, DLRM^{+}13, Hen10, NBdP11, NBdP14, PT00, RSS13]. Ranges [RR10, BLdP00, BdPNdGPS08, BdP13, KLR11, LTU96, LR98a, Rod14a, VU12]. Rank [BEH^{+}14, Tia10b, AM14, BFH05, BFH^{+}09, BFMN11, BvdHL04, BvdHL05, BBC^{+}09, BGK^{+}11, BMN^{+}13, Bea99, BCH^{+}13, Bos10, CW13, CRU09, CJ11, CDH^{+}07, CKLN14, DGH^{+}09, DGK^{+}12, DK06, DMS11, FMMN11, GR09, GHH^{+}14, Gro10, HP12a, Hog11, hLSxS12, LNP14, Lim12, LX14, MNR10, OK09, RR10, SSV04, SM01, SM04, fXyHZ09, XY14, ZZ14]. rank- [AM14, XY14, Bea99]. rank-one [CJ11, OK09]. rank-two [CKLN14]. Ranks [LPW14, Tia10a, WJ10, Zho12]. Rao [CZY02]. rate [LA02]. Rational [D̄SS09, MM12, GP03, Nof12]. ray [LMS04]. re [SW11b, WZ08]. re-nonnegative [WZ08]. re-positive [SW11b]. Reachability [BCRS04, FW10, BGRV09]. Real [CHMN15, Cof08, FGVR14, BdPdP10, D̄T01, D̄SS09, Laf04, LZFR12, Meh06a, PES11, Rod07, RŠ11, Rum02, XLJ13, dF10]. realizable [KM01]. realization [GP03, PES11, SRMB07, SSM10, SRM11].

Realizations [SZ09, JW12a]. reciprocal [BNPS08]. Recognition [Mor03]. reconstruction [Mor11, Mor13]. recovery [DDM09]. rectangular [KKP08]. recurrence [Lin13a, MR12]. recursive [MSS08, PS13]. reduced [HP12a, LPW14]. Reducing [FHJT96, MT04]. Refined [DOvdD10, GOvdD13]. refinement [LM98]. Refinements [KM13]. reflexive [WZ08]. reflexivity [Bra09]. reformulation [HM12a]. regarding [BF04, JTU98, dF05]. Regions [HZS06]. regular [Brü09, BF08, CEG12, Car12, De 10, Elz03, Fri12, Lan08, LL08b, LhLY09, LHY09, Maz10, Siv05]. regularity [Roh09c, Roh10]. related [AS06, AC05, De 10, FKS08, MSS08, Tao11]. relation [Lin13a, PT00, Tlu07]. relations [GG04]. Relationship [HM12b]. relative [JCST09, KN14]. relaxation [LX14]. remark [De 13]. repetition [BF12]. representation [GKW02, LLMZ12, RS04]. Representations [CvdD12, CIW09, BZ10, BSZW12, COvdD09, WW11]. representative [KKP09]. require [CHOvdD09, EHT09, KS08a]. respect [Gon12, MX99, WWY10]. restricted [KP11]. restriction [NR11]. restrictions [BMN^{+}13]. result [Tam99, Tam10]. Results [Fri12, BNS14, BLZ10, BZZ09, CILW11, FFJM14, LLS14, LL12, Mar08, WHT14, WW01]. RETRACTED [Gus13]. Reverse [MD11a, BLZ10, CI12, LY11, TL07, XZ11, ZX08]. reverse-order [TL07]. revisited [DS10, Dru06, RSS13]. Riccati [Böt11, LZL10, ZL11]. Riemann [YY13]. ring [CG13, Rak12]. rings [AH10, Che12, HP07a, MD09, MDKM14, SS08a, XZ10]. Rodman [JTU98]. role [KM11]. rooted [CMRR13, RRCM11]. roots [LZFR12, Psa02, RS05, Rod08a]. rotations [BTM07, Sin12]. Roth [HP07a]. row [BH04, DMLT03, Mor03, Tia05]. rule [JvdB12].

Sakamoto [Mar08]. sandwich [HS03]. scalar [CMS00, HM98, TG06]. scaled [Mat08]. Scaling [EFP12, BS05]. Schatten [Böt06]. Schur [BOvdD05, FV08, JG14, LH10, LZFR12, NR11, TT05, YF11, ZLXX09, yZLXJ09, ZZLL12]. Schwarz [Tya12]. second [AKS12, BF11, Fen01, Sch11b, SZ09, Tlu07]. second-order [Fen01]. Segrè [Cof08]. self [FKS08, GS13]. self-duality [GS13]. selfadjoint [Dra11]. semi [El 11, vdH09]. semi-définies [El 11]. semi-definite [vdH09]. Semidefinite [Hen10, BP07, Bal11, BFMN11, CW13, EEH^{+}12, El 11, HZ10a, JG14, Lin14, LvdD14, LX14, MRR04, Pet12]. semigroups [Bra09, LMR11]. semiiterative [LA02]. Semitransitive [Sem06]. Sensitivity [gLgWhX13]. Separable [MPT00, KDTP12]. separator [YL11]. sequence [MSS08, ZZ11]. sequences [BHL09, GJOvdD07, KL15, MMT11, Sol02]. serial [Mat08]. series [Dod07]. servers [GT05]. set [BE12, BF10, BGH15, Gon12, KM01, ZZ11]. sets [BJL^{+}10, Car12, DD10, ES10, Fie98]. Shadowing [FG14]. Shalom [JTU98]. shape [LHL11]. Shapes [LR98a]. Sharp [DD08]. Shells [BLM^{+}02]. shift [CN09, VU12]. shifted [Aur13]. Short [CL09b]. shorted [Pek14]. sided [AB00, HS03]. Sign [BCD^{+}09, CHOvdD09, CEH^{+}12, EHT09, KS08a, LS98, Tya12, AHL^{+}10, CW13, DHHW03, GSF09, GS11, GOvdD13, HK05, HP07b, Hog11, KOS^{+}09, Per07, Rum02, Uhl98]. sign-complex [Rum02]. Sign-consistency [LS98]. sign-real [Rum02]. signed [YY12]. signless [AKS12, DH13, HP12c, LLW11, Liu12, LWYJ13, MC13, MK10, OV10, YL11, YGX13, YFW14, Zhu11]. similar [HP07b]. similarities [Ver08]. similarity [HS03]. Simon [Bid07, FR04]. simple [Ari02, BCH^{+}13, Shu11]. Simplifications [LHL14]. Simultaneous [BL14a, BCR13, HS03]. single [CFP11, Kir04a]. single/combined [CFP11]. Singular [tCN12, DS07, GK99, CEM12, DDM09, JPS12, KM05, LR13, Loi12, MP09b, PS11, Sci07, SK00, Tam10, Tri13, WW01, ZHS08]. singularities [Böt06]. six [BJL^{+}10]. skew [BZZ09, FKS08, GLX14, KL13, LZFR12, MX99, Rod08a, SHZ12, Son13, WJ10, ZB10]. skew- [KL13, WJ10]. skew-Hamiltonian [LZFR12]. skew-self-adjoint [FKS08]. small [BLM^{+}13, BNSY14, Sta13, SZ09, Xu08]. small-world [Xu08]. smallest [HP12c, Pat01, PS11, ZHS08]. Smith [MMMM11]. Solution [HSSM05, CJ11, DD11a, DKS14, Fen01, SK00, WZ08, fXyHZ09, YLD10, ZL10, ZL11, ZLLX12]. solutions [Brü09, FFP12, LR02, LM10, Roh09b, SW11b, TMR06, WZY08, WJ10, WW01, XHZ09, fYwW12, ZCL10]. solvability [LS98, LD11]. Solvable [BSZ10, WZS09]. solvents [HD10]. solving [Roh09a]. Some [BE12, BLZ10, BZZ09, CRU10, CZY02, CILW11, De 10, DL12, LLYL11, LLS14, LL12, MH96, Mar06, TT07, Bal11, De 13, GR09, JW01, KL14, LM98, MMT11, RRCM11, SW11b, YGX13, dF10]. Soules [SM04]. space [CIW09, De 10, EPM13, FGVR14, LTU96, LR98a, MM12, NBdP11]. Spaces [Bea99, Bos10, DX12, BdPNdGPS08, BdP13, BLM^{+}02, CT04, CKLN14, D̄T01, Dra11, Fie98, GR09, LR98b, Meh06b, Meh06a, Nie10, RD12, Sch11a, TG06, XTC10]. spanning [VY12]. Special [Tia05, BGRV09, BBCP12, De 13, D̄ok02, fYwW12, HZ10b]. Spectra [FS00, RM09, RRCM11, CS09, FN02, Gus13, KS08b, LLW11, OV10, PES11, SVP11, SRMB07, SSM10, SRM11]. Spectral [Ber03, HK05, Hog05, LLMZ12, LHL11, LZ14, Ref14, AS06, BH13, CF12, GL10, GP03, HZ10a, KV14, LPR07, LZZ13, LHL14, Liu12, LL12, Rum02, Sch11a, Sne03, Sta13, SH08, SI10, Stu98, Xu10, YH05, Yu10, YSH12, YWZ12, ZzH06, ZZ11, Zha14, Zhu11]. Spectrally [AHL^{+}10, GSF09, CKSV05, DOvdD10, Per07]. Spectrum [BK97, Gil07b, Maz10, AT06, Bou09, FN98, HS10, Lot07, MK10, SH12]. splittings [CP99, JW01, SPS12b, WW01, Woź01]. spreads [De 10]. square [Ben10, D̄SZ03, HP12a, KKP08, LZFR12, Rod08a]. squares [CJ11, HL07, PL12, fYwW12, ZCL10]. stability [HM98, HM12b]. stabilization [FS98]. Stable [Gol08, BCR13, FHR04, Lan99]. stably [Rod06]. stair [FRM12]. standard [Sze07]. Starlike [OV10]. state [God12]. states [LPW14]. station [GT05]. stationary [Kir03, Kir04a, Kir14, Sch12]. statistical [MMW12, TT07]. stochastic [Kir04b, Kir09, LZ12, Rea02, Sch12]. strict [BLM^{+}13, Rod07]. strictly [BPS06]. Strong [AG12, BL14b, vdH10, BNPS08, CEH^{+}12, ERS09, GN02, Hog09, KSH11, Roh10]. Strongly [Elz03, Lan99, Car12, HR06, RD12, Siv05]. structural [ZJW11]. Structure [CVV11, KS09, LZFR12, MMD05, Sze05, CMS12, CFP11, HR06, NM02]. Structure-preserving [LZFR12]. Structured [Dav04, KL13, MMT03, MX99, SW11a, TG06, FKS08, MMX00, MX08]. structures [MM12, PLBG12]. stubborn [WS12]. study [BFH^{+}09, KKP08]. style [JPS12]. subblocks [ZB10]. subclasses [SPS12b]. Subdirect [BPS05, BPS06, ZH07]. subdivided [CCH^{+}12]. subdivision [BBC^{+}14]. subdivisions [BBC^{+}09]. Subdominant [Kir09, Kir04b]. subexponential [BL14b]. subgraphs [BLS05, RRCM11]. subgroups [Sze07]. subject [FK98]. Submatrices [BNS14, MNP03, Tia10b]. submatrix [ZJW11]. subnorms [Gol08]. subpolytopes [Mar06]. subspaces [Bru02, De 10, FGVR14, MRR04, MT04, Rod08b, RŠ11, Rod14b, Rol03, Sem06]. substantial [Tlu07]. sufficient [Cio10, Roh09c]. sum [BH04]. sums [BPS05, BPS06, DMLT03, Dub08, Fot10, Kir09, ZH07]. sun [MK10]. superstars [Fer09]. surfaces [Tlu07, YY13]. survey [Roh09c, RF11]. SVEP [Mec12]. Sylvester [DD11a, HKT14, RR10, WZS09]. Symmetric [DKS14, SRMB07, ZJW11, BF04, BFH^{+}09, CT04, CEM12, CF14a, CF14b, CJ11, CKLN14, DHHW03, HK05, JK96, JW12a, KM01, KL13, PES09, PES11, Pul12, RR02, Rea02, RŠ11, RMdAJ10, SW11a, SGL12, SH10, XHZ09, fXyHZ09, YY12, ZCL10, dF10]. symmetric/ skewsymmetric [RŠ11]. symmetrically [JK11]. symmetries [AV06, Rod07]. symmetry [DLRM^{+}13]. symplectic [Lin13a, MVPS09]. system [CJ11, RMZ14, TMR06]. systems [BGRV09, BLdP00, BCRS04, Brü09, CM00, Fen01, FFP12, FKS08, FS98, JvdB12, KKP08, Lan99, LS98, Mat03, Mor11, Mor13, NS12, SS08a, Sch11b, SK00, SW11b, WW01]. Szego [MR12].

Taylor [Lin13b]. Techniques [CGK^{+}13]. tensor [BGJ13, D̄T01, Lim12, PTZ14]. tensors [DD10, FMMN11]. term [ZZ14]. terms [JS02]. ternary [tCN12]. test [Hog09, Stu98]. tests [AG14]. th [CF14a, CF14b, Psa02]. their [AV06, BdPNdGPS08, BGJ13, Bou09, Bou11, BPS05, KM11, LZ12, LLZ13, LLW11, OV10, SH12]. Theorem [CRU09, Bru02, GKMS13, LL11, Oep03, SS08b, Tlu07]. theorems [BLS05, CL09b, CP99, HS03, LLZ13, Woź01]. theoretic [FK98]. Theory [FFB12, MST09, Hog05, Per02, KSvdD05]. there [Elz03]. third [Pat01]. Thompson [MN10]. three [DD10, D̄T01]. Tight [MNS00, RW02, Tlu07]. time [Brü09, Stu98].

Toeplitz [Ari02, BG99, Böt06, Böt11, FN06, LS13, MP09a]. tools [MMT03]. topological [BE12]. total [FFJM14]. Totally [Bro13, JN09, AG14, CRU10, FJS00, JCST09, JK11, MM13, Pin08, YF11]. tournament [DMS11, GK99]. Trace [DLRM^{+}13, Rea02, Xu10]. Traces [Gre14]. Tracial [KLR11]. trajectories [BL14b]. transfer [God12]. transform [FT02]. transformation [BP07, D̄SZ03, Nof12]. Transformations [Bal11, CT04, Gow12, NT07]. transition [Xu08]. transpose [D̄SZ03]. trapezoidal [Dra11]. tree [AHL^{+}10, BLP09, BS12, BFMN11, CDH^{+}07, FHJT96, GS11, GOvdD13, GL10, Gon12, GKMS13, JCST09, MNS00]. trees [BF05, BHL09, CL04, DGK^{+}12, EEH^{+}12, LL08a, LZ12, LHL11, LL12, MN01, OV10, RM09, SI10, ZZ11]. triangles [MP09a]. triangular [AB00, AH10, BK97, CJ08, CKLN14, Gil07b, JTU98, Mar14, Qi13, Shu11, XZ10]. tridiagonal [AC05, AG12, BO00, CN98, CN10, Kou06, MMW12, PL12, PES11]. triple [TL07]. triples [Bro13, Han05]. trivial [LR02]. Two [FR04, GT05, HP14, HK98, PES09, fYwW12, AB00, BF04, BvdHL04, BvdHL05, BGK^{+}11, BDL^{+}09, BL14a, Ber03, CKLN14, DD10, D̄ok02, Gil05, Gil11, Gus13, HJS02, HS10, LWB11, MM12, MN01, MP09b, RMZ14, Roh13, Spi06, YY13, ZB10, Ben02]. two-dimensional [MM12]. two-letter [HJS02]. two-matrix [Roh13]. two-parameter [MP09b]. two-sided [AB00]. Two-station [GT05]. type [AS06, Bro13, Dra11, HKT14, KS11, Lin13b, MNS11, Nie10, SK00, TL07, XZ11]. types [D̄ok02].

Ulam [HM12b]. Ultrametric [Fie98]. Unicyclic [BNPS08, GL10, LL12, LWYJ13, SH12, SGL11, Zha14]. unilateral [DDL06]. unions [Bou09]. unit [BF07, Roh11]. unitary [BLR01, MX99]. units [HP14]. universal [TS01]. Universally [DGH^{+}09]. unknown [RMZ14]. unspecified [AJN14, FJS00]. updates [CJ11]. updating [ZJW11]. Upper [FFP12, MKR13, Nie06, AB00, BGRV09, JTU98, Kir01, LHL14, LZL10, XZ10, Yu10]. Using [BGH15, HL07, But07, Kir07, SST04].

value [CEG12, DS07, JPS12, Loi12, PS11, Roh09a, SS08b, ZHS08]. valued [GG04, vdAtMM07]. values [Elz03, Gil07a, GK99, Tam10]. Vandermonde [BF07, MH96, MM13]. variables [Bay11]. variant [BFH05]. variants [BGH15]. Variational [Rum02]. varieties [Cof08]. Vector [Dra11, BH04, D̄T01, GG04, Kir14, LNP14, MM12, MNR10]. vector-valued [GG04]. vectors [Gon12]. Verdière [BFH05, FM13]. versus [Ber03, D̄SS09]. vertex [JS02, Lot07, VY12]. vertices [BJL^{+}10, Dru04, GL10, LHC09, Liu12, LWYJ13, WS12]. via [CN98, KDTP12, NO03, NX05, Ova04, PTZ14, Uhl98]. volume [HZ10b].

Waring [De 13]. Weak [SPS12a, SPS12b, Bid07, CP99, Lot07, Woź01]. weaker [Woź01]. weakly [DHHW03]. weighing [NWZ98]. weight [BF05, MN01, ZZ11]. Weighted [EG10, MD11b, BF11, But07, CMRR13, CN09, KP11, RM09, RRCM11, Son13, VU12, XCL11, ZZ11]. weights [CN09, VU12]. well [MMT11]. well-known [MMT11]. which [Bou11, Elz03, FGVR14, hLSxS12]. whose [BvdHL04, BvdHL05, MS14, MMT11]. Wiener [BH13]. wild [BDL^{+}09]. without [NY14, SGL12, YWZ12]. WLSEs [TT07]. words [Dru09, HJS02]. Workshop [BL04, HZ10b, KSvdD05]. world [Xu08]. ycle [dFNP13].

Zero [CCH^{+}12, VV09, Pin08, BBC^{+}14, BCH^{+}13, BGH15, DOvdD10, EEH^{+}12, GHH^{+}14, MS14, Pet12, Rea02]. Zero-nonzero [VV09, DOvdD10]. Zeros [DDL06, CJLP99]. zeta [BS12]. zone [SA12].

References

Ahmadi:2013:MND

[AAC^{+}13] Bahman Ahmadi, Fatemeh Alinaghipour, Michael S. Cavers, Shaun Fallat, Karen Meagher, and Shahla Nasserasr. Minimum number of distinct eigenvalues of graphs. Electronic Journal of Linear Algebra, 26(1):673–691, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/45.

Alvarez:2012:ECO

[ÁAFG12] Víctor Álvarez, José Andrés Armario, María Dolores Frau, and Félix Gudiel. Embedding cocyclic D-optimal designs in cocyclic Hadamard matrices. Electronic Journal of Linear Algebra, 24(1):66– 82, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/7.

Alpay:2000:TSI

[AB00] Daniel Alpay and Vladimir Bolotnikov. On two-sided interpolation for upper triangular matrices. Electronic Journal of Linear Algebra, 6(1):31–55, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol6/iss1/5.

Alnajjar:2005:FTP

[AC05] Hasan Alnajjar and Brian Curtin. A family of tridiagonal pairs related to the quantum affine algebra U_{q}(ŝl_{2}). Electronic Journal of Linear Algebra, 13(1):1–9, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/1.

Alnajjar:2010:LPE

[AC10] Hasan Alnajjar and Brian Curtin. Leonard pairs from the equitable basis of sl_{2}. Electronic Journal of Linear Algebra, 20(1):490–505, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/36.

Armandnejad:2012:SLP

[AG12] Ali Armandnejad and Zahra Gashool. Strong linear preservers of g-tridiagonal majorization on R^{N}. Electronic Journal of Linear Algebra, 23 (1):115–121, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/7.

Adm:2014:ITC

[AG14] Mohammad Adm and Jurgen Garloff. Improved tests and characterizations of totally nonnegative matrices. Electronic Journal of Linear Algebra, 27(1):588–610, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/255.

Armentia:2011:DDA

[AGV11] Gorka Armentia, Juan-Miguel Gracia, and Francisco E. Velasco. Derivatives of the diameter and the area of a connected component of the pseudospectrum. Electronic Journal of Linear Algebra, 22(1): 1004–1019, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/65.

An:2010:CJD

[AH10] Runling An and Jinchuan Hou. Characterizations of Jordan derivations on triangular rings: Additive maps Jordan derivable at idempotents. Electronic Journal of Linear Algebra, 21(1):28–42, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/5.

Arav:2010:SAT

[AHL^{+}10] Marina Arav, Frank Hall, Zhongshan Li, Krishna Kaphle, and Nilay Manzagol. Spectrally arbitrary tree sign patterns of order 4. Electronic Journal of Linear Algebra, 20 (1):180–197, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/14.

Akin:2014:CPM

[AJN14] V. Akin, C. R. Johnson, and S. Nasserasr. TP_{K} completion of partial matrices with one unspecified entry. Electronic Journal of Linear Algebra, 27(1):426–443, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/27.

Andelic:2012:NGB

[AKS12] Milica Andelić, Tamara Koledin, and Zoran Stanić. Nested graphs with bounded second largest (signless Laplacian) eigenvalue. Electronic Journal of Linear Algebra, 24 (1):181–201, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol24/iss1/14.

Armandnejad:2012:GML

[AM12] Ali Armandnejad and Asma Ilkhanizadeh Manesh. GUTmajorization and its linear preservers. Electronic Journal of Linear Algebra, 23 (1):646–654, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/46.

Aretaki:2014:RKN

[AM14] Aik. Aretaki and John Maroulas. On the rank-k numerical range of matrices polynomials. Electronic Journal of Linear Algebra, 27 (1):809–820, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/269.

Afshin:2009:PNH

[AMS09] Hamid Reza Afshin, Mohammad Ali Mehrjoofard, and Abbas Salemi. Polynomial numerical hulls of order 3. Electronic Journal of Linear Algebra, 18(1):253–263, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/22.

Arimoto:2002:SPC

[Ari02] Akio Arimoto. A simple proof of the classification of normal Toeplitz matrices. Electronic Journal of Linear Algebra, 9(1):108–111, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/9.

Ali:2006:LTE

[AS06] Mouhamad Al Sayed Ali and Miloud Sadkane. On a Lyapunov type equation related to parabolic spectral dichotomy. Electronic Journal of Linear Algebra, 15 (1):134–142, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/9.

Adam:2006:EIS

[AT06] Maria Adam and Michael J. Tsatsomeros. An eigenvalue inequality and spectrum localization for complex matrices. Electronic Journal of Linear Algebra, 15 (1):239–250, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/19.

Aurentz:2013:FIS

[Aur13] Jared L. Aurentz. A factorization of the inverse of the shifted companion matrix. Electronic Journal of Linear Algebra, 26(1):231–247, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol26/iss1/18.

Antoniou:2004:NFC

[AV04] E. N. Antoniou and S. Vologiannidis. A new family of companion forms of polynomial matrices. Electronic Journal of Linear Algebra, 11(1):78–87, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/8.

Antoniou:2006:LPM

[AV06] Efstathios N. Antoniou and Stavros Vologiannidis. Linearizations of polynomial matrices with symmetries and their applications. Electronic Journal of Linear Algebra, 15 (1):107–114, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/7.

Balaji:2011:CPS

[Bal11] R. Balaji. Characterization of P -property for some Ztransformations on positive semidefinite cone. Electronic Journal of Linear Algebra, 22 (1):1020–1030, 2011. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/66.

Bapat:2007:GIB

[Bap07] Ravindra B. Bapat. On generalized inverses of banded matrices. Electronic Journal of Linear Algebra, 16 (1):284–290, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/24.

Bayat:2011:GPE

[Bay11]  Morteza Bayat.  Gener-

alized Pascal k-eliminated functional matrix with 2n variables. Electronic Journal of Linear Algebra, 22 (1):419–429, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/24.

Balaji:2007:BDM

[BB07] R. Balaji and Ravindra B. Bapat. Block distance matrices. Electronic Journal of Linear Algebra, 16(1): 435–443, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol16/iss1/36.

Barrett:2009:MRE

[BBC^{+}09] Wayne Barrett, Ryan Bowcutt, Mark Cutler, Seth Gibelyou, and Kayla Owens. Minimum rank of edge subdivisions of graphs. Electronic Journal of Linear Algebra, 18 (1):530–563, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/42.

Barrett:2014:MNC

[BBC^{+}14] Wayne Barrett, Steve Butler, Minerva Catral, Shaun M. Fallat, H. Tracy Hall, Leslie Hogben, and Michael Young. The maximum nullity of a complete subdivision graph is equal to its zero forcing number. Electronic Journal of Linear Algebra, 27 (1):444–457, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/28.

Baksalary:2012:PSI

[BBCP12] Oskar Maria Baksalary, Natalia Bebiano, Ljiljana Cvetković, and Simo Puntanen. Preface to special issue: International Conference on Matrix Analysis and Its Applications — MatTriad 2011. Electronic Journal of Linear Algebra, 25(1):1–2, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol25/iss1/1.

Berman:2009:SPA

[BCD^{+}09] Abraham Berman, Minerva Catral, Luz Maria DeAlba, Abed Elhashash, Frank J. Hall, Leslie Hogben, In-Jae Kim, Dale D. Olesky, Pablo Tarazaga, Michael J. Tsatsomeros, and Pauline van den Driessche. Sign patterns that allow eventual positivity. Electronic Journal of Linear Algebra, 19 (1):108–120, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol19/iss1/5.

Bru:1996:PCM

[BCG96] Rafael Bru, Carmen Coll, and Josep Gelonch. Periodic coprime matrix fraction decompositions. Electronic Journal of Linear Algebra, 1(1):44–58, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol1/iss1/4.

Berliner:2013:MRM

[BCH^{+}13] Adam Berliner, Minerva Catral, Leslie Hogben, My Huynh, Kelsey Lied, and Michael Young. Minimum rank, maximum nullity, and zero forcing number of simple digraphs. Electronic Journal of Linear Algebra, 26 (1):762–780, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/52.

Bras:2013:NSB

[BCR13] I. Brás, A. C. Carapito, and P. Rocha. A note on simultaneous block diagonally stable matrices. Electronic Journal of Linear Algebra, 26 (1):90–100, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/7.

Bru:2004:RIP

[BCRS04] Rafael Bru, Carmen Coll, Sergio Romero, and Elena Sánchez. Reachability indices of positive linear systems. Electronic Journal of Linear Algebra, 11(1):88–102, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/9.

Belitskii:2009:PCA

[BDL^{+}09] Genrich Belitskii, Andrii R.

Dmytryshyn, Ruvim Lipyanski, Vladimir V. Sergeichuk, and Arkady Tsurkov. Problems of classifying associative or Lie algebras over a field of characteristic not two and finite metabelian groups are wild. Electronic Journal of Linear Algebra, 18 (1):516–529, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/41.

Bebiano:2013:NRQ

[BdP13] N. Bebiano and J. da Providência.

Numerical ranges of quadratic operators in spaces with an indefinite metric. Electronic Journal of Linear Algebra, 26 (1):141–157, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/11.

Bebiano:2010:CNH

[BdPdP10] Natália Bebiano, João da Providência, and João P. da Providência. Classes of non-Hermitian operators with real eigenvalues.

Electronic Journal of Linear Algebra, 21(1):98–109, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol21/iss1/10.

Bebiano:2008:KSN

[BdPNdGPS08] Natalia Bebiano, Joao da Providência, Ana Cristina Nata, and Maria da Graca Pereira Soares. Krein spaces numerical ranges and their computer generation. Electronic Journal of Linear Algebra, 17(1): 192–208, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/15.

Beyn:1997:IPP

[BE97] Wolf-Jurgen Beyn and Ludwig Elsner. Infinite products and paracontracting matrices. Electronic Journal of Linear Algebra, 2(1):1–8, 1997. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol2/iss1/1.

Bahrami:2012:STP

[BE12] F. Bahrami and A. Bayati Eshkaftaki. Some topological properties of the set of linear preservers of majorization. Electronic Journal of Linear Algebra, 23(1):655–663, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/47.

Beasley:1999:SRM

[Bea99] LeRoy B. Beasley. Spaces of rank-2 matrices over GF(2). Electronic Journal of Linear Algebra, 5(1):11–18, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol5/iss1/2.

Bozeman:2014:MRG

[BEH^{+}14] Chassidy Bozeman, AnnaVictoria Ellsworth, Leslie Hogben, Jephian Chin-Hung Lin, Gabi Maurer, Kathleen Nowak, Aaron Rodriguez, and James Strickland. Minimum rank of graphs with loops. Electronic Journal of Linear Algebra, 27(1): 907–934, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/1071.

Bendoukha:2002:DMP

[Ben02] Berrabah Bendoukha. Deux modèles pour les opérateurs de contraction. (French) [Two models for contraction operators]. Electronic Journal of Linear Algebra, 9(1): 171–189, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol9/iss1/18.

Benitez:2010:NDS

[Ben10] Julio Benítez. A new decomposition for square matrices. Electronic Journal of Linear Algebra, 20(1):207–225, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/16.

Bercovici:2003:SVC

[Ber03] Hari Bercovici. Spectral versus classical Nevanlinna–Pick interpolation in dimension two. Electronic Journal of Linear Algebra, 10(1):60–64, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol10/iss1/5.

Barioli:2004:TCR

[BF04] Francesco Barioli and Shaun M. Fallat. On two conjectures regarding an inverse eigenvalue problem for acyclic symmetric matrices. Electronic Journal of Linear Algebra, 11 (1):41–50, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/4.

Berman:2005:ACT

[BF05] Abraham Berman and Karl Heinz Förster. Algebraic connectivity of trees with a pendant edge of infinite weight. Electronic Journal of Linear Algebra, 13(1):175–186, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/14.

Berman:2007:PCV

[BF07] Lihu Berman and Arie Feuer.

On perfect conditioning of Vandermonde matrices on the unit circle. Electronic Journal of Linear Algebra, 16 (1):157–161, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/13.

Bueno:2008:ERP

[BF08] Maribel I. Bueno and Susana Furtado. On the exponent of r-regular primitive matrices. Electronic Journal of Linear Algebra, 17(1):28–47, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/4.

Bueno:2010:GSE

[BF10] Maribel I. Bueno and Susana Furtado. On the gaps in the set of exponents of Boolean primitive circulant matrices. Electronic Journal of Linear Algebra, 20(1):640–660, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/46.

Berman:2011:LBS

[BF11] Abraham Berman and Miriam Farber. A lower bound for the second largest Laplacian eigenvalue of weighted graphs. Electronic Journal of Linear Algebra, 22(1): 1179–1184, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/79.

Bueno:2012:PLM

[BF12] M. I. Bueno and S. Furtado. Palindromic linearizations of a matrix polynomial of odd degree obtained from Fiedler pencils with repetition. Electronic Journal of Linear Algebra, 23 (1):562–577, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/40.

Barioli:2005:VGP

[BFH05] Francesco Barioli, Shaun Fallat, and Leslie Hogben. A variant on the graph parameters of Colin de Verdière: implications to the minimum rank of graphs. Electronic Journal of Linear Algebra, 13 (1):387–404, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol13/iss1/24.

Barioli:2009:MRN

[BFH^{+}09] Francesco Barioli, Shaun M.

Fallat, H. Tracy Hall, Daniel Hershkowitz, Leslie Hogben, Hein van der Holst, Bryan Shader, Bryan Shader, Bryan Shader, Bryan Shader, and Bryan Shader. On the minimum rank of not necessarily symmetric matrices: a preliminary study. Electronic Journal of Linear Algebra, 18 (1):126–145, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/12.

Barioli:2011:MSR

[BFMN11] Francesco Barioli, Shaun M.

Fallat, Lon H. Mitchell, and Sivaram K. Narayan. Minimum semidefinite rank of outerplanar graphs and the tree cover number. Electronic Journal of Linear Algebra, 22(1):10–21, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/2.

Bottcher:1999:TBM

[BG99] Albrecht Böttcher and Sergei M.

Grudsky. Toeplitz band matrices with exponentially growing condition numbers. Electronic Journal of Linear Algebra, 5(1):104–125, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol5/iss1/8.

Butler:2015:UVZ

[BGH15] Steve Butler, Jason Grout, and H. Tracy Hall. Using variants of zero forcing to bound the inertia set of a graph. Electronic Journal of Linear Algebra, 30(1):1–18, 2015. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol30/iss1/1.

Bhatia:2013:DTP

[BGJ13] Rajendra Bhatia, Priyanka Grover, and Tanvi Jain. Derivatives of tensor powers and their norms. Electronic Journal of Linear Algebra, 26 (1):604–619, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/41.

Barrett:2011:IEI

[BGK^{+}11] Wayne Barrett, Seth Gibelyou, Mark Kempton, Nicole Malloy, Curtis Nelson, William Sexton, and John Sinković. The inverse eigenvalue and inertia problems for minimum rank two graphs. Electronic Journal of Linear Algebra, 22(1):389–418, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/23.

Bailo:2009:UBR

[BGRV09] Esteban Bailo, Josep Gelonch, and Sergio Romero-Vivó. An upper bound of the reachability index for a special class of positive 2-D systems. Electronic Journal of Linear Algebra, 18(1):1–12, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/1.

Brualdi:2004:BOC

[BH04] Richard A. Brualdi and Suk-Geun Hwang. A Bruhat order for the class of (0, 1)- matrices with row sum vector R and column sum vector S. Electronic Journal of Linear Algebra, 12(1):6–16, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol12/iss1/3.

Bottcher:2013:NMC

[BH13] Albrecht Böttcher and Martin Halwass. A Newton method for canonical Wiener–Hopf and spectral factorization of matrix polynomials. Electronic Journal of Linear Algebra, 26 (1):873–897, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/59.

Biyikoglu:2009:LED

[BHL09] Türker Biyikoğlu, Marc Hellmuth, and Josef Leydold. Largest eigenvalues of the discrete p-Laplacian of trees with degree sequences. Electronic Journal of Linear Algebra, 18(1):202–210, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/17.

Ben-Israel:2002:MMP

[BI02] Adi Ben-Israel. The Moore of the Moore–Penrose inverse. Electronic Journal of Linear Algebra, 9(1):150–157, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol9/iss1/16.

Bidard:2007:WHS

[Bid07] Christian Bidard. The weak Hawkins–Simon condition. Electronic Journal of Linear Algebra, 16(1):44–59, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/4.

Barrett:2010:ISG

[BJL^{+}10] Wayne Barrett, Camille Jepsen, Robert Lang, Emily McHenry, Curtis Nelson, and Kayla Owens. Inertia sets for graphs on six or fewer vertices. Electronic Journal of Linear Algebra, 20(1):53–78, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/4.

Berman:1997:SPL

[BK97] Abraham Berman and Mark Krupnik. Spectrum preserving lower triangular completions — the nonnegative nilpotent case. Electronic Journal of Linear Algebra, 2(1):9–16, 1997. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol2/iss1/2.

Bukovsek:2007:PCN

[BKNO07] Damjana Kokol Bukovšek, Tomaž Košir, Nika Novak, and Polona Oblak. Products of commuting nilpotent operators. Electronic Journal of Linear Algebra, 16 (1):237–247, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/22.

Berman:2004:PWN

[BL04] Abraham Berman and Thomas J.

Laffey. Preface [2004 Workshop on Nonnegative Matrices]. Electronic Journal of Linear Algebra, 12(1):1, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol12/iss1/1. Held in Bar-

berstown Castle and at the National University of Ireland, Maynooth, July 11–14,

Benitez:2014:SDT

[BL14a] Julio Benítez and Xiaoji Liu. Simultaneous decomposition of two EP matrices with applications. Electronic Journal of Linear Algebra, 27 (1):407–425, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/26.

Bochkarev:2014:SPS

[BL14b] Vladimir V. Bochkarev and Eduard Yu. Lerner. Strong power and subexponential laws for an ordered list of trajectories of a Markov chain. Electronic Journal of Linear Algebra, 27(1):534–556, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/252.

Bebiano:2000:GNR

[BLdP00] Natália Bebiano, Chi-Kwong Li, and João da Providência. Generalized numerical ranges of permanental compounds arising from quantum systems of bosons. Electronic Journal of Linear Algebra, 7(1):73–91, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol7/iss1/7.

Bolotnikov:2002:SMI

[BLM^{+}02] Vladimir Bolotnikov, Chi-Kwong Li, Patrick R. Meade, Christian Mehl, and Leiba Rodman. Shells of matrices in indefinite inner product spaces. Electronic Journal of Linear Algebra, 9(1):67–92, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/7.

Barrett:2013:CIE

[BLM^{+}13] Wayne Barrett, Anne Lazenby, Nicole Malloy, Curtis Nelson, William Sexton, Ryan Smith, John Sincković, and Tianyi Yang. The combinatorial inverse eigenvalue problems: complete graphs and small graphs with strict inequality. Electronic Journal of Linear Algebra, 26(1):656–672, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/44.

Bapat:2009:DMB

[BLP09] Ravindra B. Bapat, Arbind K.

Lal, and Sukanta Pati. The distance matrix of a bidirected tree. Electronic Journal of Linear Algebra, 18 (1):233–245, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/20.

Bolotnikov:2001:MDP

[BLR01] Vladimir Bolotnikov, Chi-Kwong Li, and Leiba Rodman. Minimal distortion problems for classes of unitary matrices. Electronic Journal of Linear Algebra, 8(1):26–46, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol8/iss1/3.

Biyikoglu:2005:NDT

[BLS05] Türker Bıyıkoğlu, Josef Leydold, and Peter F. Stadler. Nodal domain theorems and bipartite subgraphs. Electronic Journal of Linear Algebra, 13(1):344–351, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol13/iss1/21.

Benitez:2010:SRM

[BLZ10] Julio Benítez, Xiaoji Liu, and Jin Zhong. Some results on matrix partial orderings and reverse order law. Electronic Journal of Linear Algebra, 20 (1):254–273, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/19.

Barrett:2013:DER

[BMN^{+}13] Wayne Barrett, Nicole Malloy, Curtis Nelson, William Sexton, and John Sinković. Diagonal entry restrictions in minimum rank matrices. Electronic Journal of Linear Algebra, 26(1):300–332, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/21.

Barik:2008:UGS

[BNPS08] Sasmita Barik, Milan Nath, Sukanta Pati, and Bhaba Kumar Sarma. Unicyclic graphs with the strong reciprocal eigenvalue property. Electronic Journal of Linear Algebra, 17(1):139–153, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/12.

Banica:2014:SHM

[BNS14] Teo Banica, Ion Nechita, and Jean-Marc Schlenker. Submatrices of Hadamard matrices: complementation results. Electronic Journal of Linear Algebra, 27(1): 197–212, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/12.

Barrett:2014:CIE

[BNSY14] Wayne Barrett, Curtis G.

Nelson, John Henry Sinkovic, and Tianyi Yang. The combinatorial inverse eigenvalue problem II: all cases for small graphs. Electronic Journal of Linear Algebra, 27 (1):742–778, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/265.

Bar-On:2000:CNT

[BO00] Ilan Bar-On. Checking nonsingularity of tridiagonal matrices. Electronic Journal of Linear Algebra, 6(1):11–19, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol6/iss1/3.

Boston:2010:SCR

[Bos10] Nigel Boston. Spaces of constant rank matrices over GF(2). Electronic Journal of Linear Algebra, 20(1):1–5, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/1.

Bottcher:2006:SNT

[Böt06] Albrecht Böttcher. Schatten norms of Toeplitz matrices with Fisher–Hartwig singularities. Electronic Journal of Linear Algebra, 15 (1):251–259, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/20.

Bottcher:2011:ARE

[Böt11] Albrecht Böttcher. The algebraic Riccati equation with Toeplitz matrices as coefficients. Electronic Journal of Linear Algebra, 22 (1):348–362, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/20.

Boulet:2009:DUC

[Bou09] Romain Boulet. Disjoint unions of complete graphs characterized by their Laplacian spectrum. Electronic Journal of Linear Algebra, 18 (1):773–783, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/56.

Bourgeois:2011:PMO

[Bou11]  Gerald Bourgeois.  Pairs

of matrices, one of which commutes with their commutator. Electronic Journal of Linear Algebra, 22 (1):593–597, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/38.

Britz:2004:MID

[BOvdD04] Thomas Britz, Dale D.

Olesky, and Pauline van den Driessche. Matrix inversion and digraphs: the one factor case. Electronic Journal of Linear Algebra, 11 (1):115–131, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/11.

Britz:2005:SCM

[BOvdD05] Thomas Britz, Dale D.

Olesky, and Pauline van den Driessche. Schur complements of matrices with acyclic bipartite graphs. Electronic Journal of Linear Algebra, 14(1):2–11, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol14/iss1/2.

Balaji:2007:PMT

[BP07] R. Balaji and T. Parthasarathy. The Q-property of a multiplicative transformation in semidefinite linear complementarity problems. Electronic Journal of Linear Algebra, 16(1):419–428, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/34.

Barreras:2012:AED

[BP12a] A. Barreras and J. M. Peña. Accurate and efficient LDU decomposition of diagonally dominant M-matrices. Electronic Journal of Linear Algebra, 24(1):152–167, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/12.

Barreras:2012:BDM

[BP12b] A. Barreras and J. M. Peña. Bidiagonal decompositions, minors and applications. Electronic Journal of Linear Algebra, 25(1):60–71, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol25/iss1/6.

Bru:2005:SSN

[BPS05] Rafael Bru, Francisco Pedroche, and Daniel B. Szyld. Subdirect sums of nonsingular M-matrices and of their inverses. Electronic Journal of Linear Algebra, 13 (1):162–174, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/13.

Bru:2006:SSS

[BPS06] Rafael Bru, Francisco Pedroche, and Daniel B. Szyld. Subdirect sums of S-strictly diagonally dominant matrices. Electronic Journal of Linear Algebra, 15(1): 201–209, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/15.

Bracic:2009:ARS

[Bra09] Janko Bračič. Algebraic reflexivity for semigroups of operators. Electronic Journal of Linear Algebra, 18 (1):745–760, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/54.

Britz:2007:MPI

[Bri07] Thomas Britz. The Moore– Penrose inverse of a free matrix. Electronic Journal of Linear Algebra, 16 (1):208–215, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol16/iss1/19.

Brown:2013:TBA

[Bro13] George M. F. Brown. Totally bipartite/abipartite Leonard pairs and Leonard triples of Bannai/Itô type. Electronic Journal of Linear Algebra, 26 (1):258–299, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/20.

Brualdi:2002:MNS

[Bru02] Richard A. Brualdi. Maximal nests of subspaces, the matrix Bruhat decomposition, and the marriage theorem — with an application to graph coloring. Electronic Journal of Linear Algebra, 9 (1):118–121, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/11.

Brull:2009:ESR

[Brü09] Tobias Brüll. Explicit solutions of regular linear discrete-time descriptor systems with constant coefficients. Electronic Journal of Linear Algebra, 18 (1):317–338, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/28.

Butkovic:2005:AMA

[BS05] Peter Butković and Hans Schneider. Applications of max algebra to diagonal scaling of matrices. Electronic Journal of Linear Algebra, 13 (1):262–273, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/19.

Bradley:2007:CDL

[BS07] Lisa Wood Bradley and Ernie L. Stitzinger. Class, dimension and length in nilpotent Lie algebras. Electronic Journal of Linear Algebra, 16 (1):429–434, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/35.

Bapat:2012:PDM

[BS12] R. B. Bapat and S. Sivasubramanian. The product distance matrix of a tree and a bivariate zeta function of a graph. Electronic Journal of Linear Algebra, 23 (1):275–286, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/19.

Butkovic:2014:MAC

[BSS14] Peter Butkovic, Hans Schneider, and Sergei Sergeev. On the max-algebraic core of a nonnegative matrix. Electronic Journal of Linear Algebra, 27(1):951–972, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/1075.

Bai:2010:SLA

[BSZ10] Rui-Pu Bai, Cai-Hong Shen, and Yao-Zhong Zhang. Solvable 3-Lie algebras with a maximal hypo-nilpotent ideal N. Electronic Journal of Linear Algebra, 21(1):43–61, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol21/iss1/6.

Bu:2012:NBR

[BSZW12] Changjiang Bu, Lizhu Sun, Jiang Zhou, and Yimin Wei. A note on block representations of the group inverse of Laplacian matrices. Electronic Journal of Linear Algebra, 23(1):866–876, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/61.

Burk:2012:BLM

[BT12] J. Burk and Michael J. Tsatsomeros. On the Brualdi– Li matrix and its Perron eigenspace. Electronic Journal of Linear Algebra, 23 (1):212–230, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/15.

Baksalary:2013:PM

[BT13] Oskar Maria Baksalary and Götz Trenkler. On K-potent matrices. Electronic Journal of Linear Algebra, 26 (1):446–470, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/30.

Bayat:2007:GRH

[BTM07] M. Bayat, H. Teimoori, and B. Mehri. A generalization of rotations and hyperbolic matrices and its applications. Electronic Journal of Linear Algebra, 16(1):125–134, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/11.

Butler:2007:IWG

[But07] Steve Butler. Interlacing for weighted graphs using the normalized Laplacian. Electronic Journal of Linear Algebra, 16(1):90–98, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/8.

Barrett:2004:GWM

[BvdHL04] Wayne Barrett, Hein van der Holst, and Raphael Loewy. Graphs whose minimal rank is two. Electronic Journal of Linear Algebra, 11 (1):258–280, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/21.

Barrett:2005:GWM

[BvdHL05] Wayne Barrett, Hein van der Holst, and Raphael Loewy. Graphs whose minimal rank is two: The finite fields case. Electronic Journal of Linear Algebra, 14(1):32–42, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol14/iss1/4.

Bapat:2003:GIB

[BZ03] Ravindra B. Bapat and Bing Zheng. Generalized inverses of bordered matrices. Electronic Journal of Linear Algebra, 10(1):16–30, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol10/iss1/2.

Bu:2010:ERD

[BZ10] Changjiang Bu and Kuize Zhang. The explicit representations of the Drazin inverses of a class of block matrices.

Electronic Journal of Linear Algebra, 20(1):406–418, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/31.

Bu:2009:SRG

[BZZ09] Changjiang Bu, Jiemei Zhao, and Kuize Zhang. Some results on group inverses of block matrices over skew fields. Electronic Journal of Linear Algebra, 18(1): 117–125, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/11.

Carvalho:2012:NSR

[Car12] Paula Carvalho. A note on strongly regular graphs and (k, τ)-regular sets. Electronic Journal of Linear Algebra, 25(1):3–8, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol25/iss1/2.

Catral:2012:ZFN

[CCH^{+}12] Minerva Catral, Anna Cepek, Leslie Hogben, My Huynh, Kirill Lazebnik, Travis Peters, and Michael Young. Zero forcing number, maximum nullity, and path cover number of subdivided graphs. Electronic Journal of Linear Algebra, 23 (1):906–922, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/64.

Cohen:2000:QD

[CD00] Nir Cohen and Stefano De Leo. The quaternionic determinant. Electronic Journal of Linear Algebra, 7(1): 100–111, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol7/iss1/9.

Choi:2002:MCP

[CDH^{+}02] Ji Young Choi, Luz Maria DeAlba, Leslie Hogben, Mandi S. Maxwell, and Amy Wangsness. The P_{0}-matrix completion problem. Electronic Journal of Linear Algebra, 9(1):1–20, 2002. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol9/iss1/1.

Choi:2003:NMC

[CDH^{+}03] Ji Young Choi, Luz Maria DeAlba, Leslie Hogben, Benard M. Kivunge, Sandra K. Nordstrom, and Mike Shedenhelm. The nonnegative P_{0}-matrix completion problem. Electronic Journal of Linear Algebra, 10(1):46–59, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol10/iss1/4.

Chenette:2007:MRT

[CDH^{+}07] Nathan L. Chenette, Sean V.

Droms, Leslie Hogben, Rana Mikkelson, and Olga Pryporova. Minimum rank of a tree over an arbitrary field. Electronic Journal of Linear Algebra, 16(1):183–186, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/16.

Carmona:2012:JMB

[CEG12] Angeles Carmona, Andres M.

Encinas, and Silvia Gago. Jacobi matrices and boundary value problems in distanceregular graphs. Electronic Journal of Linear Algebra, 24 (1):202–214, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol24/iss1/15.

Catral:2012:SPA

[CEH^{+}12] M. Catral, C. Erickson, L. Hogben, D. D. Olesky, and P. van den Driessche. Sign patterns that allow strong eventual nonnegativity. Electronic Journal of Linear Algebra, 23(1):1–10, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/1.

Carmona:2012:MIP

[CEM12] Ángeles Carmona, Andrés M.

Encinas, and Margarida Mitjana. On the M-matrix inverse problem for singular and symmetric Jacobi matrices. Electronic Journal of Linear Algebra, 24 (1):237–254, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol24/iss1/17.

Cavers:2012:APC

[CF12] Michael S. Cavers and Shaun M. Fallat. Allow problems concerning spectral properties of patterns. Electronic Journal of Linear Algebra, 23(1):731–754, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/52.

Carvalho:2014:DIS

[CF14a] Sónia Carvalho and Pedro J. Freitas. The k-th derivatives of the immanant and the χ-symmetric power of an operator. Electronic Journal of Linear Algebra, 27 (1):284–301, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/18.

Carvalho:2014:NDS

[CF14b] Sónia Carvalho and Pedro J. Freitas. The norm of the k-th derivative of the χsymmetric power of an operator. Electronic Journal of Linear Algebra, 27 (1):302–316, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/19.

Compta:2011:GSS

[CFP11] Albert Compta, Josep Ferrer, and Marta Peña. Geometric structure of single/ combined equivalence classes of a controllable pair. Electronic Journal of Linear Algebra, 22(1):1112–1128, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/73.

Cioaba:2007:PEI

[CG07] Sebastian M. Cioabă and David A. Gregory. Principal eigenvectors of irregular graphs. Electronic Journal of Linear Algebra, 16 (1):366–379, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/31.

Castro-Gonzalez:2013:GIM

[CG13] N. Castro-González. Group inverse of modified matrices over an arbitrary ring. Electronic Journal of Linear Algebra, 26(1):201–214, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/16.

Cavers:2013:TII

[CGK^{+}13] M. S. Cavers, C. Garnett, I.-J. Kim, D. D. Olesky, P. van den Driessche, and K. N. Vander Meulen. Techniques for identifying inertially arbitrary patterns. Electronic Journal of Linear Algebra, 26(1):71–89, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol26/iss1/6.

Castro-Gonzalez:2011:EPA

[CGMSR11] N. Castro-González, M. F. Martínez-Serrano, and J. Robles. An extension of the perturbation analysis for the Drazin inverse. Electronic Journal of Linear Algebra, 22 (1):539–556, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/34.

Campbell:2013:EPC

[CH13] Stephen L. Campbell and Lise E. Holte. Eigenvalue placement in completions of DAES. Electronic Journal of Linear Algebra, 26 (1):520–534, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/33.

Chapman:2012:GPD

[Cha12] Adam Chapman. General polynomials over division algebras and left eigenvalues. Electronic Journal of Linear Algebra, 23(1):508–513, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/34.

Chen:2012:EEN

[Che12] Weixing Chen. On EP elements, normal elements and partial isometries in rings with involution. Electronic Journal of Linear Algebra, 23 (1):553–561, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/39.

Cimpric:2015:RNI

[CHMN15] J. Cimpric, J. W. Helton, S. McCullough, and C. Nelson. Real nullstellensatz and *-ideals in *-algebras. Electronic Journal of Linear Algebra, 30(1):19–50, 2015. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol30/iss1/2.

Catral:2009:SPR

[CHOvdD09] Minerva Catral, Leslie Hogben, Dale D. Olesky, and Pauline van den Driessche. Sign patterns that require or allow power-positivity. Electronic Journal of Linear Algebra, 19(1):121–128, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol19/iss1/6.

Chu:2006:CDA

[Chu06] Wenchang Chu. The Cauchy double alternant and divided differences. Electronic Journal of Linear Algebra, 15 (1):14–21, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/3.

Cvetkovic-Ilic:2012:NCR

[CI12] Dragana S. Cvetković-Ilić. New conditions for the reverse order laws for {1, 3} and {1, 4}-generalized inverses. Electronic Journal of Linear Algebra, 23 (1):231–242, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/16.

Cvetkovic-Ilic:2011:SAR

[CILW11] Dragana S. Cvetković-Ilić, Xiaoji Liu, and Yimin Wei. Some additive results for the generalized Drazin inverse in a Banach algebra. Electronic Journal of Linear Algebra, 22 (1):1049–1058, 2011. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/68.

Cioaba:2010:NSE

[Cio10] Sebastian M. Cioabă. A necessary and sufficient eigenvector condition for a connected graph to be bipartite. Electronic Journal of Linear Algebra, 20(1):351–353, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/26.

Cvetkovic-Ilic:2009:RDI

[CIW09] Dragana S. Cvetković-Ilić and Yimin Wei. Representations for the Drazin inverse of bounded operators on Banach space. Electronic Journal of Linear Algebra, 18 (1):613–627, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/46.

Cheng:2008:AMT

[CJ08] Xuehan Cheng and Wu Jing.

Additivity of maps on triangular algebras. Electronic Journal of Linear Algebra, 17 (1):597–615, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/42.

Chen:2011:MNL

[CJ11] Xuzhou Chen and Jun Ji.

The minimum-norm leastsquares solution of a linear system and symmetric rankone updates. Electronic Journal of Linear Algebra, 22 (1):480–489, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/29.

Cheon:1999:PNZ

[CJLP99] Gi-Sang Cheon, Charles R. Johnson, Sang-Gu Lee, and Ethan J. Pribble. The possible numbers of zeros in an orthogonal matrix. Electronic Journal of Linear Algebra, 5(1):19–23, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol5/iss1/3.

Chooi:2014:LSP

[CKLN14] W. L. Chooi, K. H. Kwa, M. H. Lim, and Z. C. Ng. Linear spaces and preservers of bounded rank-two persymmetric triangular matrices. Electronic Journal of Linear Algebra, 27(1): 619–651, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/257.

Cavers:2005:DMS

[CKSV05] Michael S. Cavers, In-Jae Jae Kim, Bryan L. Shader, and Kevin N. Vander Meulen. On determining minimal spectrally arbitrary patterns. Electronic Journal of Linear Algebra, 13(1):240–248, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/17.

Coppersmith:2004:CFC

[CL04] Don Coppersmith and Zvi Lotker. On Cayley’s formula for counting trees in nested interval graphs. Electronic Journal of Linear Algebra, 11 (1):241–245, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/19.

Cheng:2007:NG

[CL07] Bo Cheng and Bolian Liu.

On the nullity of graphs. Electronic Journal of Linear Algebra, 16(1):60–67, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/5.

Cao:2009:GIM

[CL09a] Chongguang Cao and Juyan Li. Group inverses for matrices over a Bézout domain. Electronic Journal of Linear Algebra, 18(1):600–612, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/45.

Carlen:2009:SPT

[CL09b] Eric A. Carlen and Elliott H.

Lieb. Short proofs of theorems of Mirsky and Horn on diagonals and eigenvalues of matrices. Electronic Journal of Linear Algebra, 18 (1):438–441, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/35.

Chen:2012:NEB

[CLW12] Fubin Chen, Yaotang Li, and Defeng Wang. A new eigenvalue bound for the Hadamard product of an Mmatrix and an inverse Mmatrix. Electronic Journal of Linear Algebra, 23 (1):287–294, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/20.

Chu:2000:DDD

[CM00] Delin Chu and Volker Mehrmann. Disturbance decoupling for descriptor systems by measurement feedback. Electronic Journal of Linear Algebra, 7(1):152–173, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol7/iss1/11.

Cisneros-Molina:2005:IM

[CM05] José Luis Cisneros-Molina. An invariant of 2×2 matrices. Electronic Journal of Linear Algebra, 13(1):146–152, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/10.

Cardoso:2013:CRW

[CMRR13] Domingos M. Cardoso, Enide A.

Martins, Maria Robbiano, and Oscar Rojo. Copies of a rooted weighted graph attached to an arbitrary weighted graph and applications. Electronic Journal of Linear Algebra, 26 (1):706–717, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/47.

Corach:2000:PDU

[CMS00] Gustavo Corach, Alejandra Maestripieri, and Demetrio Stojanoff. Polar decomposition under perturbations of the scalar product. Electronic Journal of Linear Algebra, 7(1):21–29, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol7/iss1/2.

Carvalho:2012:COB

[CMS12] Francisco Carvalho, Joao T.

Mexia, and Carla Santos. Commutative orthogonal block structure and error orthogonal models. Electronic Journal of Linear Algebra, 25 (1):119–128, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol25/iss1/11.

Chien:1998:PDT

[CN98] Mao-Ting Chien and Michael Neumann. Positive definiteness of tridiagonal matrices via the numerical range. Electronic Journal of Linear Algebra, 3(1):93–102, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol3/iss1/9.

Chien:2009:NRW

[CN09] Mao-Ting Chien and Hiroshi Nakazato. The numerical radius of a weighted shift operator with geometric weights. Electronic Journal of Linear Algebra, 18(1):58– 63, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/5.

Chien:2010:NRT

[CN10] Mao-Ting Chien and Hiroshi Nakazato. The q-numerical range of 3 × 3 tridiagonal matrices. Electronic Journal of Linear Algebra, 20 (1):376–390, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/29.

Coffman:2008:REC

[Cof08] Adam Coffman. Real equivalence of complex matrix pencils and complex projections of real Segrè varieties.

Electronic Journal of Linear Algebra, 17(1):651–698, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/45.

Catral:2008:GIM

[COvdD08] Minerva Catral, Dale D.

Olesky, and Pauline van den Driessche. Group inverses of matrices with path graphs. Electronic Journal of Linear Algebra, 17(1):219–233, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/17.

Catral:2009:BRD

[COvdD09] Minerva Catral, Dale D.

Olesky, and Pauline van den Driessche. Block representations of the Drazin inverse of a bipartite matrix. Electronic Journal of Linear Algebra, 18 (1):98–107, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/9.

Climent:1999:CTW

[CP99] Joan-Josep Josep Climent and Carmen Perea. Comparison theorems for weak nonnegative splittings of Kmonotone matrices. Electronic Journal of Linear Algebra, 5(1):24–38, 1999. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol5/iss1/4.

Chorianopoulos:2010:MIN

[CPU10] Christos Chorianopoulos, Panayiotis Psarrakos, and Frank Uhlig. A method for the inverse numerical range problem. Electronic Journal of Linear Algebra, 20 (1):198–206, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/15.

Canto:2009:FRF

[CRU09] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Full rank factorization and the Flanders theorem. Electronic Journal of Linear Algebra, 18(1):352–363, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/30.

Canto:2010:SCT

[CRU10] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Some characterizations of totally nonpositive (totally negative) matrices. Electronic Journal of Linear Algebra, 20(1):241–253, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/18.

Ccapa:2009:SPE

[CS09] Javier Ccapa and Ricardo L.

Soto. On spectra perturbation and elementary divisors of positive matrices. Electronic Journal of Linear Algebra, 18(1):462–481, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/37.

Cao:2004:DPT

[CT04] Chongguang Cao and Xiaomin Tang. Determinant preserving transformations on symmetric matrix spaces. Electronic Journal of Linear Algebra, 11(1):205–211, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/17.

Cavers:2007:IAN

[CV07] Michael S. Cavers and Kevin N. Vander Meulen. Inertially arbitrary nonzero patterns of order 4. Electronic Journal of Linear Algebra, 16(1):30–43, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/3.

Catral:2012:RDI

[CvdD12] M. Catral and P. van den Driessche. Representations for the Drazin inverse of block cyclic matrices. Electronic Journal of Linear Algebra, 23(1):445–456, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/30.

Campbell:2011:SNM

[CVV11] Natalie Campbell, Kevin N. Vander Meulen, and Adam Van Tuyl. Structure of nilpotent matrices over fields. Electronic Journal of Linear Algebra, 22(1):931–958, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/61.

Cai:2013:NPS

[CW13] Xinzhong Cai and Xinmao Wang. A note on the positive semidefinite minimum rank of a sign pattern matrix. Electronic Journal of Linear Algebra, 26(1):345–356, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/23.

Cao:2002:SIK

[CZY02] Chong-Guang Cao, Xian Zhang, and Zhong-Peng Yang. Some inequalities for the Khatri–Rao product of matrices. Electronic Journal of Linear Algebra, 9(1): 276–281, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol9/iss1/23.

Dahl:2003:NLD

[Dah03] Geir Dahl. A note on linear discrepancy. Electronic Journal of Linear Algebra, 10(1):77–80, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/7.

Dahl:2007:CPF

[Dah07]  Geir Dahl.  Combinato-

rial properties of Fourier– Motzkin elimination. Electronic Journal of Linear Algebra, 16(1):334–346, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/29.

deAbreu:2013:MIG

[dAN13] Nair Maria Maia de Abreu and Vladimir Nikiforov. Maxima of the Q-index: graphs with bounded clique number. Electronic Journal of Linear Algebra, 26 (1):121–130, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/9.

Davies:2004:SCM

[Dav04] Philip I. Davies. Structured conditioning of matrix functions. Electronic Journal of Linear Algebra, 11 (1):132–161, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/12.

deTeran:2008:SLB

[DD08] Fernando De Terán and Froilán M. Dopico. Sharp lower bounds for the dimension of linearizations of matrix polynomials. Electronic Journal of Linear Algebra, 17 (1):518–531, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/38.

Diaz:2010:LIS

[DD10] Alexander Diaz and Corey Dunn. The linear independence of sets of two and three canonical algebraic curvature tensors. Electronic Journal of Linear Algebra, 20 (1):436–448, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/33.

DeTeran:2011:CES

[DD11a] Fernando De Terán and Froilán M. Dopico. Consistency and efficient solution of the Sylvester equation for ⋆-congruence. Electronic Journal of Linear Algebra, 22 (1):849–863, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/57.

deTeran:2011:EXD

[DD11b] Fernando De Terán and Froilán M. Dopico. The equation XA + AX∗ = 0 and the dimension of *congruence orbits. Electronic Journal of Linear Algebra, 22 (1):448–465, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/27.

DeLeo:2006:ZUQ

[DDL06] Stefano De Leo, Gisele Ducati, and Vinicius Leonardi. Zeros of unilateral quaternionic polynomials. Electronic Journal of Linear Algebra, 15(1):297–313, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/25.

DeTeran:2009:LSM

[DDM09] Fernando De Terán, Froilán M.

Dopico, and D. Steven Mackey. Linearizations of singular matrix polynomials and the recovery of minimal indices. Electronic Journal of Linear Algebra, 18 (1):371–402, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/32.

DeBruyn:2010:SSP

[De 10] Bart De Bruyn. Some subspaces of the projective space PG(⋀^{K} V) related to regular spreads of PG(V). Electronic Journal of Linear Algebra, 20 (1):354–366, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/27.

DeParis:2013:RWD

[De 13] Alessandro De Paris. A remark on Waring decompositions of some special plane quartics. Electronic Journal of Linear Algebra, 26 (1):510–519, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/32.

daFonseca:2003:PPC

[dF03] C. M. da Fonseca. The path polynomial of a complete graph. Electronic Journal of Linear Algebra, 10 (1):155–162, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/12.

daFonseca:2005:CRC

[dF05] C. M. da Fonseca. On a conjecture regarding characteristic polynomial of a matrix pair. Electronic Journal of Linear Algebra, 13 (1):157–161, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/12.

daFonseca:2010:LBN

[dF10]  C. M. da Fonseca.  A

lower bound for the number of distinct eigenvalues of some real symmetric matrices. Electronic Journal of Linear Algebra, 21(1):3–11, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/2.

deFreitas:2013:MIF

[dFNP13] Maria Aguieiras A. de Freitas, Vladimir Nikiforov, and Laura Patuzzi. Maxima of the Q-index, forbidden 4-ycle and 5-cycle. Electronic Journal of Linear Algebra, 26 (1):905–916, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/61.

Dragomir:2013:BOC

[DFS13] Sever S. Dragomir, Jun Ichi Fujii, and Yuki Seo. Bounds for an operator concave function. Electronic Journal of Linear Algebra, 26(1): 192–200, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/15.

DeAlba:2009:UOM

[DGH^{+}09] Luz M. DeAlba, Jason Grout, Leslie Hogben, Rana Mikkelson, and Kaela Rasmussen. Universally optimal matrices and field independence of the minimum rank of a graph. Electronic Journal of Linear Algebra, 18(1):403–419, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/33.

DeAlba:2012:MRP

[DGK^{+}12] Luz M. DeAlba, Jason Grout, In-Jae Kim, Steve Kirkland, Judith J. McDonald, and Amy Yielding. Minimum rank of powers of trees. Electronic Journal of Linear Algebra, 23(1):151–163, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/10.

Deng:2013:MSL

[DH13] Hanyuan Deng and He Huang. On the main signless Laplacian eigenvalues of a graph. Electronic Journal of Linear Algebra, 26(1):381–393, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/25.

DeAlba:2003:WSS

[DHHW03] Luz Maria DeAlba, Timothy L. Hardy, Leslie Hogben, and Amy Wangsness. The (weakly) sign symmetric P - matrix completion problems. Electronic Journal of Linear Algebra, 10(1):257–271, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/21.

DeAlba:2009:MCP

[DHS09] Luz Maria DeAlba, Leslie Hogben, and Bhaba Kumar Sarma. The Q-matrix completion problem. Electronic Journal of Linear Algebra, 18 (1):176–191, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/15.

Dickinson:2010:ICI

[Dic10] Peter J. C. Dickinson. An improved characterisation of the interior of the completely positive cone. Electronic Journal of Linear Algebra, 20 (1):723–729, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/51.

Ding:2006:MRF

[DK06] Guoli Ding and Andreĭ Kotlov. On minimal rank over finite fields. Electronic Journal of Linear Algebra, 15 (1):210–214, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/16.

Dolinar:2008:GPQ

[DK08] Gregor Dolinar and Bojan Kuzma. General preservers of quasi-commutativity on Hermitian matrices. Electronic Journal of Linear Algebra, 17 (1):436–444, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/31.

Dolinar:2012:MDC

[DKO12] Gregor Dolinar, Bojan Kuzma, and Polona Oblak. On maximal distances in a commuting graph. Electronic Journal of Linear Algebra, 23 (1):243–256, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/17.

Dolzan:2013:DCC

[DKO13] David Dolžan, Matjaž Konvalinka, and Polona Oblak. Diameters of connected components of commuting graphs. Electronic Journal of Linear Algebra, 26 (1):433–445, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/29.

Dmytryshyn:2014:SMP

[DKS14] Andrii Dmytryshyn, Bo Kågström, and Vladimir V. Sergeichuk. Symmetric matrix pencils: codimension counts and the solution of a pair of matrix equations. Electronic Journal of Linear Algebra, 27(1):1– 18, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/1.

DeSterck:2012:SMI

[DL12] Hans De Sterck and Minghua Lin. Some majorization inequalities for coneigenvalues. Electronic Journal of Linear Algebra, 23(1):669–677, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/49.

Deaett:2013:TCS

[DLRM^{+}13] Louis Deaett, Ramiro H.

Lafuente-Rodriguez, Juan Marin, Jr., Erin Haller Martin, Linda J. Patton, Kathryn Rasmussen, and Rebekah B. Johnson Yates. Trace conditions for symmetry of the numerical range. Electronic Journal of Linear Algebra, 26(1):591–603, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/40.

Drury:2003:NMG

[DMLT03] S. W. Drury, J. K. Merikoski, V. Laakso, and T. Tossavainen. On nonnegative matrices with given row and column sums. Electronic Journal of Linear Algebra, 10 (1):280–290, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol10/iss1/23.

Doering:2011:EOT

[DMS11] Elizabeth Doering, T. S. Michael, and Bryan L. Shader. Even and odd tournament matrices with minimum rank over finite fields. Electronic Journal of Linear Algebra, 22(1):363–377, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/21.

DeAbreu:2012:MIA

[DN12] Nair M. M. De Abreu and Vladimir Nikiforov. Maxima of the Q-index: abstract graph properties. Electronic Journal of Linear Algebra, 23 (1):782–789, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/55.

Dodig:2007:CSC

[Dod07] Marija Dodig. Controllability of series connections. Electronic Journal of Linear Algebra, 16(1):135–156, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/12.

Dodig:2008:EPP

[Dod08] Marija Dodig. Eigenvalues of partially prescribed matrices.

Electronic Journal of Linear Algebra, 17(1):316–332, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/23.

Deaett:2010:RIS

[DOvdD10] Louis Deaett, Dale D. Olesky, and Pauline van den Driessche. Refined inertially and spectrally arbitrary zerononzero patterns. Electronic Journal of Linear Algebra, 20 (1):449–467, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/34.

Dragomir:2011:VOT

[Dra11] Sever S. Dragomir. Vector and operator trapezoidal type inequalities for continuous functions of selfadjoint operators in Hilbert spaces. Electronic Journal of Linear Algebra, 22(1):161–178, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/10.

Drury:2004:EVC

[Dru04] S. W. Drury. The external vertices conjecture in case n = 4. Electronic Journal of Linear Algebra, 11 (1):180–191, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/15.

Drury:2006:EHM

[Dru06] Stephen W. Drury. Essentially Hermitian matrices revisited. Electronic Journal of Linear Algebra, 15 (1):285–296, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/24.

Drury:2009:ONW

[Dru09] Stephen W. Drury. Operator norms of words formed from positive-definite matrices. Electronic Journal of Linear Algebra, 18(1):13–20, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/2.

Day:2007:SVI

[DS07] Jane Day and Wasin So. Singular value inequality and graph energy change. Electronic Journal of Linear Algebra, 16(1):291–299, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/25.

Dur:2008:IPC

[DS08] Mirjam Dür and Georg Still. Interior points of the completely positive cone. Electronic Journal of Linear Algebra, 17(1):48–53, 2008.

CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/5.

Deschamp:2010:PHM

[DS10] Brent J. Deschamp and Bryan L. Shader. Permanents of Hessenberg (0, 1)- matrices revisited. Electronic Journal of Linear Algebra, 20(1):20–52, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/3.

Dodig:2013:PGM

[DS13] Marija Dodig and Marko Stošić. On properties of the generalized majorization. Electronic Journal of Linear Algebra, 26(1):471–509, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/31.

Pazzis:2014:CHM

[dSP14] Clément de Seguins Pazzis.

Commutators from a hyperplane of matrices. Electronic Journal of Linear Algebra, 27(1):39–54, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/3.

Dubeau:2008:LAS

[Dub08] François Dubeau. Linear algebra and the sums of powers of integers. Electronic Journal of Linear Algebra, 17 (1):577–596, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/41.

Deng:2010:PGD

[DW10] Chunyuan Deng and Yimin Wei. Perturbation of the generalized Drazin inverse. Electronic Journal of Linear Algebra, 21(1):85–97, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/9.

Du:2012:PAB

[DX12] Fapeng Du and Yifeng Xue. Perturbation analysis of A_{T,S}^{(2)} on Banach spaces. Electronic Journal of Linear Algebra, 23 (1):586–598, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/42.

Du:2011:EIG

[DZ11] Zhibin Du and Bo Zhou. On the Estrada index of graphs with given number of cut edges. Electronic Journal of Linear Algebra, 22 (1):586–592, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/37.

Ekstrand:2012:NPS

[EEH^{+}12] J. Ekstrand, C. Erickson, D. Hay, L. Hogben, and J. Roat. Note on positive semidefinite maximum nullity and positive semidefinite zero forcing number of partial 2-trees. Electronic Journal of Linear Algebra, 23(1):79–87, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/5.

Edwards:2012:SPP

[EFP12] Roderick Edwards, Eric Foxall, and Theodore J. Perkins. Scaling properties of paths on graphs. Electronic Journal of Linear Algebra, 23 (1):966–988, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/68.

Elzinga:2010:WME

[EG10] Randall J. Elzinga and David A. Gregory. Weighted matrix eigenvalue bounds on the independence number of a graph. Electronic Journal of Linear Algebra, 20 (1):468–489, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/35.

Ellison:2009:SPR

[EHT09] Elisabeth M. Ellison, Leslie Hogben, and Michael J. Tsatsomeros. Sign patterns that require eventual positivity or require eventual nonnegativity. Electronic Journal of Linear Algebra, 19 (1):98–107, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol19/iss1/4.

ElKadiri:2011:CFB

[El 11] Mohamed El Kadiri. Sur une classe de formes biquadratiques semi-définies positives. (French) [On a class of positive semidefinite quadratric forms]. Electronic Journal of Linear Algebra, 22 (1):513–520, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/32.

Elzinga:2003:SRG

[Elz03] Randall J. Elzinga. Strongly regular graphs: values of λ and μ for which there are only finitely many feasible (v, k, λ, μ). Electronic Journal of Linear Algebra, 10 (1):232–239, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/19.

Eagambaram:2013:CSD

[EPM13] N. Eagambaram, K. Manjunatha Prasad, and K. S. Mohana. Column space decomposition and partial order on matrices. Electronic Journal of Linear Algebra, 26 (1):795–815, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/54.

Erovenko:2004:BGE

[Ero04] Igor V. Erovenko. SL_{n}(F [x])

is not boundedly generated by elementary matrices: explicit proof. Electronic Journal of Linear Algebra, 11 (1):162–167, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/13.

Elhashash:2009:PMS

[ERS09] Abed Elhashash, Uriel G.

Rothblum, and Daniel B. Szyld. Paths of matrices with the strong Perron– Frobenius property converging to a given matrix with the Perron–Frobenius property. Electronic Journal of Linear Algebra, 19(1):90–97, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol19/iss1/3.

Elhashash:2008:GMH

[ES08] Abed Elhashash and Daniel B.

Szyld. On general matrices having the Perron–Frobenius property. Electronic Journal of Linear Algebra, 17 (1):389–413, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/28.

Elhashash:2010:MFP

[ES10] Abed Elhashash and Daniel B. Szyld. Matrix functions preserving sets of generalized nonnegative matrices. Electronic Journal of Linear Algebra, 20(1):673–690, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/48.

Fang:2011:LMP

[Fan11] Li Fang. Linear maps preserving the idempotency of Jordan products of operators. Electronic Journal of Linear Algebra, 22(1): 767–779, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/51.

Fayziev:2000:G

[Fay00] Valeriy Fayziev. On the group GL(2, R[x]). Electronic Journal of Linear Algebra, 7(1):59–72, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol7/iss1/6.

Feng:2001:ACS

[Fen01]  Zhaosheng Feng.  Alge-

braic curve solution for second-order polynomial autonomous systems. Electronic Journal of Linear Algebra, 8(1):14–25, 2001. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol8/iss1/2.

Fernandes:2009:IEP

[Fer09]  Rosário Fernandes.  On

the inverse eigenvalue problems: the case of superstars. Electronic Journal of Linear Algebra, 18(1):442–461, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/36.

Fonseca:2012:PDM

[FFB12] Carlos M. Fonseca, Shmuel Friedland, and Oskar Maria Baksalary. Preface to Directions in Matrix Theory Conference. Electronic Journal of Linear Algebra, 24(1):1–2, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/1.

Farber:2014:ERP

[FFJM14] Miriam Farber, Mitchell Faulk, Charles R. Johnson, and Evan Marzion. Exact results for perturbation to total positivity and to total nonsingularity. Electronic Journal of Linear Algebra, 27 (1):779–797, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/266.

Freitas:2012:UBM

[FFP12] Pedro J. Freitas, Shmuel Friedland, and Gaspar Porta. Upper bounds on the magnitude of solutions of certain linear systems with integer coefficients. Electronic Journal of Linear Algebra, 24 (1):113–125, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol24/iss1/10.

Fakhari:2014:SPR

[FG14] Abbas Fakhari and Ali Golmakani. Shadowing property of random linear cocycles. Electronic Journal of Linear Algebra, 27(1):190–196, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/11.

Fourie:2014:RCI

[FGVR14] J. H. Fourie, G. J. Groenwald, D. B. Janse Van Rensburg, and A. C. M. Ran. Real and complex invariant subspaces for matrices which are H-positive real in an indefinite inner product space. Electronic Journal of Linear Algebra, 27(1):124–145, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/6.

Fricke:1996:RAM

[FHJT96] Gerd H. Fricke, Stephen T.

Hedetniemi, David P. Jacobs, and Vilmar Trevisan. Reducing the adjacency matrix of a tree. Electronic Journal of Linear Algebra, 1(1):34–43, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol1/iss1/3.

Friedland:2004:PES

[FHR04] Shmuel Friedland, Daniel Hershkowitz, and Siegfried M. Rump. Positive entries of stable matrices. Electronic Journal of Linear Algebra, 12(1):17–24, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol12/iss1/4.

Fiedler:1998:USE

[Fie98] Miroslav Fiedler. Ultrametric sets in Euclidean point spaces. Electronic Journal of Linear Algebra, 3(1):23–30, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol3/iss1/5.

Fiedler:2005:MGE

[Fie05] Miroslav Fiedler. Matrices and graphs in Euclidean geometry. Electronic Journal of Linear Algebra, 14(1):51–58, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol14/iss1/6.

Fallat:2000:GTP

[FJS00] Shaun M. Fallat, Charles R. Johnson, and Ronald L. Smith. The general totally positive matrix completion problem with few unspecified entries. Electronic Journal of Linear Algebra, 7(1):1–20, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol7/iss1/1.

Fallat:1998:EAC

[FK98] Shaun Fallat and Steve Kirkland. Extremizing algebraic connectivity subject to graph theoretic constraints. Electronic Journal of Linear Algebra, 3(1):48–74, 1998. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol3/iss1/7.

Fritzsche:2008:NCS

[FKS08] Bernd Fritzsche, Bernd Kirstein, and Alexander L. Sakhnovich. On a new class of structured matrices related to the discrete skewself-adjoint Dirac systems. Electronic Journal of Linear Algebra, 17(1):473–486, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/34.

Futorny:2014:CCC

[FKS14] Vyacheslav Futorny, Lena Klimenko, and Vladimir V. Sergeichuk. Change of the *-congruence canonical form of 2-by-2 matrices under perturbations. Electronic Journal of Linear Algebra, 27 (1):146–154, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/7.

Fallat:2013:CVP

[FM13] Shaun M. Fallat and Lon H.

Mitchell. Colin de Verdière parameters of chordal graphs. Electronic Journal of Linear Algebra, 26(1):49–62, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/4.

Friedland:2011:FLR

[FMMN11] S. Friedland, V. Mehrmann, A. Miedlar, and M. Nkengla. Fast low rank approximations of matrices and tensors. Electronic Journal of Linear Algebra, 22(1): 1031–1048, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/67.

Filipiak:2012:DMM

[FMS12] Katarzyna Filipiak, Augustyn Markiewicz, and Aneta Sawikowska. Determinants of multidiagonal matrices. Electronic Journal of Linear Algebra, 25 (1):102–118, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol25/iss1/10.

Forster:1998:NOF

[FN98] K.-H. H. Förster and B. Nagy. On nonnegative operators and fully cyclic peripheral spectrum. Electronic Journal of Linear Algebra, 3(1):13– 22, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol3/iss1/4.

Forster:2002:SEG

[FN02] Karl Heinz Förster and Bela Nagy. On spectra of expansion graphs and matrix polynomials, II. Electronic Journal of Linear Algebra, 9 (1):158–170, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/17.

Forster:2006:ITH

[FN06] Karl Heinz Förster and Bela Nagy. Irreducible Toeplitz and Hankel matrices. Electronic Journal of Linear Algebra, 15(1):274–284, 2006.

CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/23.

Foth:2010:ESP

[Fot10] Philip Foth. Eigenvalues of sums of pseudo-Hermitian matrices. Electronic Journal of Linear Algebra, 20 (1):115–125, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/9.

Fujimoto:2004:TCI

[FR04] Takao Fujimoto and Ravindra R. Ranade. Two characterizations of inverse-positive matrices: the Hawkins– Simon condition and the Le Chatelier–Braun principle. Electronic Journal of Linear Algebra, 11(1):59–65, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/6.

Friedland:2012:ROP

[Fri12]  Shmuel Friedland.  Re-

sults and open problems in matchings in regular graphs. Electronic Journal of Linear Algebra, 24(1):18–33, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/4.

Fernandes:2012:CBS

[FRM12] Célia Fernandes, Paulo Ramos, and João T. Mexia. Crossing balanced and stair nested designs. Electronic Journal of Linear Algebra, 25(1):22–48, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol25/iss1/4.

Furtado:1998:CPM

[FS98] Susana Furtado and Fernando C. Silva. On the characteristic polynomial of matrices with prescribed columns and the stabilization and observability of linear systems. Electronic Journal of Linear Algebra, 4(1):19– 31, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol4/iss1/2.

Friedland:2000:SEG

[FS00] Shmuel Friedland and Hans Schneider. Spectra of expansion graphs. Electronic Journal of Linear Algebra, 6(1):2– 10, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol6/iss1/2.

Fallat:2002:CTP

[FT02] Shaun M. Fallat and Michael J. Tsatsomeros. On the Cayley transform of positivity classes of matrices. Electronic Journal of Linear Algebra, 9 (1):190–196, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol9/iss1/19.

Friedland:2008:NSC

[FV08] Shmuel Friedland and Elena Virnik. Nonnegativity of Schur complements of nonnegative idempotent matrices. Electronic Journal of Linear Algebra, 17(1): 426–435, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/30.

Fallat:2000:MAM

[FvdD00] Shaun M. Fallat and Pauline van den Driessche. On matrices with all minors negative. Electronic Journal of Linear Algebra, 7(1):92–99, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol7/iss1/8.

Ferrante:2010:RMC

[FW10] Augusto Ferrante and Harald K. Wimmer. Reachability matrices and cyclic matrices. Electronic Journal of Linear Algebra, 20(1):95–102, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/7.

Xiao:2009:SMR

[fXyHZ09] Qing feng Xiao, Xi yan Hu, and Lei Zhang. The symmetric minimal rank solution of the matrix equation AX = B and the optimal approximation. Electronic Journal of Linear Algebra, 18 (1):264–273, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/23.

Yuan:2012:TSK

[fYwW12] Shi fang Yuan and Qing wen Wang. Two special kinds of least squares solutions for the quaternion matrix equation AXB + CXD = E. Electronic Journal of Linear Algebra, 23(1):257–274, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/18.

Gavruta:2013:FC

[Găv13] Pasc Găvruţa. On the Feichtinger conjecture. Electronic Journal of Linear Algebra, 26(1):546–552, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol26/iss1/35.

Garrity:2004:RIV

[GG04] Thomas Garrity and Zachary Grossman. On relations of invariants for vector-valued forms. Electronic Journal of Linear Algebra, 11(1):24–40, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/3.

Grood:2014:MRZ

[GHH^{+}14] Cheryl Grood, Johannes Harmse, Leslie Hogben, Thomas J. Hunter, Bonnie Jacob, Andrew Klimas, and Sharon McCathern. Minimum rank with zero diagonal. Electronic Journal of Linear Algebra, 27 (1):458–477, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/29.

Gumus:2014:ELC

[GHK14] Ibrahim Halil Gumus, Omar Hirzallah, and Fuad Kittaneh. Eigenvalue localization for complex matrices. Electronic Journal of Linear Algebra, 27(1):892–906, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/1070.

Gil:2005:NEF

[Gil05] Michael Gil. Norm estimates for functions of two commuting matrices. Electronic Journal of Linear Algebra, 13 (1):122–130, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol13/iss1/8.

Gil:2007:EAV

[Gil07a] Michael I. Gil’. Estimates for absolute values of matrix functions. Electronic Journal of Linear Algebra, 16 (1):444–450, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol16/iss1/37.

Gil:2007:SIB

[Gil07b] Michael I. Gil’. Spectrum of infinite block matrices and π-triangular operators. Electronic Journal of Linear Algebra, 16(1):216–231, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/20.

Gil:2010:PFD

[Gil10] Michael I. Gil’. Perturbations of functions of diagonalizable matrices. Electronic Journal of Linear Algebra, 20 (1):303–313, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/22.

Gil:2011:NEF

[Gil11] Michael Gil’. Norm estimates for functions of two noncommuting matrices. Electronic Journal of Linear Algebra, 22(1):504–512, 2011.

CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/31.

Gil:2014:BCN

[Gil14] Michael I. Gil’. A bound for condition numbers of matrices. Electronic Journal of Linear Algebra, 27 (1):162–171, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/9.

Grundy:2007:PMN

[GJOvdD07] David A. Grundy, Charles R.

Johnson, Dale D. Olesky, and Pauline van den Driessche. Products of M-matrices and nested sequences of principal minors. Electronic Journal of Linear Algebra, 16 (1):380–388, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/32.

Gregory:1999:SVT

[GK99] David A. Gregory and Stephen J. Kirkland. Singular values of tournament matrices. Electronic Journal of Linear Algebra, 5(1):39–52, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol5/iss1/5.

Gursoy:2013:MCT

[GKMS13] Buket Benek Gursoy, Steve Kirkland, Oliver Mason, and Sergeĭ Sergeev. On the Markov chain tree theorem in the max algebra. Electronic Journal of Linear Algebra, 26(1):15–27, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/2.

Gonzalez:2002:IRD

[GKW02] N. Castro González, J. J. Koliha, and Yimin Wei. Integral representation of the Drazin inverse. Electronic Journal of Linear Algebra, 9 (1):129–131, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/13.

Geng:2010:SRQ

[GL10] Xianya Geng and Shuchao Li. On the spectral radii of quasi-tree graphs and quasiunicyclic graphs with k pendant vertices. Electronic Journal of Linear Algebra, 20 (1):391–405, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/30.

Goldberg:2011:PE

[GL11] Moshe Goldberg and Eliahu Levy. The power equation. Electronic Journal of Linear Algebra, 22(1):810–821, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/54.

Liu:2013:SAM

[gLgWhX13] Xin guo Liu, Wei guo Wang, and Lian hua Xu. Sensitivity analysis for the multivariate eigenvalue problem. Electronic Journal of Linear Algebra, 26(1):636–655, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/43.

Guralnick:2003:MMI

[GLR03] Robert M. Guralnick, Chi-Kwong Li, and Leiba X. Rodman. Multiplicative maps on invertible matrices that preserve matricial properties. Electronic Journal of Linear Algebra, 10(1):291–319, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/24.

Gong:2014:OGM

[GLX14] Shi-Cai Gong, Xueliang Li, and Guanghui Xu. On oriented graphs with minimal skew energy. Electronic Journal of Linear Algebra, 27 (1):692–704, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/261.

Goldberger:2002:SFC

[GN02] Assaf Goldberger and Michael Neumann. On a strong form of a conjecture of Boyle and Handelman. Electronic Journal of Linear Algebra, 9(1): 138–149, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/15.

Godsil:2012:WCP

[God12] Chris Godsil. When can perfect state transfer occur? Electronic Journal of Linear Algebra, 23(1):877–890, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/62.

Goldberg:2008:SSF

[Gol08] Moshe Goldberg. Stable subnorms on finite-dimensional power-associative algebras. Electronic Journal of Linear Algebra, 17(1):359–375, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/26.

Gong:2012:CSR

[Gon12] Shi-Cai Gong. The characteristic set with respect to k-maximal vectors of a tree. Electronic Journal of Linear Algebra, 23(1):66–78, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/4.

Garnett:2013:RIT

[GOvdD13] Colin Garnett, D. D. Olesky, and Pauline van den Driessche. Refined inertias of tree sign patterns. Electronic Journal of Linear Algebra, 26 (1):620–635, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/42.

Gowda:2012:CCP

[Gow12] M. Seetharama Gowda. On copositive and completely positive cones, and Ztransformations. Electronic Journal of Linear Algebra, 23 (1):198–211, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/14.

Groenewald:2003:SFR

[GP03] Gilbert J. Groenewald and Mark A. Petersen. J-spectral factorization for rational matrix functions with alternative realization. Electronic Journal of Linear Algebra, 10 (1):240–256, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/20.

Glebsky:2009:LRP

[GR09] Lev Glebsky and Luis Manuel Rivera. On low rank perturbations of complex matrices and some discrete metric spaces. Electronic Journal of Linear Algebra, 18 (1):302–316, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/27.

Greene:2014:TMP

[Gre14] John Greene. Traces of matrix products. Electronic Journal of Linear Algebra, 27 (1):716–734, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/263.

Grout:2010:MRP

[Gro10] Jason Grout. The minimum rank problem over finite fields. Electronic Journal of Linear Algebra, 20 (1):691–716, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/49.

Goldberg:2003:MIG

[GS03] Felix Goldberg and Gregory Shapiro. The Merris index of a graph. Electronic Journal of Linear Algebra, 10 (1):212–222, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol10/iss1/17.

Gao:2011:IAT

[GS11] Yubin Gao and Yanling Shao.

Inertially arbitrary tree sign patterns of order 4. Electronic Journal of Linear Algebra, 22(1):1148–1155, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/76.

Gowda:2013:ISD

[GS13] M. Seetharama Gowda and R. Sznajder. On the irreducibility, self-duality, and non-homogeneity of completely positive cones. Electronic Journal of Linear Algebra, 26(1):177–191, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/14.

Gao:2009:SAC

[GSF09] Yubin Gao, Yanling Shao, and Yizheng Fan. Spectrally arbitrary complex sign pattern matrices. Electronic Journal of Linear Algebra, 18 (1):674–692, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/49.

Goldberg:2014:MAB

[GSM14] Felix Goldberg and Naomi Shaked-Monderer. On the maximal angle between copositive matrices. Electronic Journal of Linear Algebra, 27(1):837–850, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/1065.

Grassmann:2005:TSQ

[GT05] Winfried K. Grassmann and Javad Tavakoli. Two-station queueing networks with moving servers, blocking, and customer loss. Electronic Journal of Linear Algebra, 13(1):72–89, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/5.

Guseinov:2013:RCC

[Gus13] Gusein Sh. Guseinov. RE-TRACTED: On construction of a complex finite Jacobi matrix from two spectra. Electronic Journal of Linear Algebra, 26(1):101–120, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/8.

Han:2005:CTM

[Han05] Yongho Han. Commuting triples of matrices. Electronic Journal of Linear Algebra, 13 (1):274–343, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/20.

Handelman:2009:MPP

[Han09] David Handelman. Matrices of positive polynomials. Electronic Journal of Linear Algebra, 19(1):2–89, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol19/iss1/2.

Hardy:2014:KQ

[Har14] Yorick Hardy. On Kronecker quotients. Electronic Journal of Linear Algebra, 27 (1):172–189, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/10.

Hershkowitz:1998:P

[HB98] Daniel Hershkowitz and Richard A. Brualdi. Preface. Electronic Journal of Linear Algebra, 3(1):2–3, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol3/iss1/2.

Hashemi:2010:ECE

[HD10] Behnam Hashemi and Mehdi Dehghan. Efficient computation of enclosures for the exact solvents of a quadratic matrix equation. Electronic Journal of Linear Algebra, 20 (1):519–536, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/38.

Henrion:2010:SGN

[Hen10] Didier Henrion. Semidefinite geometry of the numerical range. Electronic Journal of Linear Algebra, 20 (1):322–332, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/24.

Hershkowitz:2000:P

[Her00] Daniel Hershkowitz. Preface. Electronic Journal of Linear Algebra, 6(1):1, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol6/iss1/1.

Hillar:2002:PET

[HJS02] Christopher J. Hillar, Charles R. Johnson, and Ilya M. Spitkovsky. Positive eigenvalues and twoletter generalized words. Electronic Journal of Linear Algebra, 9(1):21–26, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol9/iss1/2.

Horn:1998:TAB

[HK98] Roger A. Horn and Fuad Kittaneh. Two applications of a bound on the Hadamard product with a Cauchy matrix. Electronic Journal of Linear Algebra, 3(1):4–12, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/

vol3/iss1/3. Dedicated to Hans Schneider on the occasion of his 70th birthday.

Hershkowitz:2005:SPS

[HK05] Daniel Hershkowitz and Nathan Keller. Spectral properties of sign symmetric matrices. Electronic Journal of Linear Algebra, 13 (1):90–110, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/6.

Hashemi:2014:FPT

[HKT14] Benham Hashemi, Mahtab Mirzaei Khalilabadi, and Hanieh Tavakolipour. On the fixedpoint type Sylvester matrix equations over complete commutative diods. Electronic Journal of Linear Algebra, 27 (1):882–891, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/1069.

Hecker:2007:ULS

[HL07] David Hecker and Deborah Lurie. Using least-squares to find an approximate eigenvector. Electronic Journal of Linear Algebra, 16 (1):99–110, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol16/iss1/9.

Li:2012:BGW

[hLSxS12] Hong hai Li, Li Su, and Hui xian Sun. On bipartite graphs which attain minimum rank among bipartite graphs with a given diameter. Electronic Journal of Linear Algebra, 23(1):137–150, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/9.

Hershkowitz:1998:MLS

[HM98] Daniel Hershkowitz and Nira Mashal. P^{α}-matrices and Lyapunov scalar stability. Electronic Journal of Linear Algebra, 4(1):39–47, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol4/iss1/4.

Hartwig:2005:FP

[HM05] Robert E. Hartwig and John Maroulas. Focal power. Electronic Journal of Linear Algebra, 13(1):56–71, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol13/iss1/4.

Ha:2012:ARL

[HM12a] Phi Ha and Volker Mehrmann. Analysis and reformulation of linear delay differentialalgebraic equations. Electronic Journal of Linear Algebra, 23(1):703–730, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/51.

Huang:2012:RBH

[HM12b] Qianglian Huang and Mohammad Sal Moslehian. Relationship between the Hyers–Ulam stability and the Moore–Penrose inverse. Electronic Journal of Linear Algebra, 23(1):891–905, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/63.

Hogben:2005:SGT

[Hog05]  Leslie Hogben.  Spectral

graph theory and the inverse eigenvalue problem of a graph. Electronic Journal of Linear Algebra, 14(1):12– 31, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol14/iss1/3.

Hogben:2009:ECM

[Hog09] Leslie Hogben. Eventually cyclic matrices and a test for strong eventual nonnegativity. Electronic Journal of Linear Algebra, 19 (1):129–140, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol19/iss1/7.

Hogben:2011:NMR

[Hog11] Leslie Hogben. A note on minimum rank and maximum nullity of sign patterns. Electronic Journal of Linear Algebra, 22(1):203–213, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/13.

Holtz:2000:CIM

[Hol00] Olga Holtz. On convergence of infinite matrix products. Electronic Journal of Linear Algebra, 7(1):178–181, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol7/iss1/13.

Hartwig:2007:RPE

[HP07a] Robert E. Hartwig and Pedro Patricio. On Roth’s pseudo equivalence over rings. Electronic Journal of Linear Algebra, 16(1):111–124, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/10.

Hershkowitz:2007:NSE

[HP07b] Daniel Hershkowitz and Allan Pinkus. On nonnegative sign equivalent and sign similar factorizations of matrices.

Electronic Journal of Linear Algebra, 16(1):162–170, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/14.

Haemers:2012:MOR

[HP12a] W. H. Haemers and M. J. P.

Peeters. The maximum order of reduced square (0, 1)- matrices with a given rank. Electronic Journal of Linear Algebra, 24(1):3–6, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/2.

Hartwig:2012:NPB

[HP12b] R. E. Hartwig and P. Patricio. A note on power bounded matrices. Electronic Journal of Linear Algebra, 23 (1):625–645, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/45.

He:2012:SSL

[HP12c] Chang-Xiang He and Hao Pan. The smallest signless Laplacian eigenvalue of graphs under perturbation. Electronic Journal of Linear Algebra, 23(1):473–482, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/32.

Hartwig:2014:TTU

[HP14] R. E. Hartwig and P. Patrício. Two by two units. Electronic Journal of Linear Algebra, 27 (1):489–503, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/31.

Horn:2009:EPD

[HPP09] Roger A. Horn, Giuseppe Piazza, and Tiziano Politi. Explicit polar decompositions of complex matrices. Electronic Journal of Linear Algebra, 18 (1):693–699, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/50.

Hasani:2006:SLO

[HR06] Ahmad M. Hasani and Mehdi Radjabalipour. The structure of linear operators strongly preserving majorizations of matrices. Electronic Journal of Linear Algebra, 15 (1):260–268, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/21.

Hershkowitz:2003:OSS

[HS03] Daniel Hershkowitz and Hans Schneider. One-sided simultaneous inequalities and sandwich theorems for diagonal similarity and diagonal equivalence of nonnegative matrices. Electronic Journal of Linear Algebra, 10 (1):81–101, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/8.

Hou:2010:SEC

[HS10] Yaoping Hou and Wai-Chee Shiu. The spectrum of the edge corona of two graphs. Electronic Journal of Linear Algebra, 20(1):586–594, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/42.

Horn:2005:SLM

[HSSM05] Roger A. Horn, Vladimir V.

Sergeichuk, and Naomi Shaked-Monderer. Solution of linear matrix equations in a *-congruence class. Electronic Journal of Linear Algebra, 13(1):153–156, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/11.

Hogben:2012:EPM

[HW12] Leslie Hogben and Ulrica Wilson. Eventual properties of matrices. Electronic Journal of Linear Algebra, 23 (1):953–965, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/67.

Horn:2010:BSR

[HZ10a] Roger A. Horn and Fuzhen Zhang. Bounds on the spectral radius of a Hadamard product of nonnegative or positive semidefinite matrices. Electronic Journal of Linear Algebra, 20(1):90–94, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/6.

Horn:2010:PSV

[HZ10b] Roger A. Horn and Fuzhen Zhang. Preface: Special volume of the electronic journal of linear algebra on the occasion of the 3rd International Workshop on Matrix Analysis and Applications. Electronic Journal of Linear Algebra, 21(1):1–2, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/1.  Held in

Hangzhou (Lin’An), July 9– 13, 2009.

Huang:2006:RCE

[HZS06] Ting-Zhu Huang, Wei Zhang, and Shu-Qian Shen. Regions containing eigenvalues of a matrix. Electronic Journal of Linear Algebra, 15 (1):215–224, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/17.

Ilisevic:2012:WHS

[IT12] Dijana Ilišević and Néstor Thome. When is the Hermitian/skew-Hermitian part of a matrix a potent matrix? Electronic Journal of Linear Algebra, 24(1):95–112, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/9.

Jargalsaikhan:2013:ICM

[Jar13] Bolor Jargalsaikhan. Indefinite copositive matrices with exactly one positive eigenvalue or exactly one negative eigenvalue. Electronic Journal of Linear Algebra, 26 (1):754–761, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/51.

Johnson:2009:MTP

[JCST09] Charles R. Johnson, Roberto S.

Costas-Santos, and Boris Tadchiev. Matrices totally positive relative to a tree. Electronic Journal of Linear Algebra, 18(1):211–221, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/18.

Johnson:2005:PCC

[JdGM05] Charles R. Johnson and Maria da Graça Marques. Patterns of commutativity: the commutant of the full pattern. Electronic Journal of Linear Algebra, 14(1):43– 50, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol14/iss1/5.

Joshi:2014:SIH

[JG14] Babhru Joshi and Muddappa S. Gowda. Schur ideals and homomorphisms of the semidefinite cone. Electronic Journal of Linear Algebra, 27 (1):865–878, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/1067.

Johnson:2003:AFI

[JHH^{+}03] Charles R. Johnson, Yonatan Harel, Christopher J. Hillar, Jonathan M. Groves, and Patrick X. Rault. Absolutely flat idempotents. Electronic Journal of Linear Algebra, 10 (1):190–200, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/15.

Jin:2005:MCQ

[Jin05] Wen-Long Jin. A method for computing quadratic Brunovsky forms. Electronic Journal of Linear Algebra, 13(1):40–55, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/3.

Johnson:1996:CSM

[JK96] Charles R. Johnson and Brenda K. Kroschel. The combinatorially symmetric P -matrix completion problem. Electronic Journal of Linear Algebra, 1(1):59–63, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol1/iss1/5.

Johnson:2011:CTP

[JK11] Charles R. Johnson and Brenda K. Kroschel. Conditions for a totally positive completion in the case of a symmetrically placed cycle. Electronic Journal of Linear Algebra, 22(1):125–137, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/7.

Jaklic:2013:EDM

[JM13] Gašper Jaklič and Jolanda Modic. On Euclidean distance matrices of graphs. Electronic Journal of Linear Algebra, 26(1):574–589, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/38.

Johnson:2009:TPC

[JN09] Charles R. Johnson and Cris Negron. Totally positive completions for monotonically labeled block clique graphs. Electronic Journal of Linear Algebra, 18 (1):146–161, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/13.

Johnson:2012:OGS

[JPS12] Ch. R. Johnson, J. M. Peña, and T. Szulc. Optimal Geršgorin-style estimation of the singular value. Electronic Journal of Linear Algebra, 25(1):48–59, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol25/iss1/5.

Johnson:2008:CCM

[JR08] Charles R. Johnson and Robert Reams. Constructing copositive matrices from interior matrices. Electronic Journal of Linear Algebra, 17(1):9–20, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/2.

Johnson:2002:EMM

[JS02] Charles R. Johnson and Carlos M. Saiago. Estimation of the maximum multiplicity of an eigenvalue in terms of the vertex degrees of the graph of a matrix. Electronic Journal of Linear Algebra, 9(1):27– 31, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/3.

Jordan:1998:RSC

[JTU98] Cristina Jordán, Juan R.

Torregrosa, and Ana M. Urbano. On the Rodman– Shalom conjecture regarding the Jordan form of completions of partial upper triangular matrices. Electronic Journal of Linear Algebra, 3 (1):103–118, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol3/iss1/10.

Justino:2012:CRA

[JvdB12] Júlia Justino and Imme van den Berg. Cramer’s rule applied to flexible systems of linear equations. Electronic Journal of Linear Algebra, 24 (1):126–152, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol24/iss1/11.

Jedrzejec:2001:PSM

[JW01] Henryk A. Jedrzejec and Zbigniew I. Woźnicki. On properties of some matrix splittings. Electronic Journal of Linear Algebra, 8(1):47– 52, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol8/iss1/4.

Johnson:2012:CPP

[JW12a] Charles R. Johnson and Olivia Walch. Commuting pairs of patterns and symmetric realizations. Electronic Journal of Linear Algebra, 25(1):84–91, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol25/iss1/8.

Johnson:2012:CEO

[JW12b] Charles R. Johnson and Olivia Walch. Critical exponents: old and new. Electronic Journal of Linear Algebra, 25(1):72–83, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol25/iss1/7.

Kalita:2012:CDE

[Kal12] Debajit Kalita. On 3-colored digraphs with exactly one nonsingular cycle. Electronic Journal of Linear Algebra, 23 (1):397–421, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/28.

Karow:2006:ECN

[Kar06] Michael Karow. Eigenvalue condition numbers and a formula of Burke, Lewis and Overton. Electronic Journal of Linear Algebra, 15 (1):143–153, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/10.

Kechriniotis:2012:NPP

[KDTP12] Aristides I. Kechriniotis, Konstantinos K. Delibasis, Christos Tsonos, and Nicholas Petropoulos. A new parallel polynomial division by a separable polynomial via Hermite interpolation with applications. Electronic Journal of Linear Algebra, 23(1):770–781, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/54.

Kirkland:2001:UBA

[Kir01] Stephen J. Kirkland. An upper bound on algebraic connectivity of graphs with many cutpoints. Electronic Journal of Linear Algebra, 8 (1):94–109, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol8/iss1/8.

Kirkland:2003:CPS

[Kir03] Stephen J. Kirkland. Conditioning properties of the stationary distribution for a Markov chain. Electronic Journal of Linear Algebra, 10(1):1–15, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol10/iss1/1.

Kirkland:2004:CAC

[Kir04a] Stephen J. Kirkland. A combinatorial approach to the conditioning of a single entry in the stationary distribution for a Markov chain. Electronic Journal of Linear Algebra, 11(1):168–179, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol11/iss1/14.

Kirkland:2004:GSE

[Kir04b] Stephen J. Kirkland. Girth and subdominant eigenvalues for stochastic matrices. Electronic Journal of Linear Algebra, 12(1):25–41, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol12/iss1/5.

Kirkland:2006:LPN

[Kir06] Stephen J. Kirkland. Limit points for normalized Laplacian eigenvalues. Electronic Journal of Linear Algebra, 15 (1):337–344, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/28.

Kirkland:2007:NDB

[Kir07] Stephen J. Kirkland. A note on a distance bound using eigenvalues of the normalized Laplacian matrix. Electronic Journal of Linear Algebra, 16 (1):204–207, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/18.

Kirkland:2009:SES

[Kir09] Stephen J. Kirkland. Subdominant eigenvalues for stochastic matrices with given column sums. Electronic Journal of Linear Algebra, 18(1):784–800, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/57.

Kirkland:2014:KCS

[Kir14] Steve Kirkland. On the Kemeny constant and stationary distribution vector for a Markov chain. Electronic Journal of Linear Algebra, 27 (1):354–372, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/23.

Kaagstrom:2011:CCG

[KKK11] Bo Kågström, Lars Karlsson, and Daniel Kressner. Computing codimensions and generic canonical forms for generalized matrix products. Electronic Journal of Linear Algebra, 22(1):277–309, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/18.

Kalogeropoulos:2008:DIT

[KKP08] Grigoris I. Kalogeropoulos, Athanasios D. Karageorgos, and Athanasios A. Pantelous. The Drazin inverse through the matrix pencil approach and its application to the study of generalized linear systems with rectangular or square coefficient matrices. Electronic Journal of Linear Algebra, 17(1):118–138, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/11.

Kalogeropoulos:2009:AMT

[KKP09] Grigoris I. Kalogeropoulos, Athanasios D. Karageorgos, and Athanasios A. Pantelous. An angle metric through the notion of Grassmann representative. Electronic Journal of Linear Algebra, 18 (1):108–116, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/10.

Kressner:2013:SCF

[KL13] Daniel Kressner and Xin Liu. Structured canonical forms for products of (skew-) symmetric matrices and the matrix equation XAX = B. Electronic Journal of Linear Algebra, 26(1):215–230, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/17.

Kumar:2014:SPP

[KL14] G. Krishna Kumar and S. H.

Lui. On some properties of the pseudospectral radius. Electronic Journal of Linear Algebra, 27(1):342–353, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/22.

Koo:2015:EBS

[KL15] Yoo Young Koo and Jae Kun Lim. Extension of Bessel sequences to oblique dual frame sequences and the minimal projection. Electronic Journal of Linear Algebra, 30 (1):51–65, 2015. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol30/iss1/3.

Kuzma:2011:NPJ

[KLL^{+}11] Bojan Kuzma, Gorazd Lešnjak, Chi-Kwong Li, Tatjana Petek, and Leiba Rodman. Norm preservers of Jordan products. Electronic Journal of Linear Algebra, 22 (1):959–978, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/62.

Kuzma:2011:TNR

[KLR11] Bojan Kuzma, Chi-Kwong Li, and Leiba Rodman. Tracial numerical ranges and linear dependence of operators. Electronic Journal of Linear Algebra, 22(1):22–52, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/3.

Knudsen:2001:NCR

[KM01] Cindy Knudsen and Judith J. McDonald. A note on the convexity of the realizable set of eigenvalues for nonnegative symmetric matrices. Electronic Journal of Linear Algebra, 8(1):110–114, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol8/iss1/9.

Kunkel:2005:CCS

[KM05] Peter Kunkel and Volker Mehrmann. Characterization of classes of singular linear differential-algebraic equations. Electronic Journal of Linear Algebra, 13 (1):359–386, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol13/iss1/23.

Kunkel:2011:FAL

[KM11] Peter Kunkel and Volker Mehrmann. Formal adjoints of linear DAE operators and their role in optimal control. Electronic Journal of Linear Algebra, 22(1):672–693, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/44.

Kian:2013:ROJ

[KM13] Mohsen Kian and Mohammad Sal Moslehian. Refinements of the operator Jensen–Mercer inequality. Electronic Journal of Linear Algebra, 26(1):742–753, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/50.

Kiani:2014:LCL

[KM14] Dariush Kiani and Maryam Mirzakhah. On the logconcavity of Laplacian characteristic polynomials of graphs. Electronic Journal of Linear Algebra, 27 (1):392–406, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/25.

Kluza:2014:IRO

[KN14] Pawel Kluza and Marek Niezgoda. Inequalities for relative operator entropies. Electronic Journal of Linear Algebra, 27(1):851–864, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/1066.

Kolman:2000:ADF

[Kol00] Oren Kolman. Almost disjoint families: an application to linear algebra. Electronic Journal of Linear Algebra, 7(1):41–52, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol7/iss1/4.

Kim:2009:GPN

[KOS^{+}09] In-Jae Kim, Dale D. Olesky, Bryan L. Shader, Pauline van den Driessche, Hein van der Holst, and Kevin N. Vander Meulen. Generating potentially nilpotent full sign patterns. Electronic Journal of Linear Algebra, 18 (1):162–175, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/14.

Kouachi:2006:EET

[Kou06] Said Kouachi. Eigenvalues and eigenvectors of tridiagonal matrices. Electronic Journal of Linear Algebra, 15 (1):115–133, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/8.

Kirkland:2000:DLE

[KOvdD00] Stephen J. Kirkland, Dale D. Olesky, and Pauline van den Driessche. Digraphs with large exponent. Electronic Journal of Linear Algebra, 7(1):30–40, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol7/iss1/3.

Katsikis:2008:FCM

[KP08] Vasilios N. Katsikis and Dimitrios Pappas. Fast computing of the Moore–Penrose inverse matrix. Electronic Journal of Linear Algebra, 17 (1):637–650, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/44.

Katsikis:2011:RWG

[KP11] Vasilios N. Katsikis and Dimitrios Pappas. The restricted weighted generalized inverse of a matrix. Electronic Journal of Linear Algebra, 22(1): 1156–1167, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/77.

Khalooei:2008:LPL

[KRT08] Fatemeh Khalooei, Mehdi Radjabalipour, and Parisa Torabian. Linear preservers of left matrix majorization. Electronic Journal of Linear Algebra, 17(1):304–315, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/22.

Kim:2008:SPR

[KS08a] In-Jae Kim and Bryan L. Shader. Sign patterns that require a positive or nonnegative left inverse. Electronic Journal of Linear Algebra, 17 (1):209–218, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/16.

Krebs:2008:SJG

[KS08b] Mike Krebs and Anthony Shaheen. On the spectra of Johnson graphs. Electronic Journal of Linear Algebra, 17 (1):154–167, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/13.

Khalooei:2009:SLP

[KS09] Fatemeh Khalooei and Abbas Salemi. The structure of linear preservers of left matrix majorization on R^{P}. Electronic Journal of Linear Algebra, 18(1):88–97, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/8.

Kierzkowski:2011:BNM

[KS11] Jakub Kierzkowski and Alicja Smoktunowicz. Block normal matrices and Gershgorintype discs. Electronic Journal of Linear Algebra, 22(1): 1059–1069, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/69.

Kim:2011:EDS

[KSH11] Byeong Moon Kim, Byung Chul Song, and Woonjae Hwang. Exponents and diameters of strong products of digraphs. Electronic Journal of Linear Algebra, 22(1): 1106–1111, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/72.

Kirkland:2005:PDC

[KSvdD05] Stephen J. Kirkland, Bryan L.

Shader, and Pauline van den Driessche. Preface [Directions in Combinatorial Matrix Theory Workshop]. Electronic Journal of Linear Algebra, 14(1):1, 2005. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol14/iss1/1.  Held in

Banff, AB, May 7–8, 2004.

Key:2004:EMP

[KV04] Eric S. Key and Hans Volkmer. Eigenvalue multiplicities of products of companion matrices. Electronic Journal of Linear Algebra, 11 (1):103–114, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/10.

Key:2014:NSR

[KV14] Eric Key and Hans Volkmer. A note on the spectral radius of a product of companion matrices. Electronic Journal of Linear Algebra, 27 (1):879–881, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/1068.

Li:2002:CRC

[LA02] Xiezhang Li and Fangjun Arroyo. The convergence rate of the Chebyshev semiiterative method under a perturbation of the foci of an elliptic domain. Electronic Journal of Linear Algebra, 9(1):55–66, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol9/iss1/6.

Laffey:2004:PNR

[Laf04] Thomas J. Laffey. Perturbing non-real eigenvalues of nonnegative real matrices. Electronic Journal of Linear Algebra, 12(1):73–76, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol12/iss1/8.

Lancaster:1999:SSG

[Lan99] Peter Lancaster. Strongly stable gyroscopic systems. Electronic Journal of Linear Algebra, 5(1):53–66, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol5/iss1/6.

Lancaster:2008:LRM

[Lan08] Peter Lancaster. Linearization of regular matrix polynomials. Electronic Journal of Linear Algebra, 17(1):21– 27, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/3.

Liang:2011:SCM

[LD11] Mao-Lin Liang and Li-Fang Dai. The solvability conditions of matrix equations with K-involution. Electronic Journal of Linear Algebra, 22(1):1138–1147, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/75.

Liu:2010:GSC

[LH10] Jianzhou Liu and Rong Huang. Generalized Schur complements of matrices and compound matrices. Electronic Journal of Linear Algebra, 21(1):12–24, 2010.

CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/3.

Liu:2009:CGP

[LHC09] Bolian Liu, Yufei Huang, and Siyuan Chen. On the characterization of graphs with pendent vertices and given nullity. Electronic Journal of Linear Algebra, 18 (1):719–734, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/52.

Liu:2011:SCS

[LHL11] Fenjin Liu, Qiongxiang Huang, and Qinghai Liu. Spectral characterization of †-shape trees. Electronic Journal of Linear Algebra, 22 (1):822–837, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/55.

Li:2014:SOU

[LHL14] Chaoqian Li, Baohua Hu, and Yaotang Li. Simplifications of the Ostrowski upper bounds for the spectral radius of nonnegative matrices. Electronic Journal of Linear Algebra, 27(1):237–249, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/14.

Liu:2009:ENL

[LhLY09] Bolian Liu, Mu huo Liu, and Zhifu You. Erratum to “A note on the largest eigenvalue of non-regular graphs” [MR2390359]. Electronic Journal of Linear Algebra, 18(1):64–68, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/6. See

[LL08b].

Liu:2009:ENR

[LHY09] Bolian Liu, Yufei Huang, and Zhifu You. On λ_{1}- extremal non-regular graphs. Electronic Journal of Linear Algebra, 18(1):735–744, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/53.

Lim:2012:APT

[Lim12] Ming-Huat Lim. Additive preservers of tensor product of rank one Hermitian matrices. Electronic Journal of Linear Algebra, 23 (1):356–374, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/25.

Lin:2013:HRS

[Lin13a] Kai-Liang Lin. Hurwitz– Radon’s symplectic analogy and Hua’s cyclic recurrence relation. Electronic Journal of Linear Algebra, 26 (1):858–872, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/58.

Lin:2013:NAT

[Lin13b] Minghua Lin. Notes on an Anderson–Taylor type inequality. Electronic Journal of Linear Algebra, 26(1):63– 70, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/5.

Lin:2014:DIP

[Lin14] Minghua Lin. A determinantal inequality for positive semidefinite matrices. Electronic Journal of Linear Algebra, 27(1):821–826, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/1063.

Liu:2012:SLS

[Liu12] Muhuo Liu. The (signless) Laplacian spectral radii of C-cyclic graphs with N vertices and K pendant vertices. Electronic Journal of Linear Algebra, 23(1):942–952, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/66.

Li:2008:MET

[LL08a] Shuchao Li and Nana Li.

On minimal energies of trees with given diameter. Electronic Journal of Linear Algebra, 17(1):414–425, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/29.

Liu:2008:NLE

[LL08b] Bolian Liu and Gang Li. A note on the largest eigenvalue of non-regular graphs. Electronic Journal of Linear Algebra, 17(1):54–61, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/6. See erratum

[LhLY09].

Li:2011:GBL

[LL11] Chaoqian Li and Yaotang Li.

Generalizations of Brauer’s localization theorem. Electronic Journal of Linear Algebra, 22(1):1168–1178, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/78.

Liu:2012:SRS

[LL12] Muhuo Liu and Bolian Liu.

Some results on the spectral radii of trees, unicyclic, and bicyclic graphs. Electronic Journal of Linear Algebra, 23 (1):327–339, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/23.

Lin:2012:SPN

[LLMZ12] Zhiping Lin, Yongzhi Liu, Giuseppe Molteni, and Dongye Zhang. Spectral properties for a new composition of a matrix and a complex representation. Electronic Journal of Linear Algebra, 23 (1):530–539, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/37.

Lin:2014:SRL

[LLS14] Huiqiu Lin, Ruifang Liu, and Jinlong Shu. Some results on the largest and least eigenvalues of graphs. Electronic Journal of Linear Algebra, 27 (1):670–682, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/259.

Liu:2011:GDT

[LLW11] Muhuo Liu, Bolian Liu, and Fuyi Wei. Graphs determined by their (signless) Laplacian spectra. Electronic Journal of Linear Algebra, 22 (1):112–124, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/6.

Li:2011:SNL

[LLYL11] Yaotang Li, Xin Liu, Xiaoying Yang, and Chaoqian Li. Some new lower bounds for the minimum eigenvalue of the Hadamard product of an M-matrix and its inverse. Electronic Journal of Linear Algebra, 22(1):630–643, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/41.

Li:2013:EGT

[LLZ13] Shuchao Li, Yan Li, and Xixi Zhang. Edge-grafting theorems on permanents of the Laplacian matrices of graphs and their applications. Electronic Journal of Linear Algebra, 26(1):28–48, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/3.

Laffey:1998:RIJ

[LM98] Thomas J. Laffey and Eleanor Meehan. A refinement of an inequality of Johnson, Loewy and London on nonnegative matrices and some applications. Electronic Journal of Linear Algebra, 3 (1):119–128, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol3/iss1/11.

Li:2010:CES

[LM10] Aihua Li and Edward Mosteig. On the construction of explicit solutions to the matrix equation X^{2}AX = AXA. Electronic Journal of Linear Algebra, 21(1):142–153, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol21/iss1/13.

Livshits:2011:MDM

[LMR11] L. Livshits, G. Macdonald, and H. Radjavi. Multiplicative diagonals of matrix semigroups. Electronic Journal of Linear Algebra, 22 (1):252–266, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/16.

Li:2004:NEF

[LMS04] Chi-Kwong Li, Thomas Milligan, and Bryan L. Shader. Non-existence of 5 × 5 full ray nonsingular matrices. Electronic Journal of Linear Algebra, 11(1):212–240, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol11/iss1/18.

Li:2010:ECE

[LMT10] Chi-Kwong Li, Thomas Milligan, and Michael W. Trosset. Euclidean and circum-Euclidean distance matrices:

Characterizations and linear preservers. Electronic Journal of Linear Algebra, 20 (1):739–752, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/53.

Li:2014:MVR

[LNP14] Xiaowei Li, Michael Nathanson, and Rachel Phillips. Minimum vector rank and complement critical graphs. Electronic Journal of Linear Algebra, 27(1):100–123, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/5.

Loi:2012:MBV

[Loi12] Le Cong Loi. On multipoint boundary value problems for index-2 linear singular difference equations. Electronic Journal of Linear Algebra, 23 (1):88–114, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/6.

Lotker:2007:NDV

[Lot07] Zvi Lotker. Note on deleting a vertex and weak interlacing of the Laplacian spectrum. Electronic Journal of Linear Algebra, 16(1):68–72, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/6.

Li:2007:MPS

[LPR07] Chi-Kwong Li, Edward Poon, and Ashwin Rastogi. Maps preserving spectral radius, numerical radius, spectral norm. Electronic Journal of Linear Algebra, 16 (1):347–365, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol16/iss1/30.

Li:2014:RES

[LPW14] Chi-Kwong Li, Yiu-Tung Poon, and Xuefeng Wang. Ranks and eigenvalues of states with prescribe reduced states. Electronic Journal of Linear Algebra, 27(1): 935–950, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/1074.

Li:1998:SCG

[LR98a] Chi-Kwong Li and Leiba Rodman. Shapes and computer generation of numerical ranges of Krein space operators. Electronic Journal of Linear Algebra, 3(1):31–47, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol3/iss1/6.

Loewy:1998:SMB

[LR98b] Raphael Loewy and Nizar Radwan. On spaces of matrices with a bounded number of eigenvalues. Electronic Journal of Linear Algebra, 3 (1):142–152, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol3/iss1/13.

Li:2002:NTS

[LR02] Aihua Li and Duane Randall. Non-trivial solutions to certain matrix equations. Electronic Journal of Linear Algebra, 9(1):282–289, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/24.

Lal:2013:NSC

[LR13] A. K. Lal and A. Satyanarayana Reddy. Nonsingular circulant graphs and digraphs. Electronic Journal of Linear Algebra, 26 (1):248–257, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/19.

Lee:1998:SCS

[LS98] Gwang-Yeon Lee and Bryan L.

Shader. Sign-consistency and solvability of constrained linear systems. Electronic Journal of Linear Algebra, 4(1):1– 18, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol4/iss1/1.

Laffey:2008:CEN

[LŠ08] Thomas J. Laffey and Helena Šmigoc. On a classic example in the nonnegative inverse eigenvalue problem. Electronic Journal of Linear Algebra, 17(1):333–342, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/24.

Lippert:2009:JF

[LS09] Ross A. Lippert and Gilbert Strang. The Jordan forms of AB and BA. Electronic Journal of Linear Algebra, 18 (1):281–288, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/25.

Lee:2013:NIO

[LS13] Moon Ho Lee and Ferenc Szöllősi. A note on inverseorthogonal Toeplitz matrices. Electronic Journal of Linear Algebra, 26(1):898–904, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol26/iss1/60.

Lebtahi:2012:PMG

[LT12] Leila Lebtahi and Néstor Thome. Properties of a matrix group associated to a {K, S + 1}-potent matrix. Electronic Journal of Linear Algebra, 24(1):34–44, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/5.

Li:1996:NRO

[LTU96] Chi-Kwong Li, Nam-Kiu Tsing, and Frank Uhlig. Numerical ranges of an operator on an indefinite inner product space. Electronic Journal of Linear Algebra, 1(1):1–17, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol1/iss1/1.

Lin:2014:PSX

[LvdD14] Minghua Lin and Pauline van den Driessche. Positive semidefinite 3 × 3 block matrices. Electronic Journal of Linear Algebra, 27 (1):827–836, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/1064.

Liu:2011:GIL

[LWB11] Xiaoji Liu, Lingling Wu, and Julio Benítez. On the group inverse of linear combinations of two group invertible matrices. Electronic Journal of Linear Algebra, 22 (1):490–503, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/30.

Liu:2013:LES

[LWYJ13] Ruifang Liu, Haixia Wan, Jinjiang Yuan, and Huicai Jia. The least eigenvalue of the signless Laplacian of nonbipartite unicyclic graphs with K pendant vertices. Electronic Journal of Linear Algebra, 26(1):333–344, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol26/iss1/22.

Lei:2004:MIM

[LWZ04] Tian-Gang Lei, Ching-Wah Woo, and Fuzhen Zhang. Matrix inequalities by means of embedding. Electronic Journal of Linear Algebra, 11(1):66–77, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/7.

Li:2011:AMG

[LX11] Yanbo Li and Zhankui Xiao. Additivity of maps on generalized matrix algebras. Electronic Journal of Linear Algebra, 22(1):743–757, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/49.

Luo:2014:ERS

[LX14] Ziyan Luo and Naihua Xiu.

Exact relaxation for the semidefinite matrix rank minimization problem with extended Lyapunov equation constraint. Electronic Journal of Linear Algebra, 27(1):683–691, September 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/260.

Liu:2011:NRO

[LY11] Xifu Liu and Hu Yang. A note on the reverse order laws for {1, 2, 3}- and {1, 2, 4}- inverses of multiple matrix products. Electronic Journal of Linear Algebra, 22 (1):620–629, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/40.

Li:2012:OTM

[LZ12] Shuchao Li and Qin Zhao.

Ordering trees by the minimal entries of their doubly stochastic graph matrices. Electronic Journal of Linear Algebra, 23(1):295–305, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/21.

Liu:2014:SCP

[LZ14] Xiaogang Liu and Sanming Zhou. Spectral characterizations of propeller graphs. Electronic Journal of Linear Algebra, 27(1):19–38, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/2.

Liu:2012:SPS

[LZFR12] Zhongyun Liu, Yulin Zhang, Carla Ferreira, and Rui Ralha. Structure-preserving Schur methods for computing square roots of real skew-Hamiltonian matrices. Electronic Journal of Linear Algebra, 23(1):845–865, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/60.

Liu:2010:NUB

[LZL10] Jianzhou Liu, Juan Zhang, and Yu Liu. A new upper bound for the eigenvalues of the continuous algebraic Riccati equation. Electronic Journal of Linear Algebra, 20 (1):314–321, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/23.

Li:2013:SMG

[LZZ13] Shuchao Li, Huihui Zhang, and Minjie Zhang. On the spectral moment of graphs with K cut edges. Electronic Journal of Linear Algebra, 26 (1):718–731, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/48.

MarquesdeSa:2006:SSB

[Mar06] Eduardo Marques de Sá.

Some subpolytopes of the Birkhoff polytope. Electronic Journal of Linear Algebra, 15(1):1–7, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/1.

Maroulas:2008:FRC

[Mar08] John Maroulas. Further results on the Craig–Sakamoto equation. Electronic Journal of Linear Algebra, 17 (1):445–457, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/32.

Martin:2014:GTA

[Mar14] Antonio J. Calderón Martín.

Graded triangular algebras. Electronic Journal of Linear Algebra, 27(1):317–331, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/20.

Matsuura:2003:GMP

[Mat03] Masaya Matsuura. A generalization of Moore–Penrose biorthogonal systems. Electronic Journal of Linear Algebra, 10(1):146–154, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/11.

Matejas:2008:QCB

[Mat08] Josip Matejaš. Quadratic convergence bounds of scaled iterates by the serial Jacobi methods for indefinite Hermitian matrices. Electronic Journal of Linear Algebra, 17(1):62–87, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/7.

Mazko:2010:SLR

[Maz10] A. G. Mazko. Spectrum localization of regular matrix polynomials and functions. Electronic Journal of Linear Algebra, 20(1):333–350, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/25.

Maden:2012:GCI

[MÇ12] A. Dilek Maden and A. Sinan Çevik. A generalization for the clique and independence numbers. Electronic Journal of Linear Algebra, 23 (1):164–170, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/11.

Maden:2013:MBS

[MC13] A. Dilek Maden and A. Sinan Cevik. Majorization bounds for signless Laplacian eigenvalues. Electronic Journal of Linear Algebra, 26 (1):781–794, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/53.

Mosic:2009:PIE

[MD09] Dijana Mosić and Dragan S.

Djordjević. Partial isometries and EP elements in rings with involution. Electronic Journal of Linear Algebra, 18 (1):761–772, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/55.

Mosic:2011:ROL

[MD11a] Dijana Mosić and Dragan S.

Djordjević. Reverse order law for the Moore–Penrose inverse in C^{∗}-algebras. Electronic Journal of Linear Algebra, 22(1):92–111, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/5.

Mosic:2011:WEE

[MD11b] Dijana Mosić and Dragan S.

Djordjević. Weighted EP elements in C^{∗}-algebras. Electronic Journal of Linear Algebra, 22(1):912–930, 2011.

CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/60.

Mosic:2014:IKI

[MDKM14] Dijana Mosić, Dragan S. Djordjević, and Gabriel Kantún-Montiel. Imagekernel (P, Q)-inverses in rings. Electronic Journal of Linear Algebra, 27(1): 272–283, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/17.

Mecheri:2012:BPS

[Mec12]  Salah Mecheri.  Bishop’s

property (β), SVEP and Dunford property (C). Electronic Journal of Linear Algebra, 23(1):523–529, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/36.

Mehl:2006:EDP

[Meh06a] Christian Mehl. Essential decomposition of polynomially normal matrices in real indefinite inner product spaces. Electronic Journal of Linear Algebra, 15(1):84–106, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/6.

Mehl:2006:CNM

[Meh06b] Christian Mehl. On classification of normal matrices in indefinite inner product spaces. Electronic Journal of Linear Algebra, 15(1):50–83, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/5.

Mani:1996:SPA

[MH96] Vaidyanath Mani and Robert E.

Hartwig. Some properties of the q-adic Vandermonde matrix. Electronic Journal of Linear Algebra, 1(1):18–33, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol1/iss1/2.

Mirzakhah:2010:SGD

[MK10] Maryam Mirzakhah and Dariush Kiani. The sun graph is determined by its signless Laplacian spectrum. Electronic Journal of Linear Algebra, 20(1):610–620, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/44.

Merikoski:2013:UBL

[MKR13] Jorma K. Merikoski, Ravinder Kumar, and Ram Asrey Rajput. Upper bounds for the largest eigenvalue of a bipartite graph. Electronic Journal of Linear Algebra, 26 (1):168–176, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/13.

Masque:2012:RIS

[MM12] J. Muñoz Masqué and M. Eugenia Rosado María. Rational invariants on the space of all structures of algebras on a two-dimensional vector space. Electronic Journal of Linear Algebra, 23 (1):483–507, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/33.

Marco:2013:ACT

[MM13] Ana Marco and José-Javier Martínez. Accurate computations with totally positive Bernstein–Vandermonde matrices. Electronic Journal of Linear Algebra, 26 (1):357–380, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/24.

Mackey:2005:SPA

[MMD05] D. Steven Mackey, Niloufer Mackey, and Daniel M. Dunlavy. Structure preserving algorithms for perplectic eigenproblems. Electronic Journal of Linear Algebra, 13(1):10–39, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/2.

Mackey:2011:SFP

[MMMM11] D. Steven Mackey, Niloufer Mackey, Christian Mehl, and Volker Mehrmann. Smith forms of palindromic matrix polynomials. Electronic Journal of Linear Algebra, 22(1):53–91, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/4.

Mackey:2003:STS

[MMT03] D. Steven Mackey, Niloufer Mackey, and Françoise Tisseur. Structured tools for structured matrices. Electronic Journal of Linear Algebra, 10(1):106–145, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/10.

Moghaddamfar:2011:NFI

[MMT11] Ali Reza Moghaddamfar, Kambiz Moghaddamfar, and Hadiseh Tajbakhsh. New families of integer matrices whose leading principal minors form some well-known sequences. Electronic Journal of Linear Algebra, 22 (1):598–619, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/39.

Mazilu:2012:ATM

[MMW12] Irina Mazilu, Dan A. Mazilu, and H. Thomas Williams. Applications of tridiagonal matrices in non-equilibrium statistical physics. Electronic Journal of Linear Algebra, 24(1):7–17, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol24/iss1/3.

Mehl:2000:CFD

[MMX00] Christian Mehl, Volker Mehrmann, and Hongguo Xu. Canonical forms for doubly structured matrices and pencils. Electronic Journal of Linear Algebra, 7(1):112–151, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol7/iss1/10.

Molitierno:2001:ACT

[MN01] Jason J. Molitierno and Michael Neumann. The algebraic connectivity of two trees connected by an edge of infinite weight. Electronic Journal of Linear Algebra, 8(1):1–13, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol8/iss1/1.

Molnar:2010:TIP

[MN10] Lajos Molnár and Gergő Nagy. Thompson isometries on positive operators: The 2dimensional case. Electronic Journal of Linear Algebra, 20(1):79–89, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/5.

Marks:2003:MNO

[MNP03] Michael Marks, Rick Norwood, and George Poole. The maximum number of 2 × 2 odd submatrices in (0, 1)- matrices. Electronic Journal of Linear Algebra, 10 (1):223–231, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/18.

Mitchell:2010:MVR

[MNR10] Lon H. Mitchell, Sivaram K.

Narayan, and Ian Rogers. On the minimum vector rank of multigraphs. Electronic Journal of Linear Algebra, 20 (1):661–672, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/47.

Molitierno:2000:TBA

[MNS00] Jason J. Molitierno, Michael Neumann, and Bryan L. Shader. Tight bounds on the algebraic connectivity of a balanced binary tree. Electronic Journal of Linear Algebra, 6(1):62–71, 2000. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol6/iss1/7.

Moslehian:2011:DMT

[MNS11] Mohammad Sal Moslehian, Ritsuo Nakamoto, and Yuki Seo. A Diaz–Metcalf type inequality for positive linear maps and its applications. Electronic Journal of Linear Algebra, 22(1):179–190, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/11.

Molnar:2009:MPO

[Mol09] Lajos Molnár. Maps on positive operators preserving Lebesgue decompositions. Electronic Journal of Linear Algebra, 18(1):222–232, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/19.

Molnar:2011:MPG

[Mol11] Lajos Molnár. Maps preserving general means of positive operators. Electronic Journal of Linear Algebra, 22 (1):864–874, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/58.

Morris:2003:RHP

[Mor03] Walter D. Morris, Jr. Recognition of hidden positive row diagonally dominant matrices. Electronic Journal of Linear Algebra, 10(1): 102–105, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/9.

Morillas:2011:GRS

[Mor11] Patricia Mariela Morillas.

Group reconstruction systems. Electronic Journal of Linear Algebra, 22(1): 875–911, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/59.

Morillas:2013:HRS

[Mor13] Patricia Mariela Morillas.

Harmonic reconstruction systems. Electronic Journal of Linear Algebra, 26(1): 692–705, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/46.

McDonald:1998:P

[MOS^{+}98] Judith J. McDonald, Dale Dale Olesky, Hans Schneider, Michael J. Tsatsomeros, and Pauline van den Driessche. Z-pencils. Electronic Journal of Linear Algebra, 4(1):32– 38, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol4/iss1/3.

Mond:2000:IIH

[MP00] Bertram Mond and Josip E. Pecaric. On inequalities involving the Hadamard product of matrices. Electronic Journal of Linear Algebra, 6(1):56–61, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol6/iss1/6.

Moghaddamfar:2009:GPT

[MP09a] Ali Reza Moghaddamfar and S. M. H. Pooya. Generalized Pascal triangles and Toeplitz matrices. Electronic Journal of Linear Algebra, 18 (1):564–588, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/43.

Muhic:2009:STP

[MP09b] Andrej Muhič and Bor Plestenjak. On the singular two-parameter eigenvalue problem. Electronic Journal of Linear Algebra, 18 (1):420–437, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/34.

Moghaddamfar:2007:MCD

[MPSS07] A. R. Moghaddamfar, S. M. H. Pooya, S. Navid Salehy, and S. Nima Salehy. More calculations on determinant evaluations. Electronic Journal of Linear Algebra, 16(1):19–29, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/2.

Maroulas:2000:SCP

[MPT00] John Maroulas, Panayiotis J.

Psarrakos, and Michael J. Tsatsomeros. Separable characteristic polynomials of pencils and property L. Electronic Journal of Linear Algebra, 7(1):182–190, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol7/iss1/14.

Marrero:2012:SMR

[MR12] J. Abderramán Marrero and M. Rachidi. The Szegő matrix recurrence and its associated linear non-autonomous area-preserving map. Electronic Journal of Linear Algebra, 24(1):168–180, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/13.

Mehl:2004:HMS

[MRR04] Christian Mehl, André C. M.

Ran, and Leiba Rodman. Hyponormal matrices and semidefinite invariant subspaces in indefinite inner products. Electronic Journal of Linear Algebra, 11 (1):192–204, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/16.

Mason:2004:GCC

[MS04] Oliver Mason and Robert Shorten. The geometry of convex cones associated with the Lyapunov inequality and the common Lyapunov function problem. Electronic Journal of Linear Algebra, 12(1):42–63, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol12/iss1/6.

Majstorovic:2014:NGW

[MS14] Snježana Majstorović and Dragan Stevanović. A note on graphs whose largest eigenvalue of the modularity matrix equals zero. Electronic Journal of Linear Algebra, 27(1):611–618, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol27/iss1/256.

Moghaddamfar:2008:CMR

[MSS08] Ali Reza Moghaddamfar, S. Navid Salehy, and S. Nima Salehy. Certain matrices related to the Fibonacci sequence having recursive entries. Electronic Journal of Linear Algebra, 17(1): 543–576, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/40.

McDonald:2009:PNM

[MST09] Judith J. McDonald, Hans Schneider, and Michael J. Tsatsomeros. Preface: Nonnegative Matrix Theory: Generalizations and Applications American Institute of Mathematics, Palo Alto, CA, December 1–5, 2008. Electronic Journal of Linear Algebra, 19(1):1, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol19/iss1/1.

Monov:2004:RDS

[MT04] Vladimir Monov and Michael Tsatsomeros. On reducing and deflating subspaces of matrices. Electronic Journal of Linear Algebra, 11 (1):246–257, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol11/iss1/20.

McCarthy:2010:PIN

[MT10] John E. McCarthy and Richard M. Timoney. Polynomial inequalities for noncommuting operators. Electronic Journal of Linear Algebra, 20(1):506–518, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/37.

Miklavic:2011:LMH

[MT11] Štefko Miklavič and Paul Terwilliger. The A-like matrices for a hypercube. Electronic Journal of Linear Algebra, 22(1):796–809, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/53.

Macias-Virgos:2009:LES

[MVPS09] E. Macías-Virgós and M. J. Pereira-Sáez. Left eigenvalues of 2 × 2 symplectic matrices. Electronic Journal of Linear Algebra, 18 (1):274–280, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/24.

Mehrmann:1999:SJC

[MX99] Volker Mehrmann and Hongguo Xu. Structured Jordan canonical forms for structured matrices that are Hermitian, skew Hermitian or unitary with respect to indefinite inner products. Electronic Journal of Linear Algebra, 5(1):67–103, 1999. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol5/iss1/7.

Mehrmann:2008:PPI

[MX08] Volker Mehrmann and Hongguo Xu. Perturbation of purely imaginary eigenvalues of Hamiltonian matrices under structured perturbations. Electronic Journal of Linear Algebra, 17(1):234–257, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/18.

Nabben:2000:CIM

[Nab00] Reinhard Nabben. A class of inverse M-matrices. Electronic Journal of Linear Algebra, 7(1):53–58, 2000. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol7/iss1/5.

Nakazato:2011:NRL

[NBdP11] Hiroshi Nakazato, Natália Bebiano, and João da Providência. The numerical range of linear operators on the 2dimensional Krein space. Electronic Journal of Linear Algebra, 22(1):430–442, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/25.

Nakazato:2014:NCI

[NBdP14] Hiroshi Nakazato, Natalia Bebiano, and Joao da Providencia. A note on the convexity of the indefinite joint numerical range. Electronic Journal of Linear Algebra, 27 (1):705–715, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/262.

Niezgoda:2006:UBC

[Nie06]  Marek Niezgoda.  Upper

bounds on certain functionals defined on groups of linear operators. Electronic Journal of Linear Algebra, 15 (1):191–200, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/14.

Niezgoda:2010:KTI

[Nie10]  Marek Niezgoda.  Kan-

torovich type inequalities for ordered linear spaces. Electronic Journal of Linear Algebra, 20(1):103–114, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/8.

Niezgoda:2013:CDJ

[Nie13]  Marek Niezgoda.  Choi–

Davis–Jensen’s inequality and generalized inverses of linear operators. Electronic Journal of Linear Algebra, 26 (1):406–416, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/27.

Nikiforov:2006:LCG

[Nik06] Vladimir Nikiforov. Linear combinations of graph eigenvalues. Electronic Journal of Linear Algebra, 15 (1):329–336, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/27.

Nikiforov:2014:MID

[Nik14] Vladimir Nikiforov. Maxima of the Q-index: degenerate graphs. Electronic Journal of Linear Algebra, 27 (1):250–257, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/15.

Naqvi:2002:CSE

[NM02] Sarah Carnochan Naqvi and Judith J. McDonald. The combinatorial structure of eventually nonnegative matrices. Electronic Journal of Linear Algebra, 9(1): 255–269, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol9/iss1/21.

Neumann:2003:BGE

[NO03] Michael Neumann and Nic Ormes. Bounds for graph expansions via elasticity. Electronic Journal of Linear Algebra, 10(1):163–178, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/13.

Noferini:2012:BCE

[Nof12] Vanni Noferini. The behaviour of the complete eigenstructure of a polynomial matrix under a generic rational transformation. Electronic Journal of Linear Algebra, 23(1): 607–624, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/44.

Niroomand:2011:RSM

[NR11] Peyman Niroomand and Francesco G. Russo. A restriction on the Schur multiplier of nilpotent Lie algebras. Electronic Journal of Linear Algebra, 22(1):1–9, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/1.

Nechepurenko:2012:GMI

[NS12] Yuri M. Nechepurenko and Miloud Sadkane. A generalization of matrix inversion with application to linear differential-algebraic systems. Electronic Journal of Linear Algebra, 23(1): 831–844, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/59.

Nomura:2007:ATL

[NT07] Kazumasa Nomura and Paul Terwilliger. Affine transformations of a Leonard pair. Electronic Journal of Linear Algebra, 16(1):389–417, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/33.

Neubauer:1998:OWD

[NWZ98] Michael G. Neubauer, William Watkins, and Joel Zeitlin. Doptimal weighing designs for four and five objects. Electronic Journal of Linear Algebra, 4(1):48–73, 1998. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol4/iss1/5.

Neumann:2005:PAC

[NX05] Michael Neumann and Jianhong Xu. A parallel algorithm for computing the group inverse via Perron complementation. Electronic Journal of Linear Algebra, 13 (1):131–145, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/9.

Neumann:2006:NNN

[NX06] Michael Neumann and Jianhong Xu. A note on Newton and Newton-like inequalities for M-matrices and for Drazin inverses of Mmatrices. Electronic Journal of Linear Algebra, 15 (1):314–328, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/26.

Nikiforov:2014:MIG

[NY14] Vladimir Nikiforov and Xiying Yuan. Maxima of the Q-index: graphs without long paths. Electronic Journal of Linear Algebra, 27 (1):504–514, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/32.

Oepomo:2003:CCE

[Oep03] Tedja Santanoe Oepomo. A contribution to Collatz’s eigenvalue inclusion theorem for nonnegative irreducible matrices. Electronic Journal of Linear Algebra, 10(1):31– 45, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol10/iss1/3.

Dokovic:2002:PAC

[D̄ok02] Dragomir Ž. D̄oković. Proof of Atiyah’s conjecture for two special types of configurations. Electronic Journal of Linear Algebra, 9(1): 132–137, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/14.

Orel:2009:ARO

[OK09] Marko Orel and Bojan Kuzma. Additive rank-one nonincreasing maps on Hermitian matrices over the field GF(2^{2}). Electronic Journal of Linear Algebra, 18 (1):482–499, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/38.

Olesky:2009:MGM

[OTvdD09] Dale D. Olesky, Michael J.

Tsatsomeros, and Pauline van den Driessche. M_{∨}- matrices: a generalization of M-matrices based on eventually nonnegative matrices. Electronic Journal of Linear Algebra, 18(1):339–351, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/29.

Omidi:2010:STM

[OV10] Gholam R. Omidi and Ebrahim Vatandoost. Starlike trees with maximum degree 4 are determined by their signless Laplacian spectra. Electronic Journal of Linear Algebra, 20 (1):274–290, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/20.

Ovall:2004:AGL

[Ova04] Jeffrey S. Ovall. An analysis of GCD and LCM matrices via the LDL^{T} -factorization. Electronic Journal of Linear Algebra, 11(1):51–58, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol11/iss1/5.

Pati:2001:TSE

[Pat01] Sukanta Pati. The third smallest eigenvalue of the Laplacian matrix. Electronic Journal of Linear Algebra, 8 (1):128–139, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol8/iss1/11.

Pekarev:2014:NCS

[Pek14] Edward L. Pekarev. A note on characterization of the shorted operation. Electronic Journal of Linear Algebra, 27 (1):155–161, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/8.

Peretz:2002:ICM

[Per02] Ronen Peretz. Iterations of concave maps, the Perron– Frobenius theory, and applications to circle packings. Electronic Journal of Linear Algebra, 9(1):197–254, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/20.

Pereira:2007:NMS

[Per07] Rajesh J. Pereira. Nilpotent matrices and spectrally arbitrary sign patterns. Electronic Journal of Linear Algebra, 16(1):232–236, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/21.

Pickmann:2009:TIE

[PES09] Hubert Pickmann, Juan C.

Egaña, and Ricardo L. Soto. Two inverse eigenproblems for symmetric doubly arrow matrices. Electronic Journal of Linear Algebra, 18 (1):700–718, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/51.

Pickmann:2011:ESR

[PES11] Hubert Pickmann, Juan C.

Egaña, and Ricardo L. Soto. Extreme spectra realization by real symmetric tridiagonal and real symmetric arrow matrices. Electronic Journal of Linear Algebra, 22 (1):780–795, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/52.

Peters:2012:PSM

[Pet12]  Travis Peters.  Positive

semidefinite maximum nullity and zero forcing number. Electronic Journal of Linear Algebra, 23(1):815–830, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/58.

Pinkus:2004:IM

[Pin04] Allan Pinkus. Interpolation by matrices. Electronic Journal of Linear Algebra, 11 (1):281–291, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol11/iss1/22.

Pinkus:2008:ZMT

[Pin08] Allan Pinkus. Zero minors of totally positive matrices. Electronic Journal of Linear Algebra, 17(1):532–542, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/39.

Pan:2012:IML

[PL12] Huawei Pan and Yuan Lei. Iterative method for the least squares problem of a matrix equation with tridiagonal matrix constraint. Electronic Journal of Linear Algebra, 23(1):1001–1022, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/70.

Pan:2012:PAS

[PLBG12] Yu Pan, Qing Liu, Chengming Bai, and Li Guo. PostLie algebra structures on the Lie algebra SL(2, C). Electronic Journal of Linear Algebra, 23(1):180–197, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/13.

Perinan:2012:GMH

[PM12] María J. Aragón Periñán and Antonio J. Calderón Martín. On graded matrix Homalgebras. Electronic Journal of Linear Algebra, 24(1):45– 65, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol24/iss1/6.

Poloni:2010:CMG

[Pol10] Federico G. Poloni. Constructing matrix geometric means. Electronic Journal of Linear Algebra, 20 (1):419–435, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/32.

Papathanasiou:2008:NMP

[PP08] Nikolaos Papathanasiou and Panayiotis Psarrakos. Normal matrix polynomials with nonsingular leading coefficients. Electronic Journal of Linear Algebra, 17(1): 458–472, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/33.

Pena:2011:LMO

[PS11] J. M. Peña and T. Szulc. Loewner matrix ordering in estimation of the smallest singular value. Electronic Journal of Linear Algebra, 22 (1):466–479, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/28.

Petkovic:2013:BRC

[PS13] Marko D. Petković and Predrag S. Stanimirović. Block recursive computation of generalized inverses. Electronic Journal of Linear Algebra, 26 (1):394–405, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/26.

Psarrakos:2002:RCM

[Psa02] Panayiotis J. Psarrakos. On the m-th roots of a complex matrix. Electronic Journal of Linear Algebra, 9(1):32–41, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/4.

Psarrakos:2000:RBN

[PT00] Panayiotis Psarrakos and Michael J. Tsatsomeros. On the relation between the numerical range and the joint numerical range of matrix polynomials. Electronic Journal of Linear Algebra, 6(1):20–30, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol6/iss1/4.

Paksoy:2014:IGM

[PTZ14] Vehbi E. Paksoy, Ramazan Turkmen, and Fuzhen Zhang. Inequalities of generalized matrix functions via tensor products. Electronic Journal of Linear Algebra, 27 (1):332–341, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/21.

Pultarova:2012:EPF

[Pul12] Ivana Pultarová. Error propagation formula of multilevel iterative aggregationdisaggregation methods for non-symmetric problems. Electronic Journal of Linear Algebra, 25(1):9–21, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol25/iss1/3.

Qi:2013:CLL

[Qi13] Xiaofei Qi. Characterizing Lie (ξ-Lie) derivations on triangular algebras by local actions. Electronic Journal of Linear Algebra, 26 (1):816–835, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/55.

Rakic:2012:DRI

[Rak12] Dragan S. Rakić. Decomposition of a ring induced by minus partial order. Electronic Journal of Linear Algebra, 23(1):1040–1059, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/72.

Radojevic:2012:QSQ

[RD12] Ivana M. Radojević and Dragan S. Djordjević. Quasihyponormal and strongly quasihyponormal matrices in inner product spaces. Electronic Journal of Linear Algebra, 23(1):1023–1039, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/71.

Reams:2002:CTZ

[Rea02] Robert Reams. Constructions of trace zero symmetric stochastic matrices for the inverse eigenvalue problem. Electronic Journal of Linear Algebra, 9(1):270–275, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/22.

Reff:2014:SPO

[Ref14] Nathan Reff. Spectral properties of oriented hypergraphs. Electronic Journal of Linear Algebra, 27 (1):373–391, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/24.

Rohn:2011:IIM

[RF11] Jiri Rohn and Raena Farhadsefat. Inverse interval matrix: a survey. Electronic Journal of Linear Algebra, 22 (1):704–719, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/46.

Ren:2006:NGP

[RHC06] Zhi-Gang Ren, Ting-Zhu Huang, and Xiao-Yu Cheng. A note on generalized Perron complements of Z-matrices. Electronic Journal of Linear Algebra, 15(1):8–13, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/2.

Rojo:2009:SWC

[RM09] Oscar Rojo and Luis Medina. Spectra of weighted compound graphs of generalized Bethe trees. Electronic Journal of Linear Algebra, 18(1):30–57, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/4.

Rojo:2010:EAC

[RMdAJ10] Oscar Rojo, Luis Medina, Nair M. M. de Abreu, and Claudia Justel. Extremal algebraic connectivities of certain caterpillar classes and symmetric caterpillars. Electronic Journal of Linear Algebra, 20(1):136–157, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/11.

Rivaz:2014:ISM

[RMZ14] A. Rivaz, M. Mohseni Moghadam, and S. Zangoei Zadeh. Interval system of matrix equations with two unknown matrices. Electronic Journal of Linear Algebra, 27(1):478–488, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/30.

Rodman:2006:BSB

[Rod06] Leiba Rodman. Bounded and stably bounded palindromic difference equations of first order. Electronic Journal of Linear Algebra, 15(1):22–49, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/4.

Rodman:2007:CCS

[Rod07] Leiba X. Rodman. Comparison of congruences and strict equivalences for real, complex, and quaternionic matrix pencils with symmetries. Electronic Journal of Linear Algebra, 16(1):248–283, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/23.

Rodman:2008:HSR

[Rod08a] Leiba X. Rodman. Hamiltonian square roots of skew Hamiltonian quaternionic matrices. Electronic Journal of Linear Algebra, 17 (1):168–191, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/14.

Rodman:2008:ISP

[Rod08b] Leiba X. Rodman. Isotropic subspaces for pairs of Hermitian matrices. Electronic Journal of Linear Algebra, 17 (1):508–517, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/37.

Rodman:2014:HOM

[Rod14a] Leiba Rodman. Hermitian octonion matrices and numerical ranges. Electronic Journal of Linear Algebra, 27 (1):515–533, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/33.

Rodman:2014:INS

[Rod14b] Leiba Rodman. Invariant neutral subspaces for Hamiltonian matrices. Electronic Journal of Linear Algebra, 27(1):55–99, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol27/iss1/4.

Rohn:2009:ASA

[Roh09a] Jiri Rohn. An algorithm for solving the absolute value equation. Electronic Journal of Linear Algebra, 18 (1):589–599, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/44.

Rohn:2009:DAS

[Roh09b] Jiri Rohn. Description of all solutions of a linear complementarity problem. Electronic Journal of Linear Algebra, 18(1):246–252, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/21.

Rohn:2009:FNS

[Roh09c] Jiri Rohn. Forty necessary and sufficient conditions for regularity of interval matrices: a survey. Electronic Journal of Linear Algebra, 18 (1):500–512, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/39.

Rohn:2010:CSR

[Roh10] Jiri Rohn. A characterization of strong regularity of interval matrices. Electronic Journal of Linear Algebra, 20 (1):717–722, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/50.

Rohn:2011:EII

[Roh11] Jiri Rohn. Explicit inverse of an interval matrix with unit midpoint. Electronic Journal of Linear Algebra, 22 (1):138–150, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/8.

Rohn:2013:TMA

[Roh13] Jiri Rohn. A two-matrix alternative. Electronic Journal of Linear Algebra, 26 (1):836–841, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/56.

Rolle:2003:OSC

[Rol03] Jean-Daniel Daniel Rolle. Optimal subspaces and constrained principal component analysis. Electronic Journal of Linear Algebra, 10 (1):201–211, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol10/iss1/16.

Ran:2002:SLM

[RR02] André C. M. Ran and Martine C. B. Reurings. The symmetric linear matrix equation. Electronic Journal of Linear Algebra, 9 (1):93–107, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol9/iss1/8.

Ran:2010:RSM

[RR10] Andre C. M. Ran and Leiba X. Rodman. Ranges of Sylvester maps and a minimal rank problem. Electronic Journal of Linear Algebra, 20 (1):126–135, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/10.

Rojo:2011:SWR

[RRCM11] Oscar Rojo, Maria Robbiano, Domingos M. Cardoso, and Enide A. Martins. Spectra of weighted rooted graphs having prescribed subgraphs at some levels. Electronic Journal of Linear Algebra, 22 (1):653–671, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/43.

Rovnyak:2004:KLI

[RS04] James L. Rovnyak and Lev A. Sakhnovich. On the Kreĭn–Langer integral representation of generalized Nevanlinna functions. Electronic Journal of Linear Algebra, 11(1):1–15, 2004. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/1.

Rodman:2005:ARI

[RS05] Leiba X. Rodman and Ilya M. Spitkovsky. Analytic roots of invertible matrix functions. Electronic Journal of Linear Algebra, 13 (1):187–196, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/15.

Rodman:2011:NSP

[RŠ11] Leiba Rodman and Peter Šemrl. Neutral subspaces of pairs of symmetric/ skewsymmetric real matrices. Electronic Journal of Linear Algebra, 22(1):979–999, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/63.

Rault:2013:MEN

[RSS13] Patrick X. Rault, Tsvetanka Sendova, and Ilya M. Spitkovsky. 3-by-3 matrices with elliptical numerical range revisited. Electronic Journal of Linear Algebra, 26 (1):158–167, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol26/iss1/12.

Rump:2002:VCS

[Rum02] Siegfried M. Rump. Variational characterizations of the sign-real and the signcomplex spectral radius. Electronic Journal of Linear Algebra, 9(1):112–117, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol9/iss1/10.

Raja:2009:CGM

[RV09] P. Raja and S. M. Vaezpour. On C-commuting graph of matrix algebra. Electronic Journal of Linear Algebra, 18 (1):364–370, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol18/iss1/31.

Reams:2002:ITF

[RW02] Robert Reams and Shayne Waldron. Isometric tight frames. Electronic Journal of Linear Algebra, 9(1): 122–128, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol9/iss1/12.

Shabat:2012:IZM

[SA12] Gil Shabat and Amir Averbuch. Interest zone matrix approximation. Electronic Journal of Linear Algebra, 23 (1):678–702, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/50.

Soto:2008:NMP

[SC08] Ricardo L. Soto and Javier Ccapa. Nonnegative matrices with prescribed elementary divisors. Electronic Journal of Linear Algebra, 17 (1):287–303, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/21.

Schep:2011:BSR

[Sch11a] Anton R. Schep. Bounds on the spectral radius of Hadamard products of positive operators on ℓ_{p}-spaces. Electronic Journal of Linear Algebra, 22(1):443–447, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/26.

Scholz:2011:DAA

[Sch11b]  Lena Scholz.  A deriva-

tive array approach for linear second order differentialalgebraic systems. Electronic Journal of Linear Algebra, 22 (1):310–347, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/19.

Schomburg:2012:AEP

[Sch12] Bernd Schomburg. On the approximation of ergodic projections an d stationary distributions of stochastic matrices. Electronic Journal of Linear Algebra, 23(1): 989–1000, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/69.

Sciriha:2007:CSG

[Sci07] Irene Sciriha. A characterization of singular graphs. Electronic Journal of Linear Algebra, 16(1):451–462, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/38.

Seeger:2013:AAA

[See13] Alberto Seeger. Aperture angle analysis for ellipsoids.

Electronic Journal of Linear Algebra, 26(1):732–741, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/49.

SWGLB:2006:SSM

[Sem06] Semitransitivity Working Group at LAW’05, Bled [Slovenia]. Semitransitive subspaces of matrices. Electronic Journal of Linear Algebra, 15(1):225–238, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/18. The Work-

ing Group was coordinated by Heydar Radjavi and its members were: Janez Bernik, Roman Drnovšek, Don Hadwin, Ali Jafarian, Damjana Kokol Bukovšek, Tomaž Košir, Marjeta Kramar Fijavž, Thomas Laffey, Leo Livshits, Mitja Mastnak, Roy Meshulam, Vladimir Müller, Eric Nordgren, Jan Okniński, Matjaž Omladič, Ahmed Sourour and Richard Timoney.

Shiu:2011:MAC

[SGL11] Wai Chee Shiu, Ji-Ming Guo, and Jianxi Li. The minimum algebraic connectivity of caterpillar unicyclic graphs. Electronic Journal of Linear Algebra, 22 (1):838–848, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/56.

Shao:2012:CIS

[SGL12] Yanling Shao, Yubin Gao, and Zhongshan Li. The Mcompetition indices of symmetric primitive digraphs without loops. Electronic Journal of Linear Algebra, 23 (1):457–472, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/31.

Stevanovic:2008:MSR

[SH08] Dragan Stevanović and Pierre Hansen. The minimum spectral radius of graphs with a given clique number. Electronic Journal of Linear Algebra, 17(1):110–117, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol17/iss1/10.

Shen:2010:SME

[SH10] Shu-Qian Shen and Ting-Zhu Huang. On symmetric matrices with exactly one positive eigenvalue. Electronic Journal of Linear Algebra, 20 (1):158–167, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/12.

Shen:2012:CUG

[SH12] Xiaoling Shen and Yaoping Hou. A class of unicyclic graphs determined by their Laplacian spectrum. Electronic Journal of Linear Algebra, 23(1):375–386, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/26.

Shader:2010:EC

[Sha10] Bryan L. Shader. From the Editor-in-Chief. Electronic Journal of Linear Algebra, 21 (1):172–173, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol21/iss1/15.

Shang:2012:LBE

[Sha12] Yilun Shang. Lower bounds for the Estrada index of graphs. Electronic Journal of Linear Algebra, 23 (1):664–668, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/48.

Shader:2013:NEC

[Sha13] Bryan Shader. Note from Editor-in-Chief. Electronic Journal of Linear Algebra, 26(1):590, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/39.

Shur:2011:SCF

[Shu11] Walter Shur. A simple closed form for triangular matrix powers. Electronic Journal of Linear Algebra, 22(1): 1000–1003, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol22/iss1/64.

Shen:2012:BDE

[SHZ12] Xiaoling Shen, Yaoping Hou, and Chongyan Zhang. Bicyclic digraphs with extremal skew energy. Electronic Journal of Linear Algebra, 23 (1):340–355, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol23/iss1/24.

Stevanovic:2010:DSR

[SI10] Dragan Stevanović and Aleksandar Ilić. Distance spectral radius of trees with fixed maximum degree. Electronic Journal of Linear Algebra, 20 (1):168–179, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/13.

Singer:2012:OBR

[Sin12] Sanja Singer. Orthosymmetric block rotations. Electronic Journal of Linear Algebra, 23 (1):306–326, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/22.

Singh:2014:IIE

[Sin14] Mandeep Singh. Inequalities involving eigenvalues for difference of operator means. Electronic Journal of Linear Algebra, 27(1):557–568, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/253.

Sivasubramanian:2005:AIN

[Siv05] Sivaramakrishnan Sivasubramanian. Approximating the isoperimetric number of strongly regular graphs. Electronic Journal of Linear Algebra, 13(1):111–121, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/7.

Sidi:2000:BCT

[SK00] Avramand Sidi and Vladimir Kluzner. A Bi-CG type iterative method for Drazininverse solution of singular inconsistent nonsymmetric linear systems of arbitrary index. Electronic Journal of Linear Algebra, 6(1):72–94, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol6/iss1/8.

Shaked-Monderer:2001:MCR

[SM01] Naomi Shaked-Monderer. Minimal CP rank. Electronic Journal of Linear Algebra, 8 (1):140–157, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol8/iss1/12.

Shaked-Monderer:2004:NCR

[SM04] Naomi Shaked-Monderer. A note on the CP-rank of matrices generated by Soules matrices. Electronic Journal of Linear Algebra, 12(1):2–5, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol12/iss1/2.

Sayyed:2003:GHM

[SMB03] K. A. M. Sayyed, M. S. Metwally, and Raed S. Batahan. On generalized Hermite matrix polynomials. Electronic Journal of Linear Algebra, 10 (1):272–279, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol10/iss1/22.

Snellman:2003:MSR

[Sne03] Jan Snellman. The maximal spectral radius of a digraph with (m+1)^{2}−s edges. Electronic Journal of Linear Algebra, 10(1):179–189, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol10/iss1/14.

Soltys:2002:BAC

[Sol02] Michael Soltys. Berkowitz’s algorithm and clow sequences. Electronic Journal of Linear Algebra, 9(1):42– 54, 2002. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol9/iss1/5.

Song:2013:CWD

[Son13] Guang Jing Song. Characterization of the W -weighted Drazin inverse over the quaternion skew field with applications. Electronic Journal of Linear Algebra, 26(1):1–14, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/1.

Spitkovsky:2006:PTP

[Spi06] Ilya M. Spitkovsky. On polynomials in two projections. Electronic Journal of Linear Algebra, 15(1):154–158, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/11.

Sushama:2012:WMI

[SPS12a] Agarwal N. Sushama, K. Premakumari, and K. C. Sivakumar. Weak monotonicity of interval matrices. Electronic Journal of Linear Algebra, 25 (1):92–101, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol25/iss1/9.

Sushama:2012:WMM

[SPS12b] Agarwal N. Sushama, K. Premakumari, and K. C. Sivakumar. Weak monotonicity of matrices and subclasses of proper splittings. Electronic Journal of Linear Algebra, 24(1):83–94, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol24/iss1/8.

Soto:2011:NRP

[SRM11] Ricardo L. Soto, Oscar Rojo, and Cristina B. Manzaneda. On nonnegative realization of partitioned spectra. Electronic Journal of Linear Algebra, 22(1):557–572, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/35.

Soto:2007:SNR

[SRMB07] Ricardo L. Soto, Oscar Rojo, Julio Moro, and Alberto Borobia. Symmetric nonnegative realization of spectra. Electronic Journal of Linear Algebra, 16(1):1–18, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/1.

Saez-Schwedt:2008:MDS

[SS08a] Andrés Sáez-Schwedt. Matricial decomposition of systems over rings. Electronic Journal of Linear Algebra, 17 (1):493–507, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/36.

Sakhnovich:2008:MVT

[SS08b] Alexander L. Sakhnovich and Lev A. Sakhnovich. On a mean value theorem in the class of Herglotz functions and its applications. Electronic Journal of Linear Algebra, 17(1):102–109, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/9.

Severini:2008:FLC

[SS08c] Simone Severini and Ferenc Szöllősi. A further look into combinatorial orthogonality. Electronic Journal of Linear Algebra, 17(1):376–388, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/27.

Dokovic:2009:ROV

[D̄SS09] Dragomir Ž. D̄oković, Simone Severini, and Ferenc Szöllősi.

Rational orthogonal versus real orthogonal. Electronic Journal of Linear Algebra, 18 (1):649–673, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/48.

Soto:2010:NRC

[SSM10] Ricardo L. Soto, Mario Salas, and Cristina Manzaneda. Nonnegative realization of complex spectra. Electronic Journal of Linear Algebra, 20 (1):595–609, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/43.

Stipanovic:2004:MAA

[SST04] Dušan M. Stipanović, Sriram Shankaran, and Claire J. Tomlin. Multi-agent avoidance control using an Mmatrix property. Electronic Journal of Linear Algebra, 12(1):64–72, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol12/iss1/7.

Saunders:2004:MRC

[SSV04] B. David Saunders, Arne Storjohann, and Gilles Villard. Matrix rank certification. Electronic Journal of Linear Algebra, 11(1):16–23, 2004. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol11/iss1/2.

Stanic:2013:GSS

[Sta13] Zoran Stanić. Graphs with small spectral gap. Electronic Journal of Linear Algebra, 26 (1):417–432, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol26/iss1/28.

Stuart:1998:PTS

[Stu98] Jeffrey L. Stuart. A polynomial time spectral decomposition test for certain classes of inverse M-matrices. Electronic Journal of Linear Algebra, 3(1):129–141, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol3/iss1/12.

Sury:2000:EPH

[Sur00] B. Sury. An elementary proof of the Hilbert–Mumford criterion. Electronic Journal of Linear Algebra, 7(1): 174–177, 2000. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol7/iss1/12.

Seeger:2011:CPS

[SVP11] Alberto Seeger and José Vicente-Pérez. On cardinality of Pareto spectra. Electronic Journal of Linear Algebra, 22(1):758–766, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/50.

Salam:2011:SAH

[SW11a] A. Salam and D. S. Watkins. Structured QR algorithms for Hamiltonian symmetric matrices. Electronic Journal of Linear Algebra, 22 (1):573–585, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/36.

Song:2011:PRP

[SW11b] Guang Jing Song and Qing Wen Wang. Positive and repositive solutions to some systems of adjointable operator equations over Hilbert C^{∗}-modules. Electronic Journal of Linear Algebra, 22(1): 1070–1084, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/70.

Shan:2014:EIC

[SWK14] Erfang Shan, Hongzhuan Wang, and Liying Kang. On the Estrada index of cacti. Electronic Journal of Linear Algebra, 27(1):798–808, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/268.

Dokovic:2003:ACS

[D̄SZ03] Dragomir Ž. D̄oković, F. Szechtman, and K. Zhao. An algorithm that carries a square matrix into its transpose by an involutory congruence transformation. Electronic Journal of Linear Algebra, 10 (1):320–340, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/25.

Stitzinger:2009:RCN

[SZ09] Ernie Stitzinger and Laurie Zack. Realizations of the complex nilpotent Lie algebras with small second derived quotient. Electronic Journal of Linear Algebra, 18 (1):513–515, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/40.

Szechtman:2005:SGP

[Sze05] Fernando Szechtman. Structure of the group preserving a bilinear form. Electronic Journal of Linear Algebra, 13 (1):197–239, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/16.

Szechtman:2007:ECM

[Sze07] Fernando Szechtman. Equivalence and congruence of matrices under the action of standard parabolic subgroups. Electronic Journal of Linear Algebra, 16 (1):325–333, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/28.

Dokovic:2001:NGA

[D̄T01] Dragomir Ž. D̄oković and Peter W. Tingley. Natural group actions on tensor products of three real vector spaces with finitely many orbits. Electronic Journal of Linear Algebra, 8(1):60–82, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol8/iss1/6.

Tam:1999:ERL

[Tam99] Tin-Yau Tam. An extension of a result of Lewis. Electronic Journal of Linear Algebra, 5(1):1–10, 1999. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol5/iss1/1.

Tam:2010:HRM

[Tam10] Tin-Yau Tam. A. Horn’s result on matrices with prescribed singular values and eigenvalues. Electronic Journal of Linear Algebra, 21 (1):25–27, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol21/iss1/4.

Tao:2011:PRP

[Tao11] Jiyuan Tao. Pseudomonotonicity and related properties in Euclidean Jordan algebras. Electronic Journal of Linear Algebra, 22 (1):225–251, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/15.

Tosic:2011:MPI

[TCID11] Marina Tošić, Dragana S.

Cvetković-Ilić, and Chunyuan Deng. The Moore– Penrose inverse of a linear combination of commuting generalized and hypergeneralized projectors. Electronic Journal of Linear Algebra, 22 (1):1129–1137, 2011. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/74.

Chien:2012:SPT

[tCN12] Mao ting Chien and Hiroshi Nakazato. Singular points of the ternary polynomials associated with 4-by-4 matrices. Electronic Journal of Linear Algebra, 23(1):755–769, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/53.

Tisseur:2006:SCN

[TG06] Françoise Tisseur and Stef Graillat. Structured condition numbers and backward errors in scalar product spaces. Electronic Journal of Linear Algebra, 15 (1):159–177, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol15/iss1/12.

Tian:2010:IME

[TH10] Gui-Xian Tian and Ting-Zhu Huang. Inequalities for the minimum eigenvalue of Mmatrices. Electronic Journal of Linear Algebra, 20 (1):291–302, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol20/iss1/21.

Editors:1998:D

[The98] The Editors. Dedication. Electronic Journal of Linear Algebra, 3(1):1, 1998. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol3/iss1/1.

Tian:2005:SFG

[Tia05] Yongge Tian. Special forms of generalized inverses of row block matrices. Electronic Journal of Linear Algebra, 13 (1):249–261, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/18.

Tian:2010:CBH

[Tia10a]  Yongge Tian.  Complet-

ing block Hermitian matrices with maximal and minimal ranks and inertias. Electronic Journal of Linear Algebra, 21 (1):124–141, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol21/iss1/12.

Tian:2010:RIS

[Tia10b] Yongge Tian. Rank and inertia of submatrices of the Moore–Penrose inverse of a Hermitian matrix. Electronic Journal of Linear Algebra, 20 (1):226–240, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/17.

Tian:2012:EFI

[Tia12] Yongge Tian. On an equality and four inequalities for generalized inverses of Hermitian matrices. Electronic Journal of Linear Algebra, 23(1):11– 42, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/2.

Tian:2007:GMT

[TL07] Yongge Tian and Yonghui Liu. On a group of mixedtype reverse-order laws for generalized inverses of a triple matrix product with applications. Electronic Journal of Linear Algebra, 16(1):73–89, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/7.

Tlusty:2007:RBM

[Tlu07] Tsvi Tlusty. A relation between the multiplicity of the second eigenvalue of a graph Laplacian, Courant’s nodal line theorem and the substantial dimension of tight polyhedral surfaces. Electronic Journal of Linear Algebra, 16(1):315–324, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol16/iss1/27.

Taher:2006:FHD

[TMR06] R. Ben Taher, M. Mouline, and Mustapha Rachidi. Fibonacci–Horner decomposition of the matrix exponential and the fundamental system of solutions. Electronic Journal of Linear Algebra, 15 (1):178–190, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol15/iss1/13.

Triantafillou:2013:ESG

[Tri13] Irene Triantafillou. On the energy of singular graphs.

Electronic Journal of Linear Algebra, 26(1):535–545, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/34.

Tian:2001:HEU

[TS01] Yongge Tian and George P. H. Styan. How to establish universal block-matrix factorizations. Electronic Journal of Linear Algebra, 8(1): 115–127, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol8/iss1/10.

Tian:2005:SCB

[TT05] Yongge Tian and Yoshio Takane. Schur complements and Banachiewicz– Schur forms. Electronic Journal of Linear Algebra, 13 (1):405–418, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol13/iss1/25.

Tian:2007:SAS

[TT07] Yongge Tian and Yoshio Takane. Some algebraic and statistical properties of WLSEs under a general growth curve model. Electronic Journal of Linear Algebra, 16(1):187–203, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/17.

Tyaglov:2012:SPS

[Tya12] Mikhail Tyaglov. Sign patterns of the Schwarz matrices and generalized Hurwitz polynomials. Electronic Journal of Linear Algebra, 24 (1):215–236, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol24/iss1/16.

Uhlig:1998:PCS

[Uhl98] Frank Uhlig. Partial convergence and sign convergence of matrix powers via eigen cads. Electronic Journal of Linear Algebra, 3(1):75–92, 1998. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol3/iss1/8.

vanderAa:2007:CEE

[vdAtMM07] N. P. van der Aa, H. G. ter Morsche, and R. R. M. Mattheij. Computation of eigenvalue and eigenvector derivatives for a general complex-valued eigensystem. Electronic Journal of Linear Algebra, 16(1):300–314, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol16/iss1/26.

vandenDriessche:1996:EPD

[vdDW96] Pauline van den Driessche and Harald K. Wimmer. Explicit polar decomposition of companion matrices. Electronic Journal of Linear Algebra, 1(1):64–69, 1996. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol1/iss1/6.

vanderHolst:2009:MPS

[vdH09] Hein van der Holst. On the maximum positive semidefinite nullity and the cycle matroid of graphs. Electronic Journal of Linear Algebra, 18 (1):192–201, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/16.

vanderHolst:2010:SAH

[vdH10] Hein van der Holst. On the strong Arnol’d hypothesis and the connectivity of graphs. Electronic Journal of Linear Algebra, 20 (1):574–585, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/41.

Vermeer:2008:CSC

[Ver08] J. Vermeer. On (con)similarities and congruences between A and A^{∗}, A^{T} or Ā. Electronic Journal of Linear Algebra, 17 (1):258–283, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/19.

Vandanjav:2012:NRW

[VU12] Adiyasuren Vandanjav and Batzorig Undrakh. On the numerical ranges of the weighted shift operators with geometric and harmonic weights. Electronic Journal of Linear Algebra, 23 (1):578–585, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/41.

Meulen:2009:ZNP

[VV09] Kevin N. Vander Meulen and Adam Van Tuyl. Zerononzero patterns for nilpotent matrices over finite fields. Electronic Journal of Linear Algebra, 18(1): 628–648, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/47.

Vlasev:2012:FVQ

[VY12] Aleksandar Vlasev and Karen Yeats. A four-vertex, quadratic, spanning forest polynomial identity. Electronic Journal of Linear Algebra, 23(1):923–941, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/65.

Wanless:2008:BLC

[Wan08] Ian M. Wanless. On the Brualdi–Liu conjecture for the even permanent. Electronic Journal of Linear Algebra, 17(1):284–286, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/20.

Wan:2014:NRN

[WHT14] Jixiang Wan, Shuangnian Huang, and Qianrong Tan. New results on nonsingular power LCM matrices. Electronic Journal of Linear Algebra, 27(1):652–669, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/258.

Wang:2010:ERS

[WJ10] Qing Wen Wang and Jing Jiang. Extreme ranks of (skew-)Hermitian solutions to a quaternion matrix equation. Electronic Journal of Linear Algebra, 20(1): 552–573, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/40.

Woznicki:2001:BCT

[Woź01] Zbigniew I. Woźnicki. Basic comparison theorems for weak and weaker matrix splittings. Electronic Journal of Linear Algebra, 8(1):53– 59, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol8/iss1/5.

Wu:2012:ODS

[WS12] Yaokun Wu and Jian Shen. Opinion dynamics with stubborn vertices. Electronic Journal of Linear Algebra, 23 (1):790–800, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/56.

Wang:2012:NGL

[WSH12] Jianfeng Wang, Yufa Shen, and Qiongxiang Huang. Notes on graphs with least eigenvalue at least −2. Electronic Journal of Linear Algebra, 23(1):387–396, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/27.

Wei:2001:ARI

[WW01] Yimin Wei and Hebing Wu. Additional results on index splittings for Drazin inverse solutions of singular linear systems. Electronic Journal of Linear Algebra, 8(1):83– 93, 2001. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol8/iss1/7.

Wang:2011:NRM

[WW11] Hongxing Wang and Musheng Wei. New representations for the Moore–Penrose inverse. Electronic Journal of Linear Algebra, 22(1):720–728, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/47.

Wu:2010:MPB

[WWY10] Huazhang Wu, Xiaoxuan Wu, and Zhenghong Yang. More on polynomial Bézoutians with respect to a general basis. Electronic Journal of Linear Algebra, 21(1): 154–171, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol21/iss1/14.

Wang:2008:RRN

[WZ08] Qing-Wen Wang and Fei Zhang. The reflexive renonnegative definite solution to a quaternion matrix equation. Electronic Journal of Linear Algebra, 17 (1):88–101, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/8.

Wu:2012:IMC

[WZ12] Yaokun Wu and Shizhen Zhao. Incidence matrix and cover matrix of nested interval orders. Electronic Journal of Linear Algebra, 23(1):43– 65, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/3.

Wang:2009:NSC

[WZS09] Qing-Wen Wen Wang, Hua-Sheng Zhang, and Guang-Jing Song. A new solvable condition for a pair of generalized Sylvester equations. Electronic Journal of Linear Algebra, 18(1):289–301, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol18/iss1/26.

Wang:2008:SQM

[WZY08] Qing-Wen Wang, Hua-Sheng Zhang, and Shao-Wen Yu. On solutions to the quaternion matrix equation AXB + CY D = E. Electronic Journal of Linear Algebra, 17 (1):343–358, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol17/iss1/25.

Xu:2011:OPB

[XCL11] Wei-Wei Xu, Li-Xia Cai, and Wen Li. The optimal perturbation bounds for the weighted Moore–Penrose inverse. Electronic Journal of Linear Algebra, 22 (1):521–538, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/33.

Xu:2013:ELE

[XD13] Kexiang Xu and Kinkar Ch.

Das. Extremal Laplacianenergy-like invariant of graphs with given matching number. Electronic Journal of Linear Algebra, 26(1):131–140, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/10.

Xiao:2009:ASO

[XHZ09] Qing-Feng Xiao, Xi-Yan Hu, and Lei Zhang. The antisymmetric ortho-symmetric solutions of the matrix equation A^{T} XA = D. Electronic Journal of Linear Algebra, 18(1):21–29, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/3.

Xie:2013:CNC

[XLJ13] Ze-Jia Xie, Wen Li, and Xiao-Qing Jin. On condition numbers for the canonical generalized polar decomposition of real matrices. Electronic Journal of Linear Algebra, 26(1):842–857, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/57.

Xiong:2011:IPO

[XQ11] Zhiping Xiong and Yingying Qin. Invariance properties of an operator product involving generalized inverses. Electronic Journal of Linear Algebra, 22(1):694–703, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/45.

Xu:2010:IPM

[XTC10] Jin-Li Xu, Xiao-Min Tang, and Chong-Guang Cao. Idempotence-preserving maps between matrix spaces over fields of characteristic 2. Electronic Journal of Linear Algebra, 21(1):110–123, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol21/iss1/11.

Xu:2008:MCS

[Xu08]  Jianhong Xu.  Markov

chain small-world model with asymmetric transition probabilities. Electronic Journal of Linear Algebra, 17 (1):616–636, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/43.

Xu:2010:TCJ

[Xu10] Jianhong Xu. On the trace characterization of the joint spectral radius. Electronic Journal of Linear Algebra, 20 (1):367–375, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/28.

Xu:2014:CMR

[XY14] Xiao-Wei Xu and Xiaofei Yi.

Commuting maps on rank-k matrices. Electronic Journal of Linear Algebra, 27 (1):735–741, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/264.

Xu:2010:JLD

[XZ10] Xiao-Wei Wei Xu and Hong-Ying Ying Zhang. Jordan left derivations in full and upper triangular matrix rings. Electronic Journal of Linear Algebra, 20(1):753–759, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/54.

Xiong:2011:ROL

[XZ11] Zhiping Xiong and Bing Zheng. The reverse order laws and the mixed-type reverse order laws for generalized inverses of multiple matrix products. Electronic Journal of Linear Algebra, 22 (1):1085–1105, 2011. CO-DEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/71.

Yang:2011:SCI

[YF11] Zhongpeng Yang and Xiaoxia Feng. On a Schur complement inequality for the Hadamard product of certain totally nonnegative matrices. Electronic Journal of Linear Algebra, 22(1):214–224, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/14.

Yu:2014:QFG

[YFW14] Gui-Dong Yu, Yi-Zheng Fan, and Yi Wang. Quadratic forms on graphs with application to minimizing the least eigenvalue of signless Laplacian over bicyclic graphs. Electronic Journal of Linear Algebra, 27(1):213–236, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol27/iss1/13.

Yu:2013:LSL

[YGX13] Guanglong Yu, Shuguang Guo, and Meiling Xu. On the least signless Laplacian eigenvalue of some graphs. Electronic Journal of Linear Algebra, 26(1):560–573, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/37.

Yang:2005:NES

[YH05] Shi-Ming Yang and Ting-Zhu Huang. A note on estimates for the spectral radius of a nonnegative matrix. Electronic Journal of Linear Algebra, 13(1):352–358, 2005. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol13/iss1/22.

You:2011:SLS

[YL11] Zhifu You and Bolian Liu.

The signless Laplacian separator of graphs. Electronic Journal of Linear Algebra, 22 (1):151–160, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/9.

Yao:2010:PDS

[YLD10] Guozhu Yao, Anping Liao, and Xuefeng Duan. Positive definite solution of the matrix equation X = Q + A^{H}(I ⊗ X − C)^{(} − δ)A. Electronic Journal of Linear Algebra, 21(1):76–84, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/8.

Yu:2012:BSR

[YSH12] Guanglong Yu, Jinlong Shu, and Yuan Hong. Bounds of spectral radii of K_{2,3}-minor free graphs. Electronic Journal of Linear Algebra, 23 (1):171–179, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/12.

Yu:2010:NUB

[Yu10]  Aimei Yu.  A new up-

per bound for the Laplacian spectral radius of a graph. Electronic Journal of Linear Algebra, 20(1):730–738, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/52.

Yuan:2012:SRG

[YWZ12] Wanlian Yuan, Bing Wang, and Mingqing Zhai. On the spectral radii of graphs without given cycles. Electronic Journal of Linear Algebra, 23 (1):599–606, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/43.

You:2012:CPS

[YY12] Lihua You and Shuyong Yi. The characterization of primitive symmetric loopfree signed digraphs with the maximum base. Electronic Journal of Linear Algebra, 23 (1):122–136, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/8.

Yang:2013:DGA

[YY13] Qingjie Yang and Dan Yang.

Dihedral groups of automorphisms of compact Riemann surfaces of genus two. Electronic Journal of Linear Algebra, 26(1):553–559, 2013. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol26/iss1/36.

Zhang:2009:SCG

[yZLXJ09] Cheng yi Zhang, Shuanghua Luo, Chengxian Xu, and Hongying Jiang. Schur complements of generally diagonally dominant matrices and a criterion for irreducibility of matrices. Electronic Journal of Linear Algebra, 18 (1):69–87, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol18/iss1/7.

Zhao:2010:GIB

[ZB10] Jiemei Zhao and Changjiang Bu. Group inverse for the block matrix with two identical subblocks over skew fields. Electronic Journal of Linear Algebra, 21(1):63–75, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol21/iss1/7.

Zhu:2003:PCM

[ZCB03] Zewen Zhu, Daniel C. Coster, and Leroy B. Beasley. Properties of a covariance matrix with an application to Doptimal design. Electronic Journal of Linear Algebra, 10(1):65–76, 2003. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol10/iss1/6.

Zhao:2010:LSO

[ZCL10] Lin-Lin Zhao, Guo-Liang Chen, and Qing-Bing Liu. Least squares (P, Q)-orthogonal symmetric solutions of the matrix equation and its optimal approximation. Electronic Journal of Linear Algebra, 20(1):537–551, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol20/iss1/39.

Zhu:2007:SSD

[ZH07] Yan Zhu and Ting-Zhu Huang. Subdirect sums of doubly diagonally dominant matrices. Electronic Journal of Linear Algebra, 16 (1):171–182, 2007. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol16/iss1/15.

Zhang:2014:DSR

[Zha14] Xiao Ling Zhang. On the distance spectral radius of unicyclic graphs with perfect matchings. Electronic Journal of Linear Algebra, 27 (1):569–587, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/254.

Zhong:2012:BRB

[Zho12] Jin Zhong. Binary ranks and binary factorizations of nonnegative integer matrices. Electronic Journal of Linear Algebra, 23(1):540–552, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/38.

Zhang:2008:PLB

[ZHS08] Wei Zhang, Zheng-Zhi Han, and Shu-Qian Shen. A parameterized lower bound for the smallest singular value. Electronic Journal of Linear Algebra, 17(1):487–492, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol17/iss1/35.

Zhu:2011:SLS

[Zhu11] Zhongxun Zhu. The signless Laplacian spectral radius of bicyclic graphs with a given girth. Electronic Journal of Linear Algebra, 22 (1):378–388, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol22/iss1/22.

Zhu:2012:NLE

[Zhu12] Bao-Xuan Zhu. A note on the least eigenvalue of a graph with given maximum degree. Electronic Journal of Linear Algebra, 23(1):514–522, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol23/iss1/35.

Zhao:2011:SIG

[ZJW11] Meixiang Zhao, Zhigang Jia, and Musheng Wei. Symmetric inverse generalized eigenvalue problem with submatrix constraints in structural dynamic model updating. Electronic Journal of Linear Algebra, 22(1):729–742, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/48.

Zhang:2010:NES

[ZL10] Juan Zhang and Jianzhou Liu. New estimates for the solution of the Lyapunov matrix differential equation. Electronic Journal of Linear Algebra, 20(1):6– 19, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol20/iss1/2.

Zhang:2011:NLS

[ZL11] Juan Zhang and Jianzhou Liu. New lower solution bounds for the continuous algebraic Riccati equation. Electronic Journal of Linear Algebra, 22(1):191–202, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/12.

Zhang:2010:EDB

[ZLHL10] Cheng-Yi Zhang, Shuanghua Luo, Aiqun Huang, and Junxiang Lu. The eigenvalue distribution of block diagonally dominant matrices and block H-matrices. Electronic Journal of Linear Algebra, 20 (1):621–639, 2010. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol20/iss1/45.

Zhang:2012:ECM

[ZLLX12] Cheng-Yi Zhang, Shuanghua Luo, Jicheng Li, and Fengmin Xu. An extension of the class of matrices arising in the numerical solution of PDEs. Electronic Journal of Linear Algebra, 23 (1):422–444, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol23/iss1/29.

Zhang:2009:EDS

[ZLXX09] Cheng-Yi Zhang, Shuanghua Luo, Fengmin Xu, and Chengxian Xu. The eigenvalue distribution of Schur complements of nonstrictly diagonally dominant matrices and general H-matrices. Electronic Journal of Linear Algebra, 18(1):801–820, 2009. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol18/iss1/58.

Zheng:2008:NEC

[ZX08] Bing Zheng and Zhiping Xiong. A new equivalent condition of the reverse order law for G-inverses of multiple matrix products. Electronic Journal of Linear Algebra, 17(1):1–8, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

                                        http://repository.uwyo.
                                        edu/ela/vol17/iss1/1.

Zhang:2011:LSR

[ZZ11] Guang-Jun Zhang and Xiao-Dong Zhang. The p-Laplacian spectral radius of weighted trees with a degree sequence and a weight set. Electronic Journal of Linear Algebra, 22(1):267–276, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

                                        repository.uwyo.edu/ela/
                                        vol22/iss1/17.

Zhang:2014:PNN

[ZZ14] Yun Zhang and Jin Zhong. Possible numbers of nonzero entries in a matrix with a given term rank. Electronic Journal of Linear Algebra, 27 (1):258–271, 2014. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL

http://repository.uwyo.
edu/ela/vol27/iss1/16.

Zhang:2006:BSR

[ZzH06] Wei Zhang and Zheng zhi Han. Bounds for the spectral radius of block H-matrices. Electronic Journal of Linear Algebra, 15(1):269–273, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol15/iss1/22.

Zhu:2011:PPI

[ZZL11] Yan Zhu, Cheng-Yi Zhang, and Jun Liu. Path product and inverse M-matrices. Electronic Journal of Linear Algebra, 22(1):644–652, 2011. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol22/iss1/42.

Zhang:2012:GSC

[ZZLL12] Cheng-Yi Zhang, Yan Zhu, Shuanghua Luo, and Jicheng Li. On generalized Schur complement of nonstrictly diagonally dominant matrices and general H-matrices. Electronic Journal of Linear Algebra, 23(1):801–814, 2012. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://

repository.uwyo.edu/ela/
vol23/iss1/57.